### v1

# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

Abstract
We give an overview of the theoretical description of nuclear reactions involving weakly-bound nuclei. Some of the more widespread reaction formalisms employed in the analysis of these reactions are briefly introduced, including various recent developments. We put special emphasis on the continuum-discretized coupled-channel (CDCC) method and its extensions to incorporate core and target excitations as well as its application to three-body projectiles. The role of the continuum for one-nucleon transfer reactions is also discussed. The problem of the evaluation of inclusive breakup cross sections is addressed within the Ichimura-Austern-Vincent (IAV) model. Other methods, such as those based on a semiclasical description of the scattering process, are also briefly introduced and some of their applications are discussed and a brief discussion on topics of current interest, such as nucleon-nucleon correlations, uncertainty evaluation and non-locality is presented.
pacs:
24.10.-i Nuclear reaction models and methods and 24.50.+gDirect reactions and 25.45.De Elastic and inelastic scattering and 25.70.DeCoulomb excitation and 25.70.Hi Transfer reactions
1 Introduction
Nuclear reactions are key tools to extract information about the structure of atomic nuclei, and the dynamical phenomena arising from the nucleus-nucleus forces. When one of the colliding partners is weakly bound, new phenomena and mechanisms arise, requiring in some cases specific reaction frameworks tailored to the peculiarities of these systems and their interactions.
Weakly-bound nuclei appear in the proximity to the neutron and proton driplines, a region where new exotic structures and phenomena are found. Promiment examples are halo nuclei, weakly bound nuclei composed of a compact core and one or two loosely bound nucleons with an unusually large matter radius, or Borromean systems, three-body systems with no bound binary subsystems, such as ${}^{9}\text{Be}~{}(\alpha+\alpha+n)$ or ${}^{6}\text{He}~{}(\alpha+n+n)$.
Although the field has experienced a great impulse in the last decades, the physics of nuclear scattering with weakly-bound projectiles is not new and many of the problems that are being addressed were already recognized much earlier as, for example, in the context of low-energy deuteron scattering. The deuteron, while being a stable nucleus, displays many halo-like features, such as weak binding ($E_{b}=2.22$ MeV), large spatial extension (the proton-neutron separation is about 3.8 fm) and no bound excited states. Let us consider the scattering of low-energy deuterons (a few MeV) by a heavy target nucleus, like 208Pb. Initially, when the deuteron is far apart from the target nucleus, it is in an internal state given by the ground-state of the proton-neutron Hamiltonian. As the deuteron approaches the target, it will feel its Coulomb repulsion. If the deuteron were a point-like particle, the effect of this interaction would be to distort the trajectory of the deuteron, without altering its internal structure. However, this is not the case. The Coulomb interaction acts on the proton, whose distance from the center-of-mass of the deuteron is $\sim$2 fm. Consequently, in addition to the monopole Coulomb potential ($\propto 1/R$) the deuteron will feel higher-order Coulomb multipoles arising from the expansion, valid for $r\ll R$, given by Cle62b
$V_{C}(\vec{r},\vec{R})=\frac{Z_{t}e^{2}}{|\vec{R}+\vec{r}/2|}=\frac{Z_{t}e^{2}}{R}-\frac{Z_{t}e^{2}}{2R^{2}}r\cos(\theta)+\ldots$ (1)
where $\vec{R}$ is the coordinate from the target to the deuteron c.m.  $\vec{r}$ the proton-neutron relative coordinate and $\theta$ is the angle between them (see Fig. 1). The main deviation from the point Coulomb interaction is caused by the dipole term which depends on the orientation of the proton-neutron relative coordinate with respect to the deuteron-target coordinate. Specifically, when the proton is closer to the target nucleus ($\cos(\theta)<0$), the dipole potential will add a positive contribution, whereas when the proton is farther from the c.m. of the deuteron the contribution will be negative. Classically, we may think of this problem as an electric dipole moving in a slowly-varying electric field, in which the tidal force exerted on the dipole will favour the configuration of Fig. 1. Quantum-mechanically, the perturbation induced by the dipole force will couple the deuteron ground state (positive parity) with negative-parity states. Since the ground-state is the only bound state, these states necessarily appear in the continuum. Therefore, the polarization effect will modify the state of the system, producing a new one in which the ground state is mixed with continuum states of negative parity. This has a twofold effect on the outcome of the scattering process. First, it will produce deviations of the elastic scattering with respect to the Rutherford formula. Second, the coupling with the positive-energy states will give rise to some dissociation probability, that is, the breakup of the deuteron into a proton and a neutron.
The effect of dipole polarizability can be included by means of a polarization potential that, added to the monopole Coulomb potential, describes the motion of the center of mass of the nucleus. In the adiabatic limit, it is assumed that the excitation energies are high enough so the characteristic time for a transition to a state is small compared to the characteristic time for the collision. Applying second-order perturbation theory, one gets the following expression for this adiabatic polarization potential Ald75 :
$V_{\mathrm{ad}}(R)=-\sum_{n\neq 0}\frac{|\langle n|V_{\mathrm{dip}}|0\rangle|^{2}}{E_{n}-E_{0}}=-\frac{1}{2}\alpha\frac{(Z_{T}e)^{2}}{R^{4}},$ (2)
where $V_{\mathrm{dip}}$ is the second term of Eq. (1) and $\alpha$ is the dipole polarizability parameter Cle62b , which measures the electric response of the nucleus (the deuteron in this case) to an external electric field. It is therefore a structure property of any nucleus.
To conclude with this example, it is to be noted that the adiabiatic polarization potential given by Eq. (2) is purely real and, as such, does not describe the effect of the deuteron breakup that would remove flux from the elastic channel. This point will be revisited in Sec. 7.3, where another version of the polarization potential, which accounts for the deuteron breakup, will be presented.
The polarization potential adds an attractive contribution to the Coulomb point-particle potential that, when used in the Schrödinder equation, will produce a small, but measurable deviation of the elastic cross section with respect to the well-known Rutherford formula. An accurate measurement of sub-Coulomb elastic scattering data can therefore be used to extract the dipole polarizability of the deuteron (and other polarizable systems). The idea has in fact been successfully applied to the case of deuteron scattering Rod82 and constitutes a neat and beautiful example in which an important structure quantity can be inferred from reaction observables.
Figure 1: Relevant coordinates for a deuteron+target scattering problem.
The discussed example, albeit describing a specific problem, exhibits some general features commonly found in the scattering of weakly bound nuclei. In particular, the coupling to the breakup channels will play a role, to a larger or smaller extent, in essentially all reaction observables. Therefore, reaction models employed to describe these observables will have to incorporate this effect.
We enumerate some fingerprints of the weak binding on reaction observables:
  • Large interaction cross sections in nuclear collisions at high energies. Historically, the first evidence of the unusual properties of halo nuclei came from the pioneering experiments performed by Tanihata and co-workers at Berkeley using very energetic (800 MeV/nucleon) secondary beams of radioactive species Tan85a ; Tan85b . At these high energies, interaction cross sections are approximately proportional to the size of the colliding nuclei. It was found that some exotic isotopes of light nuclei (6He, 11Li, 14Be) presented much higher interaction cross sections than their neighbour isotopes, which was interpreted as an abnormally large radius.
  • Narrow momentum distributions of residues following fast nucleon removal. Momentum distributions of the residual nucleus following the removal of one or more nucleons of a energetic projectile colliding with a target nucleus are closely related to the momentum distribution of the removed nucleon(s) in the original projectile. Kobayashi et al. Kob88 found that the momentum distributions of 9Li following the fragmentation process ${}^{11}\text{Li}+{{}^{12}}\text{C}\rightarrow{{}^{9}}\text{Li}+\text{X}$ were abnormally narrow, which, according to the Heisenberg’s uncertainty principle, suggested a long tail in the density distribution of the 11Li nucleus. This result was later found in other weakly bound nuclei.
  • Abnormal elastic scattering cross sections. Elastic scattering is affected by the coupling to non-elastic processes. In particular, when coupling to breakup channels is important, elastic scattering cross sections are depleted with respect to the case of tightly bound nuclei. Some other key signatures are the departure of the elastic cross section with respect to the Rutherford cross section at sub-Coulomb energies and the disappearance of the Fresnel peak at near-barrier energies in reactions induced by halo nuclei on heavy targets San08 ; Pie12 ; Cub12 ; Pes17 .
  • Enhanced near-threshold breakup cross section in Coulomb dissociation experiments of neutron-halo nuclei. When a neutron-halo nucleus, composed of a charged core and one or two weakly-bound neutrons (11Be, 6He, 11Li,…) collides with a high-$Z$ target nucleus, the projectile structure is heavily distorted due to the tidal force originated from the uneven action of the Coulomb interaction on the charged core and the neutrons. This produces a stretching which may eventually break up the loosely bound projectile. This gives rise to a large population of the continuum states close to the breakup threshold.
  • Complete fusion suppression
    Experiments with weakly bound light stable nuclei (such as 6,7Li and 9Be) have shown a systematic suppression of complete fusion (CF) cross sections (defined as capture of the complete charge of the projectile) of $\sim$20-30compared to the case of tightly bound nuclei Das99 ; Tri02 ; Das02 ; Das04 ; Muk06 ; Rat09 ; LFC15 . The effect has been attributed to the presence of strong competing channels, such as the breakup of the weakly bound projectile prior to reaching the fusion barrier, with the subsequent reduction of capture probability. This interpretation is supported by the presence of large $\alpha$ yields as well as target-like residues which are consistent with the capture of one of the fragment constituents of the projectile, a process which is usually termed as incomplete fusion (ICF).
A proper and quantitative understanding of these and other phenomena requires the use of an appropriate reaction theory. When dealing with weakly-bound systems, one expects a certain decoupling between the degrees of freedom describing the relative motion between the weakly-bound nucleon (or cluster) from that of the internal excitations of the clusters themselves. Following this argument, one may be tempted to adopt an extreme model in which only the degree of freedom for the relative motion of the weakly-bound nucleon(s) or clusters are considered, while the others are simply ignored or frozen. This approach, which reminds of the separation between active nucleons and core nucleons in the shell-model, has in fact become very useful to understand the main features of the structure of halo, and other weakly-bound, nuclei and their dynamics. In some cases, such as the deuteron system or one-nucleon halos, such as 19C or 11Be, a two-body model may provide a reasonable starting point for the description of these nuclei. However, in some other systems, such as the Borromean nuclei 11Li, 6He or 9Be, it will be in general mandatory to resort to (at least) a three-body model for a meaningful description of their structure and reactions.
One of the most successful models for describing reactions involving weakly-bound nuclei, which incorporates in a natural way the few-body structure of these systemas as well as its breakup is the Continuum Discretized Coupled-Channels (CDCC) Aus87 . The model and its recent extensions will be discussed in Sec. 3. Other models, tailored to specific processes will be discussed along this review.
It is the purpose of this paper to review some of the theories developed and applied for describing such nuclear reactions with weakly bound nuclei. As any review, the present one will be necessarily incomplete and possibly biased by the expertise and personal taste of the authors with the hope that the reader will obtain, at least, a flavour of how reaction theory has helped the interpretation of experiments with these nuclear species.
2 The impact of weak-binding on the elastic cross section: optical model approach
Elastic scattering provides a valuable source of information on the structure and dynamics of nuclei. As explained in the Introduction, it was long realised that the elastic scattering of deuterons, admittedly the simplest example of weakly bound nucleus, does not follow the expected behavior for tightly bound nuclei.
As in the case of well-bound nuclei, the natural framework to study elastic scattering is the Optical Model (OM), which consists in solving a two-body Schrödinger equation with an effective nucleus-nucleus potential, referred to as the optical model potential (OMP) Sat83 . The form of this OMP can be formally derived from the microscopic nucleus-nucleus interaction. In his seminal work, Feshbach Fesh58 showed that the OMP can be formally written as a sum of two terms, a bare potential, which is the expectation value of the nucleus-nucleus potential in the ground state of the projectile+target system, and an additional term, the polarization potential, which accounts for the effect of non-elastic channels (inelastic scattering, transfer, breakup, fusion, …) on the elastic scattering: Explicitly:
$\mathcal{V}=V_{00}+\mathbf{V}_{0}\frac{1}{E^{(+)}-\mathbf{H}}\mathbf{V}_{0}^{\dagger}.$ (3)
where the first and second terms correspond, respectively, to the bare and polarization potentials. Here, $\mathbf{H}$ stands for the full, microscopic Hamiltonian, $\mathbf{V}_{0}$ is an operator with non-vanishing matrix elements between the ground-state and other states of the system , $\mathbf{V}_{0}=(V_{01},V_{02},\ldots)$. This effective effective potential gives rise to the Schrödinger equation for the relative-motion wavefunction:
$\left[T_{\vec{R}}+\mathcal{V}-E\right]\chi_{0}(\vec{R})=0$ (4)
where $T_{\vec{R}}$ is the kinetic energy operator and $\chi_{0}$ the wave-function for the relative motion between projectile and target in their ground states.
The effect of weak binding of any of the colliding partners affects both the bare and polarization potentials. The ground-state of the projectile will exhibit an extended tail (as compared to the case of tightly-bound nuclei) and this will affect the form of $V_{00}$. In addition, the weak-binding will enhance certain couplings described by the $\mathbf{V}_{0}$ operator, very prominently the breakup channels, and this will also modify the polarization term. An example is the Coulomb polarization potential described in the Introduction.
Most practical applications of the OM rely on approximate forms of the Feshbach potential, consisting typically of a central potential (with possibly spin-orbit and tensor terms) whose radial parts are parametrized in terms of simple analytical forms, such as the popular Woods-Saxon potential. Because the Feshbach potential is complex, so is the phenomenological optical potential. In principle, the potential should be also non-local and angular-momentum and energy dependent. While the energy dependence is a common feature of phenomenological potentials, nonlocalilty and angular-momentum dependence is more rarely taken into account (see Sec. 10).
In the presence of strong absorption, elastic scattering angular distributions between composite nuclei display some common features regardless of the interacting nuclei. At near-barrier energies, elastic distributions present two distinct regions. At small angles, where the Coulomb interaction dominates, the cross section remains close to the Rutherford prediction, but oscillates around it. The amplitudes of the oscillations increases as the scattering angle increases but remain roughly 25% within the Rutherford prediction. This is the so-called “illuminated” region. After this maximum, the cross section drops rapidly, becoming much smaller than the Rutherford cross sections at large angles (the “shadow” region). An example of this behaviour can be seen in the elastic scattering data of $\alpha$ particles on lead, shown in the left panel of Fig. 2.
These features have been commonly interpreted by invoking classical models. The oscillatory behaviour in the illuminated region can be qualitatively interpreted within a Fresnel picture as an interference between distant pure Coulomb trajectories with closer trajectories affected by the nuclear attractive potential, while the shadow region can be understood as a consequence of the existence of a maximum angle in the classical deflection function. A complete understanding of the observed pattern requires also the introduction of refractive effects which arise in nuclear scattering as a result of absorption Fri89 .
All these effects are naturally accommodated within the OM. For example, the solid line in the left panel of Fig. 2 represents an OM fit of the data using a standard (complex) Woods-Saxon potential for the real and imaginary parts. Although the OM is not unique, one may find potentials with radii close to the sum of the geometrical radii of the interacting nuclei and diffuseness parameters close to that of the nuclear densities.
In the case of reactions involving weakly bound nuclei, the observed elastic scattering angular distributions display notable deviations with respect to the aforementioned behaviour. The effect becomes more evident in the case of halo nuclei, such as 6He or 11Li. The oscillations in the illuminated region are damped (or even absent). The drop of the cross section with respect to the Rutherford prediction starts at smaller angles and hence the cross section exhibits a smoother angular dependence. An example is given by the 6He+208Pb data taken at $E_{\mathrm{lab}}=22$ MeV shown in the right panel of Fig. 2.
This behaviour can be also described within the OM but with significant modifications with respect to the of standard parametrizations extracted from well-bound nuclei. In particular, one needs very diffuse potentials with long-range absorptive tails. The effect of this long-range absorption is twofold: (i) it damps or suppresses the elastic Coulomb amplitude at small angles (i.e., at very large impact parameters), and (ii) it reduces the nuclear amplitude. The result is that the forward cross-section not only becomes smaller overall, but also loses its Coulomb-nuclear oscillations.
In order to reproduce the elastic scattering data of weakly-bound nuclei within the OM framework using standard Woods-Saxon potentials, a very long-range imaginary part is needed. Table 1 lists the OMP parameters for 4He+208Pb and 6He+208Pb at the same incident energy (22 MeV). The radii of the real and imaginary parts are kept the same for both systems. To reproduce the data, the 6He OM requires a significantly larger imaginary diffuseness parameter. The result of the OM calculation for the 6He is compared with the data for this reaction in the right panel of Fig. 2.
System $V_{0}$ $r_{0}$ $a_{0}$ $W_{v}$ $r_{i}$ $a_{i}$
[MeV] [fm] [fm] [MeV] [fm] [fm]
4He+208Pb 96.44 1.085 0.625 32 0.958 0.42
6He+208Pb 124.8 1.085 0.564 6.8 0.958 1.91
Table 1: Woods-Saxon parameters for 4,6He+208Pb optical models. Reduced radii ($r_{x}$) are converted into absolute (physical) radii as $R_{x}=r_{x}(A^{1/3}_{p}+A^{1/3}_{t})$.
Figure 2: Experimental data for 4,6He+208Pb compared with OM calculations, with Woods-Saxon forms with parameters given in Table 1. Experimental data from Bar74 ; Aco11
This property is a clear indication of the presence of reaction channels dominated by long-range couplings. Natural candidates are the breakup of the projectile and the neutron transfer (either one or two) from the projectile to the target nucleus. Both processes are of a peripheral nature, which explains the long-range imaginary potential. To identify the actual channels producing this effect, one should go beyond the single-channel OM scheme and incorporate those channels into the reaction framework. In the case of the breakup of the projectile, this can be efficiently accomplished with the Continuum-Discretized Coupled-Channels (CDCC) method, which is described in the next sections.
3 Inclusion of breakup: the CDCC method
3.1 Brief resumè of the CDCC method
Figure 3: Left: Application of the CDCC method to $d$+58Ni elastic scattering at $E_{d}=80$ MeV. The solid line is the full CDCC calculation. The dashed line is the calculation omitting the breakup channels. The calculations were performed ignoring the internal spins of the proton and neutron and so $\ell=j$. Right: Illustration of continuum discretization for the same problem.
The CDCC method was originally introduced by G. Rawitscher Raw74 and later refined by the Pittsburgh-Kyushu collaboration Yah86 ; Aus87 to describe the effect of the breakup channels on the elastic scattering of deuterons. Denoting the reaction by $a+A$, with $a=b+x$ (referred hereafter as the core and valence particles, respectively), the method assumes the effective three-body Hamiltonian
$H=H_{\text{proj}}+T_{\vec{R}}+U_{bA}(\vec{r}_{bA})+U_{xA}(\vec{r}_{xA}),$ (5)
with $H_{\mathrm{proj}}=T_{\vec{r}}+V_{bx}$ the projectile internal Hamiltonian, $T_{\vec{r}}$ and $T_{\vec{R}}$ are kinetic energy operators, $V_{bx}$ the inter-cluster interaction and $U_{bA}$ and $U_{xA}$ are the core-target and valence-target optical potentials (complex in general) describing the elastic scattering of the corresponding $b+A$ and $x+A$ sub-systems, at the same energy per nucleon of the incident projectile. In the CDCC method the three-body wave function of the system is expanded in terms of the eigenstates of the Hamiltonian $H_{\mathrm{proj}}$ including both bound and unbound states. Since the latter form a continuum, a procedure of discretization is applied, consisting in representing this continuum by a finite and discrete set of square-integrable functions. In actual calculations, this continuum must be truncated in excitation energy and limited to a finite number of partial waves $\ell$ associated to the relative co-ordinate $\vec{r}$. Normalizable states representing the continuum should be obtained for each $\ell,j$ values. Two main methods are used for this purpose:
  • The pseudo-state method, in which the $b+x$ Hamiltonian is diagonalized in a basis of square-integrable functions, such as Gaussians Kaw86a or transformed harmonic oscillator functions Mor09b . Negative eigenvalues correspond to the bound states of the systems, whereas positive eigenvalues are regarded as a finite representation of the continuum.
  • The binning method, in which normalizable states are obtained by constructing a wave packet (bin) by linear superposition of the actual continuum states over a certain energy interval Aus87 .
We describe the latter method in some more detail. Assuming for simplicity a spinless core, these discretized functions are denoted as
$\phi_{n}(\vec{r})={u_{\ell_{n}j_{n}}(r)\over r}[Y_{\ell}(\hat{r})\otimes\chi_{s}]_{j_{n}m_{j_{n}}},$ (6)
where $n\equiv\{[k_{n},k_{n+1}]\ell_{n},s,j_{n},m_{j_{n}}\}$ specifies the $n$-th bin, with $[k_{n},k_{n+1}]$ the wavenumber interval of the bin, $\ell$ the valence-core orbital angular momentum, $s$ the valence spin, and $\vec{j}=\vec{\ell}+\vec{s}$ the total angular momentum. The symbol $\otimes$ denotes the angular momentum coupling. The radial part of the bin is obtained as a linear combination (i.e., a wave packet) of scattering states as
[…]
where $u_{k,\ell j}(r)$ are the scattering states for a continuum energy $\varepsilon=\hbar^{2}k^{2}/2\mu$ and angular momentum quantum numbers $\ell,s,j$, $w_{n}(k)$ is a weight function (for non-resonant continuum $w_{n}(k)$ is usually taken as $e^{i\delta_{\ell}}$, where $\delta_{\ell}$ are the phase shifts Joa75 of the scattering states within the bin) and $N_{n}$ is a normalization constant. The effect of this averaging is to damp the oscillations at large distances, making the bin wavefunction normalizable.
Assuming a single bound state for simplicity, the CDCC wavefunction can be written
$\Psi^{\mathrm{CDCC}}(\vec{R},\vec{r})=\chi_{0}^{(+)}(\vec{R})\phi_{0}(\vec{r})+\sum_{n=1}^{N}\chi_{n}^{(+)}(\vec{R})\phi_{n}(\vec{r}),$ (8)
where the index $n=0$ denotes the ground state of the $b+x$ system.
This model wave function must verify the Schrödinger equation: $[H-E]\Psi^{\mathrm{CDCC}}(\vec{R},\vec{r})=0$. This gives rise to a set of coupled differential equations for the unknowns $\chi_{n}^{(+)}(\vec{R})$
$\left[E-\varepsilon_{n}-T_{\vec{R}}-U_{nn}(\vec{R})\right]\chi^{(+)}_{n}(\vec{R})=\sum_{m\neq n}U_{nm}(\vec{R})\chi^{(+)}_{m}(\vec{R}),$ (9)
where $\varepsilon_{n}=\langle\phi_{n}|H_{\text{proj}}|\phi_{n}\rangle$ and $U_{nm}(\vec{R})$ are the coupling potentials given by
[…]
The inclusion of the breakup channels in the expansion (8) will affect the elastic channel wavefunction through the coupling potentials $U_{0n}$ ($n>0$). The entrance flux, given by the norm of the plane wave associated with the entrance channel, will be distributed among the elastic and breakup channels and this will naturally reduce the elastic cross section with respect to the situation in which those breakup channels are omitted. This effect turns out to be essential for a correct description of the elastic cross sections, as illustrated in Fig.  3 for the d+58Ni reaction at 80 MeV. The figure depicts also the discretization escheme employed in these calculations, which comprises $\ell=0,2$ continuum states.
In actual calculations, such as those performed by the popular coupled-channels code FRESCO fresco , the total three-body wavefunction is expanded in states with good total angular momentum ($J_{T}$), i.e.,
$\Psi({\vec{R}},{\vec{r}})=\sum_{\beta_{i},J_{T},M_{T}}C_{\beta_{i},J_{T},M_{T}}\Psi_{\beta_{i},J_{T},M_{T}}(\vec{R},\vec{r})$ (11)
with
$\Psi_{\beta_{i},J_{T},M_{T}}(\vec{R},\vec{r})=\sum_{\beta}\frac{\chi_{\beta,\beta_{i}}^{J_{T}}(R)}{R}\left[Y_{L}({\hat{R}})\otimes\phi_{n,J}(\vec{r})\right]_{J_{T},M_{T}},$ (12)
where, for the internal states, $\phi_{n,J,M}$, we have included an additional subscript $J$ to write explicitly the projectile spin and hence $n$ is now an index to enumerate states with the same $J$. Also, in the former equation $\vec{R}$ is the relative coordinate between the projectile center of mass and the target (assumed by now to be structureless), see Fig. 4. The index $\beta$ gathers the quantum numbers compatible with a given total angular momentum $J_{T}$, $\beta\equiv\{L,J,n\}$, where $L$ (projectile-target orbital angular momentum) and $J$ both couple to the total spin of the three-body system $J_{T}$. The spin of the target is ignored for simplicity of the notation. Each of these sets is called channel. In particular, $\beta_{i}$ denote the channels compatible with the initial state of the system (typically, the ground state of the projectile and target nuclei).
Figure 4: Relevant coordinates for the description of the scattering of a two-body composite nucleus by a target.
The radial coefficients, $\chi_{\beta,\beta_{i}}^{J_{T}}(R)$, from which the scattering observables are extracted, are calculated by inserting (12) in the Schrödinger equation, giving rise to a system of coupled differential equations:
$\left(-{\hbar^{2}\over 2\mu}{d^{2}\over dR^{2}}+{\hbar^{2}L(L+1)\over 2\mu R^{2}}+\varepsilon_{n}-E\right)\chi^{J_{T}}_{\beta,\beta_{i}}(R)$
$+\sum_{\beta^{\prime}}U_{\beta,\beta^{\prime}}^{J_{T}}(R)\chi^{J_{T}}_{\beta^{\prime},\beta_{i}}(R)$ $=0$ (13)
with the coupling potentials:
$U_{\beta,\beta^{\prime}}^{J_{T}}(R)=\langle\beta;J_{T}|V_{bA}(\vec{R},\vec{r})+V_{xA}(\vec{R},\vec{r})|\beta^{\prime};J_{T}\rangle,$ (14)
where
$\langle\hat{R},\vec{r}|\beta;J_{T}\rangle=\left[Y_{L}({\hat{R}})\otimes\phi_{n,J}(\vec{r})\right]_{J_{T}}.$ (15)
The system of equations (13) is to be solved numerically following the methods described elsewhere (see, e.g. Hag22 and references therein) and subject to the asymptotic boundary conditions:
$\chi^{J_{T}}_{\beta:\beta_{i}}(K_{\beta},R)$ $\rightarrow\frac{i}{2}\left[H^{(-)}_{L}(K_{\beta}R)\delta_{\beta,\beta_{i}}-S^{J_{T}}_{\beta,\beta_{i}}H^{(+)}_{L}(K_{\beta}R)\right]$ (16)
$\rightarrow\left[F_{L}(K_{\beta}R)\delta_{\beta,\beta_{i}}+T^{J_{T}}_{\beta,\beta_{i}}H^{(+)}_{L}(K_{\beta}R)\right]$ (17)
where $K_{\beta}=\sqrt{2\mu(E-\epsilon_{n})}$, $S^{J_{T}}_{\beta,\beta_{i}}$ are the S-matrix elements, $T^{J_{T}}_{\beta,\beta_{i}}$ are the T-matrix elements, $H^{\pm)}(KR)$ are the ingoing (--) and outgoing ($+$) Coulomb functions and $F_{L}(KR)$ is the regular Coulomb function Sat83 . Because of the relation between the Coulomb functions, the two equations above are fully equivalent. In fact, the coefficients $S^{J_{T}}_{\beta,\beta_{i}}$ and $T^{J_{T}}_{\beta,\beta_{i}}$ are related by $S^{J_{T}}_{\beta,\beta_{i}}=\delta_{\beta,\beta_{i}}+2iT^{J_{T}}_{\beta,\beta_{i}}$. Scattering amplitudes and differential cross sections are expressed in terms of these coefficients (see, e.g., Glen04 ; Sat83 ).
Figure 5: Application of the CDCC method to 6Li+40Ca elastic scattering at 156 MeV. Experimental data are from Ref. Maj78 . Adapted from Mor09b .
The standard CDCC method is based on a strict three-body model of the reaction ($b+x+A$), and has proven rather successful to describe elastic and breakup cross sections of deuterons and other weakly bound two-body nuclei, such as 6,7Li and 11Be (see Fig. 5). However, it has limitations. The assumption of inert bodies is not always justified, since excitations of the projectile constituents ($b$ and $x$) and of the target ($A$) may take place along with the projectile dissociation. Furthermore, the two-body picture may be inadequate for some nuclei as, for example, in the case of the Borromean systems (e.g. 6He, 11Li). Some extensions of the CDCC method to deal with the these situations are outlined below.
3.2 The role of closed channels
In principle, the expansion (8) may include internal states with excitation energies below ($\epsilon_{i}<E$) or above ($\epsilon_{i}>E$) the total energy, which are referred to as open and closed channels, respectively. Closed channels violate energy conservation and, as such, cannot contribute asympotically to the scattering wavefunction. However, due to the couplings with other channels, they can affect the open channels and hence the cross sections.
The solution of the CDCC equations (13) in the presence of closed channels is totally analogous to the case with open channels, except for the modification of the boundary conditions of the radial functions associated with closed channels, which now reads
$\chi^{J_{T}}_{\beta:\beta_{i}}(K_{\beta},R)$ $\rightarrow C_{\beta,\beta_{i}}W_{-\eta_{i},L+1/2}(-2iK_{\beta}R),$ (18)
where $W_{-\eta_{i},L+1/2}$ is a Whittaker function with $\eta$ the Sommerfeld parameter, defined as $\eta=Z_{p}Z_{t}e^{2}/\hbar v$. In general, the effect of closed channels is small at sufficiently high energies. However, at low energies, they can have a strong influence on the scattering observables. An example is shown in Fig. 6, corresponding to the reaction $d+^{12}\mathrm{C}\rightarrow p+n+^{12}\mathrm{C}$ at 12 MeV Oga16 . CDCC calculations with and without the inclusion of closed channels are compared with Faddeev calculations, which account also for the effect of closed channels. The upper and lower panels are for the angular and excitation energy distributions. The importance of closed channels is clearly seen.
Figure 6: (a) Angular distribution and (b) breakup energy distribution of the elastic breakup cross section for 12C($d,pn$)12C at 12 MeV. The solid, dashed, and dash-dotted lines in each panel show the converged CDCC result, the result of the CDCC method calculated with only the open channels, and the result of FAGS theory taken from Ref. Upa12 , respectively. The dotted line in (a) is the same as the solid line but omitting the odd partial waves between $p$ and $n$. Taken from Oga16 , with authorization from APS.
3.3 Inclusion of core excitations
Excitations of the projectile constituents ($b$ and $x$ in our case) may take place along with the projectile breakup. This mechanism is neglected in the standard formulation of the CDCC method. For example, for the scattering of halo nuclei, the core fragment $b$ is assumed to be inert and, as such, the projectile states are described by pure single-particle or pure cluster states. Except for very specific cases (e.g. when $b$ is an alpha particle), the structure of a composite system of the form ($a=b+x$) is known to contain in general admixtures of core-excited components. Additionally, the interaction of the core with the target will produce excitations and deexcitations of the former during the collision, thereby modifying the reaction observables to some extent. These two effects (structural and dynamical) have been investigated within extended versions of the DWBA and CDCC methods Cre11 ; Mor12 ; Sum06 ; Die14 . For example, considering only possible excitations of $b$, the effective three-body Hamiltonian is generalized as follows:
$H=H_{\rm proj}(\vec{r},\xi_{b})+T_{\vec{R}}+U_{bA}(\vec{r}_{bA},\xi_{b})+U_{xA}(\vec{r}_{xA}).$ (19)
The potential $U_{bA}(\vec{r}_{bA},\xi_{b})$ is now meant to describe both elastic and inelastic scattering of the $b+A$ system (for example, it could be represented by a deformed potential as those used in the context of inelastic scattering with collective models). Note that the core degrees of freedom ($\xi_{b}$) appear in the projectile Hamiltonian (structure effect) as well as in the core-target interaction (dynamic effect).
In the weak-coupling limit, the projectile Hamiltonian can be written more explicitly as
$H_{\rm proj}=T_{\vec{r}}+V_{bx}(\vec{r},\xi_{b})+h_{\rm core}(\xi_{b}),$ (20)
where $h_{\rm core}(\xi_{b})$ is the internal Hamiltonian of the core. The eigenstates of this Hamiltonian are of the form
$\phi_{n,J,M}(\xi_{b},\vec{r})\equiv\sum_{\alpha}\left[\varphi^{n}_{\alpha}(\vec{r})\otimes\Phi_{I}(\xi_{b})\right]_{JM},$ (21)
where $n$ is an index labeling the states with angular momentum $J$ (with projection $M$), $\alpha\equiv\{\ell,s,j,I\}$, with $I$ the core intrinsic spin, $\vec{j}=\vec{\ell}+\vec{s}$ and $\vec{j}_{p}=\vec{j}+\vec{I}$. The functions $\Phi_{I}(\xi_{b})$ and $\varphi_{\alpha}(\vec{r})$ describe, respectively, the core states and the valence–core relative motion. This expansion is in principle also valid for continuum states, in which case a procedure of continuum discretization must be used.
Once the projectile states (21) have been calculated, the three-body wave function is expanded in a basis of such states, as in the standard CDCC method (c.f. Eq. (12)),
$\Psi_{\beta_{i},J_{T},M_{T}}(\vec{R},\vec{r},\xi)=\sum_{\beta}\chi_{\beta,\beta_{i}}^{J_{T}}(R)\left[Y_{L}({\hat{R}})\otimes\phi^{(N)}_{n,J}(\vec{r},\xi)\right]_{J_{T},M_{T}},$ (22)
where the superscript $N$ is the number of internal states retained in the unavoidable basis truncation.
Early calculations using this extended CDCC method (referred to as XCDCC) were first performed by Summers et al. Sum06 ; Sum07 for 11Be and 17C on 9Be and 11Be+$p$, finding a very little core excitation effect in all these cases. However, later calculations for the 11Be+$p$ reaction based on a alternative implementation of the XCDCC method using a pseudo-state representation of the projectile states Die14 suggested much larger effects. The discrepancy was found to be due to an inconsistency in the numerical implementation of the XCDCC formalism presented in Ref. Sum06 , as clarified in Sum14 . For heavier targets, such as 64Zn or 208Pb, the calculations of Die14 suggest that the core excitation mechanism plays a minor role, although its effect on the structure of the projectile is still important.
Figure 7: Breakup angular distributions for the relative-energy cuts indicated in the labels. Adapted from Ref. Die14 ; Die17 .
[…]
Figure 8: Elastic (left), inelastic (middle) and breakup (right) cross sections for the reaction 11Be+197Au at near barrier energies ($V_{b}\sim 40$ MeV). Experimental data are compared with CDCC, XCDCC and first-order semiclassical pure $E1$ calculations (labeled EPM). Adapted from Ref. Pes17 .
The importance of the deformation on the structure of the projectile is clearly evidenced in the elastic and inelastic scattering of 11Be on 197Au at near-barrier energies Pes17 , as shown in Fig. 8. XCDCC calculations based on the particle-plus-rotor model of Ref. Tar03 (solid lines) are able to reproduce simultaneously the elastic, inelastic and breakup angular distributions. By contrast, standard CDCC calculations using single-particle wave functions fail to describe the elastic and inelastic data, despite describing well the breakup. This is due to the overestimation of the $B(E1)$ connecting the ground state with the bound excited state, as discussed in Pes17 .
A simpler DWBA, no-recoil version of the formalism (XDWBA) has also been proposed in refs. Cre11 ; Mor12 . An application of this formalism to the 11Be+12C reaction at 69 MeV/u showed that the core excitation mechanism may interfere with the single-particle excitation mechanism, producing a conspicuous effect on the interference pattern of the resonant breakup angular distributions, as illustrated in Fig. 9 for the population of the low-lying $5/2^{+}$ and $3/2^{+}$ resonances of 11Be Mor12b .
Figure 9: Valence (dashed line) and core (dot- dashed line) contributions to the breakup of the 1.78 and 3.41 MeV resonances populated in the 11Be+12C reaction at 70 MeV/nucleon, using a particle-core description of the 11Be nucleus. The solid line is the coherent sum of both contribution. An overall normalization factor is included in both plots, to match the magnitude of the data. Adapted from Ref. Mor12b , with authorization from APS.
3.4 Inclusion of target excitations
In addition to the excitations of the projectile constituents, excitations of the target nucleus may also take place and compete with the projectile breakup mechanism. Note that, within CDCC, the projectile breakup is treated as a inelastic excitation of the projectile to its continuum states and, thus, inclusion of target excitation amounts at including, simultaneously, projectile plus target excitations so their relative importance, and mutual influence, can be assessed. These target excitations can be treated with the collective models mentioned in the previous section. It is worth noting that, within this three-body reaction model, target excitation arises from the non-central part of the valence-target and core-target interactions. To incorporate this effect, the effective Hamiltonian, eq. (5), must be now generalized as:
$H=H_{\text{proj}}(\vec{r})+T_{\vec{R}}+U_{bA}(\vec{r}_{bA},\xi_{t})+U_{xA}(\vec{r}_{xA},\xi_{t})+H_{A}(\xi_{t})$ (23)
where $H_{A}(\xi_{t})$ is the target internal Hamiltonian, and the $b-A$ and $x-A$ interactions depend now, in addition to the corresponding relative coordinate, on the target degrees of freedom (denoted as $\xi_{t}$). Ideally, these $U_{xA}$ and $U_{bA}$ potentials should reproduce simultaneously the elastic and inelastic scattering for the $x+A$ and $b+A$ reactions, respectively. Once a specific model has been chosen for the target, its eigenstates ($\Phi^{i}_{J_{t}}(\xi_{t})$) can be computed. They are characterized by their angular momentum $J_{t}$ and the index $i$.
In the presence of these target excitations, the CDCC wavefunction is now expanded on the generalized channel basis Mor12b ; Die14 ; fresco containing the target eigenstates
$\Psi^{\text{CDCC}}_{J_{T},M_{T}}(\vec{R},\vec{r},\xi)$ $=\sum_{\beta}\chi_{\beta}^{J_{T}}(R)\left\{\left[Y_{L}(\hat{R})\otimes\phi^{n}_{J_{p}}(\vec{r})\right]_{J}\otimes\Phi^{i}_{J_{t}}(\xi_{t})\right\}_{J_{T},M_{T}},$ (24)
where $\beta$ denotes all the quantum numbers necessary to define the channel $\beta=\{L,n,J_{p},J,J_{t},i\}$. The explicit inclusion of target excitation was first done by the Kyushu group in the 1980s Yah86 , which considered the case of deuteron scattering. The motivation was to compare the roles of target excitation and deuteron breakup in the elastic and inelastic scattering of deuterons. They applied the formalism to the $d$+58Ni reaction at $E_{d}=22$ and 80 MeV, including the ground state and the first excited state of 58Ni ($2^{+}$) and finding that, in this case, the deuteron breakup process is more important than the target-excitation.
Recently, the problem has been also addressed by some authors Chau11 ; Gom17a , also in the context of deuteron elastic and inelastic scattering. A recent application of the formalism is shown in Fig. 10, which corresponds to the reaction ${}^{24}{\rm Mg}(d,d)^{24}{\rm Mg}^{*}$ at $E_{d}=70$ MeV, including the ground and first excited states of ${}^{24}{\rm Mg}$, in addition to the deuteron breakup. The data are from Ref. Kis76 . The target excitation was treated within the collective model, using a quadrupole deformation parameter of $\beta_{2}=0.5$. Also included are Faddeev calculations performed by A. Deltuva Del16 . Both calculations reproduce equally well the elastic differential cross section. The calculated inelastic angular distributions are slightly out of phase with the data, but they agree well with each other, pointing to some inadequacy of the structure or potential inputs. Although these inelastic cross section can be also well reproduced with standard DWBA calculations based on a deformed deuteron-target potential, it was shown in Del16 that the extracted deformation parameter obtained with the three-body approach is more consistent with that derived from nucleon-nucleus inelastic scattering.
Figure 10: Elastic (upper) and $(2^{+})$ excitation (lower) differential angular cross sections for $d$+24Mg $E_{d}=70$ ${\rm MeV}$. CDCC calculations are compared with Faddeev calculations from ref. Del16 and with the data of ref. Kis76 . The calculations employ CH89 CH89 parameterization for the $p$+24Mg and $n$+24Mg potentials and treat the target excitation with a collective model, assuming a deformation of $\beta=0.5$ for the 24Mg nucleus. The plot is adapted from ref. Gom17a .
3.5 Four-body CDCC
The extension of the CDCC formalism to include three-body projectiles is typically referred to as four-body CDCC and was developed in Refs. Mat04a ; Mat04b ; manoli08 ; manoli09 to describe reactions induced by the Borromean two-neutron halo nuclei 6He and 11Li. Within the few-body scheme, the method amounts to incorporate the three-body structure in the corresponding coupling potentials of the CDCC equations, thus solving the same coupled-channel problem but typically with a considerably larger model space.
Figure 11: Schematic representation of the four-body CDCC, where the projectile is described in Jacobi coordinates $\{\vec{x},\vec{y}\}$. Adapted from Ref. casal15 .
In simple terms, the generalization implies the inclusion of three-body projectile wave functions in the channel states of Eq. (12). These can be written in Jacobi coordinates $\left\{\vec{x},\vec{y}\right\}$ (see Fig. 11), in such a way that the coupled basis (15) becomes
$\langle\hat{R},\vec{x},\vec{y}|\beta;J_{T}\rangle=\left[Y_{L}({\hat{R}})\otimes\Phi_{n,J}(\vec{x},\vec{y})\right]_{J_{T}},$
and the coupling potentials, now involving the interaction between each of the three bodies and the target, reads schematically as
$U_{\beta,\beta^{\prime}}^{J_{T}}(R)=\langle\beta;J_{T}|V_{1t}+V_{2t}+V_{3t}|\beta^{\prime};J_{T}\rangle_{\vec{x},\vec{y}}.$ (25)
As discussed, for instance, in Ref. manoli08 ; casal15 , the fragment-target potentials can be expressed each in a different $q$-Jacobi system, so that its radial dependence involves only $\vec{R}$ and $\vec{y}_{q}$. Then, a multipole expansion of these potentials is introduced to obtain the corresponding form factors. The effect of different multipole couplings can therefore be studied.
The construction of the three-body projectile states can be performed using various techniques. In Refs. Mat06 ; Matsumoto19 , for instance, a Gaussian expansion method in Jacobi coordinates is employed. Here we outline another typical choice, based on the standard hyperspherical harmonics formalism Zhu93 . By introducing the hyperspherical coordinates $\rho=\sqrt{x^{2}+y^{2}}$ and $\alpha=\arctan\left(x/y\right)$, one can write
$\Phi_{n,J,M}(\rho,\alpha,\hat{x},\hat{y})=\frac{1}{\rho^{5/2}}\sum_{\eta}u_{n\eta}(\rho)\mathcal{Y}_{\eta JM}(\alpha,\hat{x},\hat{y}),$ (26)
with
$\mathcal{Y}_{\eta JM}(\alpha,\hat{x},\hat{y})=\varphi_{K}^{l_{x},l_{y}}(\alpha)\left\{\left(\left[Y_{l_{x}}(\hat{x})\otimes Y_{l_{y}}(\hat{y})\right]_{l}\otimes\kappa_{S_{x}}\right)_{j}\otimes\chi_{I}\right\}_{JM}.$ (27)
Here, $\eta=\left\{K,l_{x},l_{y},l,S_{x},j,I\right\}$ are the possible sets of quantum numbers associated to the orbital angular momenta and spins coupled to a total $J$, where $K$ is the so-called hypermomentum, and $\varphi_{K}^{l_{x},l_{y}}(\alpha)$ are obtained in terms of Jacobi polynomials. The three-body hyperradial functions $u_{n\eta}(\rho)$ can be obtained within different discretization methods, including the binning procedure manoli09 ; Cub12 and the pseudostate method with different bases manoli08 ; Desc15 ; casal15 . This is done for a fixed value of $K_{max}$, up to convergence. Note that the above wave-function expansion and its coupling order is not unique, and other choices exist in the literature.
Figure 12 shows an application of the four-body CDCC method to describe the ${}^{11}\text{Li}+{{}^{208}}\text{Pb}$ reaction at near-barrier energies Cub12 , where the projectile states and the corresponding coupling potentials are computed with a ${}^{9}\text{Li}+n+n$ model using the binning discretization method. In that work, it is shown that the coupling to low-energy continuum states is crucial to explain the rapid decrease of the elastic differential cross section, and the calculations indicate the presence of a low-lying dipole resonance in 11Li for a proper description of the experimental data.
Figure 12: Experimental ${}^{11}\text{Li}+{{}^{208}}\text{Pb}$ elastic cross section at $E_{lab}=29.8$ MeV, relative to Rutherford, compared to four-body CDCC calculations (solid) and the result without the coupling to continuum states (dashed). Adapted from Ref. Cub12 .
Figure 13 presents results for the ${}^{11}\text{Li}(p,p^{\prime})$ reaction, focusing on the breakup cross section extracted from the four-body CDCC Matsumoto19 . These calculations employ pseudostates to describe the 11Li states, and the so-called complex scaling method is adopted to build the corresponding energy distributons. Again, a dipole resonance is found to be crucial for a proper understanding of the data. Other calculations for the same reaction exist in the literature Descouvemont20 .
Figure 13: (a) Angular distribution and (b) breakup-energy distribution for ${}^{11}\text{Li}(p,p^{\prime})$. The calculation labeled ‘1 step’ represents the result without continuum states, and the solid black line is the total calculation including $0^{+},1^{-},2^{+}$ states. Taken from Ref. Matsumoto19 . Data are from Ref. tanaka17 .
Four-body CDCC calculations typically involve very large channel sets, for which the pseudostate method is computationally less demanding than the binning method. However, it was shown in e.g., Ref. mrg11proc that the results for weakly bound projectiles on heavy targets may exhibit a slow convergence when using pseudostates, due to the very strong continuum couplings involving low-energy states (see Sec. 3.7.3 below). Still, pseudostates are the preferred choice in many cases, especially for three-body projectiles comprising several charged particles, for which the calculation of the true three-body continuum states (and the corresponding bins) is not feasible. In Refs. Desc15 ; casal15 , four-body CDCC calculations were performed for the scattering of ${}^{9}\text{Be}(\alpha+\alpha+n)$, showing reasonable agreement with the available experimental data. Figure 14 shows some ${}^{9}\text{Be}+{{}^{208}}\text{Pb}$ results from Ref. casal15 , where the relevance of dipole couplings for a proper description of the observations was highlighted.
Figure 14: Experimental ${}^{9}\text{Be}+{{}^{208}}\text{Pb}$ elastic cross section at $E_{\mathrm{lab}}=44$ MeV, relative to Rutherford, compared to four-body CDCC calculations (solid black) and the result without the coupling to continuum states (dashed magenta). Adapted from Ref. casal15 . Data are from Refs. Wooll2004 ; Yu2010 .
3.6 Microscopic CDCC
The standard CDCC method assumes a cluster (two-body or three-body) description of the projectile nucleus. This simplification has of course limitations and drawbacks. For example: (i) it requires cluster-target optical potentials, which are are not always well determined; (ii) the extension to more than three bodies is very challenging and currently not available; (iii) excitations of the fragments are ignored altogether or, at most, approximately included with some collective model. To overcome these problems, a microscopic version of the CDCC method (MCDCC) has been proposed by Descouvemont and co-workers Des13 ; Des18 . The method uses a many-body description of the projectile states, based on a cluster approximation, known as resonating group method (RGM). In the RGM, an eigenstate of the projectile Hamiltonian is written as an antisymmetric product of cluster wave functions. For example, for a 7Li projectile, described as $\alpha+t$, the RGM wave function is expressed as:
$\Phi_{i,J,M}(\xi_{p})$ $={\mathcal{A}}\bigl{[}[\phi_{\alpha}\otimes\phi_{t}]^{1/2}\otimes Y_{\ell}(\Omega_{\rho})\bigr{]}^{J,M}g^{\ell J}_{i}(\rho),$ (28)
where $\phi_{\alpha}$ and $\phi_{\alpha}$ are shell model wave functions of the $\alpha$ and $t$ clusters, $\ell$ their relative orbital angular momentum and $J$ the total spin. In Eq. (28), $\vec{\rho}$ is the relative coordinate (see Fig. 15), and ${\mathcal{A}}$ is the 7-body antisymmetrization operator which takes into account the Pauli principle among the 7 nucleons of the projectile. The function $g^{\ell J}_{i}(\rho)$ is determined from a Schrödinger equation associated with the projectile Hamiltonian. Continuum states are included using a pseudo-state basis.
In MCDCC, the projectile-target interaction is given by the sum of nucleon-nucleus interactions (instead of cluster-target interactions), for which reliable parametrizations are available. Thus, the full projectile+target Hamiltonian is written as:
$H=H_{0}+T_{\vec{R}}+\sum_{j=1}^{A}U_{tj}({\vec{r}}_{j}-{\vec{R}}),$ (29)
where $U_{tj}({\vec{r}}_{j}-{\vec{R}})$ is the interaction of the $j$-th projectile constituent with the target.
Figure 15 shows an application of the method to the reaction 7Li+208Pb at near-barrier energies. Experimental data are compared with a one-channel calculation (only 7Li g.s.), two-channels calculations (ground plus first excited state) and several CDCC calculations including continuum states up to a certain projectile angular momentum. It is seen that a good description of the data is achieved when a sufficiently large number of continuum states is included.
Figure 15: Top: Schematic picture of the projectile-target system, with a microscopic cluster structure of the projectile. Bottom: 7Li+208Pb elastic cross section Mar96 , relative to Rutherford, compared with microscopic CDCC calculations. Dotted lines represent the calculations without breakup channels and the solid lines are the full calculations with increasing $\alpha+t$ maximum angular momentum of the projectile. Taken from Ref. Des13 , with authorization from APS.
Other applications of the microscopic CDCC method include its extension for three-body projectiles including a single nucleon as one of the three particles, such as 8Li or 8B as $(^{4}\text{He}+t+n)$ and $(^{4}\text{He}+{{}^{3}}\text{He}+p)$, respectively Descouvemont2018 . Figure 16 shows results for the elastic scattering of 7Li on 12C and 209Bi targets at low energies. The dashed line corresponds to the calculation including only the projectile ground state, while the solid line are the full microscopic CDCC calculations that exhibit a better agreement with the data.
Figure 16: 8Li+12C (top panel) and 8Li+209Bi (bottom panel) elastic cross section, relative to Rutherford, computed within the microscopic CDCC approach for three-body projectiles. Dotted lines represent the calculations without breakup channels and the solid lines are the full calculations. Experimental data from Refs. Coo18 ; Barioni09 . Figure from Ref. Descouvemont2018 , with authorization from Springer.
3.7 Interpretation of the CDCC wavefunction and cross sections
For a meaningful comparison with the experimental data, it is important to properly understand which information and observables are (and which are not!) provided by the CDCC wavefunction. Of course, elastic scattering is the most direct outcome of the method and the one that can be compared with the data more easily. By construction, the elastic scattering produced by the CDCC calculation incorporates the effect of the coupling to the inelastic channels and the breakup channels included in the CDCC modelspace. According to our former notation, these breakup channels correspond to the three-body final states of the form $b+x+A$ in which the target and the projectile subsystems $b$ and $x$ escape the interaction region and remain in their ground state. Morover, in addition to these channels explicitly included in the modelspace, the CDCC includes effectively the effects of other channels associated with possible excitations of the target and/or of the fragments via the optical potentials.
Regarding breakup observables, in principle, the CDCC wavefunction provides only the elastic breakup ones. The most immediate observables are the breakup angular distributions for specific bin intervals, i.e.,
$\frac{d\sigma_{n}}{d\Omega_{\text{c.m.}}}=|f_{0,n}(\theta)|^{2}.$ (30)
When the binning method is employed, a double differential cross section (with respect to the c.m. angle and relative energy) can be obtained by dividing each angular distribution by the bin width, as follows:
$\frac{d^{2}\sigma}{d\Omega_{\text{c.m.}}d\epsilon}\simeq\frac{1}{\Delta_{n}}\frac{d\sigma_{n}}{d\Omega_{\text{c.m.}}}.$ (31)
Regarding the nonelastic breakup contributions, which are effectively included in the fragment-target imaginary potentials, the CDCC method by itself does not provide detailed breakup cross sections involving excitations. However, one can define and compute a total absorption cross section that accounts for their overall contribution. For that, one first derives a reaction cross section in terms of the CDCC S-matrices asnve [^1] See Ref. Des17 for a different convention.:
$\sigma_{R}=\frac{\pi}{k^{2}_{i}}\sum_{J_{T}\pi}\sum_{L_{i}}\frac{2J_{T}+1}{(2J^{i}_{p}+1)(2J^{i}_{t}+1)}\left(1-|S^{J_{T}}_{\beta_{i},\beta_{i}}|^{2}\right),$
and an integrated elastic breakup cross section
$\sigma_{\text{bu}}=\frac{\pi}{k^{2}_{i}}\frac{k}{k_{i}}\sum_{J_{T}\pi}\sum_{LL_{i}}\frac{2J_{T}+1}{(2J^{i}_{p}+1)(2J^{i}_{t}+1)}|S^{J_{T}}_{\beta,\beta_{i}}|^{2}.$
Then, one may define an absorption cross section as $\sigma_{\text{abs}}=\sigma_{\text{R}}-\sigma_{\text{bu}}$. This absorption cross section will account for all nonelastic processes associated with the absorptive part of the fragment-target potentials, such as target excitation (i.e., inelastic scattering), complete fusion (i.e., capture of the whole projectile) or incomplete fusion (i.e., capture of one of the projectile fragments). Very often, comparison with experimental data requires more detailed cross sections, such as the angular or energy distributions of the fragment that survives after the capture (absorption) of the other fragment. As noted above, these observables are not directly provided by the CDCC method, at least in its standard formulation. However, as will be shown in Sect. 3.7.2, the CDCC wavefunction can be used as an input for a suitable formalism capable of providing such observables.
The absorption cross section plays an important role in understanding the CDCC results. To quantify the importance of the continuum on the elastic cross sections it is common to compare full CDCC results with no-continuum calculations, in which only the ground-state to ground-state coupling potential is retained. An example is shown in Fig. 17, which depicts the results for the 6He+208Pb reaction at 18 MeV and the 8B+90Zr reaction at 26.5 MeV. Dashed and solid lines denote the no-continuum and full CDCC calculations, respectively (in the 6He case, the improved dineutron model of Ref. Mor07 was used). Table 2 lists the reaction, absorption, and (elastic) breakup cross sections from the CDCC calculations. Clearly, inclusion of the continuum has a much larger effect on the elastic cross section for 6He. This is in apparent contraction with the fact that the elastic breakup cross section is larger in the 8B case, as shown in the Table. The apparent inconsistency can be explained by looking at the absorption and reaction cross sections. In the 8B reaction, the inclusion of the continuum couplings produces only a small increase of the absorption cross section, and the reaction cross section is essentially increased by the elastic breakup cross section. By contrast, in the 6He reaction, the absorption cross section increases by almost a factor of 3 when adding the continuum couplings and this results also in an equivalent increase of the reaction cross section, hence the enhanced effect on the elastic cross section. One may anticipate that the increased absorption cross section for 6He is associated with processes in which the valence neutrons experience some kind of nonelastic interaction with the target, such as transfer to bound states.
Reaction $\sigma_{\mathrm{reac}}$ $\sigma_{\mathrm{abs}}$ $\sigma_{\mathrm{bu}}$
(mb) (mb) (mb)
8B+90Zr Full 362 136 226
no cont. 136 124
6He+208Pb Full 477 359 117
no cont. 134 134
Table 2: Comparison of reaction, absorption and elastic breakup cross sections for 6He+208Pb at 18 MeV and 8B+90Zr at 26.5 MeV.
Figure 17: Comparison of one-channel (i.e., no continuum) calculations and full CDCC calculations for (a) 8B+90Zr and (b) 6He+208Pb at near-barrier energies. See text for details.
The different effect of the breakup channels on the elastic channel for neutron versus proton halo nuclei has been addressed in the literature by several authors Bon04 ; Kum12 ; Yan16 .
3.7.1 Two-body observables
The solution of the coupled equations provides the $S-$matrix elements connecting the ground state with the continuum states, from which breakup observables can be constructed. In principle, these breakup $S-$matrix elements corresponding to a given set of quantum numbers would be a continuous function of the excitation energy (or momentum). However, due to the discretization procedure, the S-matrices calculated in CDCC correspond to discrete values of the energy. In the binning method, an approximation to this continuous $S-$matrix can be obtained dividing the discrete $S-$matrix by the square root of the bin width, i.e.,
$S^{J_{T}}_{\beta,\beta_{i}}(k)\approx\frac{1}{\sqrt{\Delta E_{n}}}\hat{S}^{J_{T}}_{\beta,\beta_{i}}$ (32)
with $\beta=\{L,J,n\}$, $\beta_{i}=\{L_{i},J_{0},n_{0}\}$ and $\Delta E_{n}$ is the width of the $n$-th bin.
In the PS method, one could apply a similar procedure by assigning a width to each pseudostate. In fact, this approach was used, for example, in Ref. Mor02 to calculate the differential breakup cross section from the cross section to individual pseudostates, assuming that the width of the $i$th pseudostate is approximately given by $\Delta E_{n}=(\varepsilon_{n+1}-\varepsilon_{n-1})/2$. A more accurate procedure, previously proposed in Ref. Mat03 is to obtain continuous $S-$matrix elements $S_{\beta,\beta_{i}}(k)$, depending on the continuous variable $k$, as well as on the initial and final angular momenta, by an appropriate superposition of the discrete $S-$matrix elements $\hat{S}_{\beta,\beta_{i}}$ resulting from the solution of the coupled channel equations. Ignoring for simplicity intrinsic spins of the projectile constituents as Mat03 ; Tos01 one gets
$S^{J_{T}}_{\beta,\beta_{i}}(k)\approx\sum_{n=1}^{N}\langle\phi^{(-)}_{k,\ell}|\phi^{(N)}_{n,\ell}\rangle\hat{S}^{J_{T}}_{\beta,\beta_{i}},$ (33)
where $\phi^{(-)}_{k,\ell}(r)$ and $\phi^{(N)}_{n,\ell}(r)$ are the radial parts of the exact and pseudostate wavefunctions, respectively. The sum runs over the set pseudostates included in the coupled–channels calculation.
[…]
Figure 18: (Color online) Modulus of the breakup $S$ matrix elements for the total angular momentum $J=17$ for the reaction $d$+58Ni at 80 MeV as a function of the p–n relative momentum in the final state. The histogram is the CDCC calculation with binning method. The filled and open circles represent the CDCC calculations using the analytical THO pseudostate basis of Ref. Mor09b . The solid lines are obtained folding the discrete $S$-matrices with the true continuum wavefunctions. Adapted from Ref. Mor09b .
3.7.2 Three-body observables
Within the CDCC and XCDCC reaction formalisms, breakup is treated as an excitation of the projectile to the continuum, so the theoretical cross sections are described in terms of the c.m. scattering angle of the projectile and the relative energy of the constituents, using two-body kinematics. For comparison with experimental data, it is useful to have also the cross sections in terms of the angle and energy of the projectile fragments, since these quantities are more directly connected with the actual measurements.
In the case of the standard CDCC framework, fivefold fully exclusive cross sections have been derived and presented by several authors Ise86 ; Tos01 . The method was generalized in Ref. Die17 to the case of XCDCC. We briefly review the main formulas of the latter, noting that the case without core excitation is recovered when a single core state is considered. For simplicity, we ignore the target spin.
In XCDCC, the solution of the coupled-channel equations provides some discrete breakup transition amplitudes, denoted $T_{M_{0},M^{\prime}}^{i,J_{0},J^{\prime}}(\theta_{i},K_{i})$, connecting an initial state $|J_{0}M_{0}\rangle$ with a three-body final state comprised by the target (assumed to be structureless), the valence particle and the core, and some discrete value of the final projectile-target c.m. momentum $\vec{K}_{i}=\{\theta_{i},K_{i}\}$. The first step of the formalism is to relate these discrete amplitudes with the actual breakup scattering amplitudes, without continuum discretization, that we denote as $T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})$, where $\vec{k}$ is the projectile internal relative momentum. Formally, these breakup transition amplitudes can be written in integral form as:
$T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})=\langle\phi_{\vec{k};I\mu;s\sigma}^{(-)}e^{i\vec{K}\cdot\vec{R}}|U|\Psi_{J_{0},M_{0}}(\vec{K}_{0})\rangle,$ (34)
where $U=U_{bA}(\vec{r},\vec{R},\xi)+U_{xA}(\vec{r},\vec{R})$ and $\phi_{\vec{k};I\mu;s\sigma}^{(-)}$ are two-body exact scattering wave functions of the $b+x$ system for a relative final momentum $\vec{k}$ and given core and valence spins.
In order to relate the discrete and continuous amplitudes, one can approximate the exact wavefunction $\Psi_{J_{0},M_{0}}$ in the equation above by its XCDCC counterpart and introduce the approximate completeness relation in the truncated discrete basis $\{\phi^{(N)}_{n,J,M};i=1,\ldots,N\}$:
$T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})$ $\simeq\sum_{n,J^{\prime},M^{\prime}}\langle\phi_{\vec{k};I\mu;s\sigma}^{(-)}|\phi^{(N)}_{n,J^{\prime},M^{\prime}}\rangle\langle\phi^{(N)}_{n,J^{\prime},M^{\prime}}e^{i\vec{K}\cdot\vec{R}}|U|\Psi_{J_{0},M_{0}}^{\mathrm{CDCC}}(\vec{K}_{0})\rangle$
$=\sum_{n,J^{\prime},M^{\prime}}\langle\phi_{\vec{k};I\mu;s\sigma}^{(-)}|\phi^{(N)}_{n,J^{\prime},M^{\prime}}\rangle T_{M_{0},M^{\prime}}^{n,J_{0},J^{\prime}}(\vec{K}),$ (35)
where the transition matrix elements $T_{M_{0},M^{\prime}}^{n,J_{0},J^{\prime}}(\vec{K})$ are to be interpolated from the discrete ones $T_{M_{0},M^{\prime}}^{n,J_{0},J^{\prime}}(\theta_{i},K_{i})$. Expressions for the overlaps between the final scattering states and the discrete states, $\langle\phi_{\vec{k};I\mu;s\sigma}^{(-)}|\phi^{(N)}_{n,J^{\prime},M^{\prime}}\rangle$, are given explicitly in Refs. Tos01 and Die17 for bin and PS functions, respectively.
The transition amplitudes of Eq. (35), $T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})$, contain the dynamics of the process in the coordinates describing the relative and center of mass motion of the core and the valence particle. From these amplitudes one can derive two-body observables for a fixed spin of the core, $I$, the solid angles describing the orientations of $\vec{k}$ ($\Omega_{k}$) and $\vec{K}$ ($\Omega_{K}$), as well as the relative energy between the valence and the core, $E_{\mathrm{rel}}$. These observables factorize into the transition matrix elements and a kinematical factor:
$\frac{d^{3}\sigma^{(I)}}{d\Omega_{k}d\Omega_{K}dE_{\mathrm{rel}}}=$ $\frac{\mu_{bx}k_{I}}{(2\pi)^{5}\hbar^{6}}\frac{K}{K_{0}}\frac{\mu_{aA}^{2}}{2J_{0}+1}\sum_{\mu,\sigma,M_{0}}|T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})|^{2},$ (36)
where $\mu_{bx}$ and $\mu_{aA}$ are the valence-core and projectile-target reduced masses.
The three-body observables, assuming the energy of the core is measured, are given by Tos01 :
$\frac{d^{3}\sigma^{(I)}}{d\Omega_{b}d\Omega_{x}dE_{b}}$ $=\frac{2\pi\mu_{aA}}{\hbar^{2}K_{0}}\frac{1}{2J_{0}+1}$
×$\times\sum_{\mu,\sigma,M_{0}}|T_{\mu\sigma;M_{0}}^{Is;J_{0}}(\vec{k},\vec{K})|^{2}\rho(\Omega_{b},\Omega_{x},E_{b}),$ (37)
where the phase space term $\rho(\Omega_{b},\Omega_{x},E_{b})$, i.e., the number of states per unit core energy interval at solid angles $\Omega_{b}$ and $\Omega_{x}$, takes the form Fuc82 :
$\rho(\Omega_{b},\Omega_{x},E_{b})$ $=\frac{m_{b}m_{x}\hbar k_{b}\hbar k_{x}}{(2\pi\hbar)^{6}}$
×$\times\left[\frac{m_{A}}{m_{x}+m_{A}+m_{x}(\vec{k}_{b}-\vec{K}_{tot})\cdot\vec{k}_{x}/k_{x}^{2}}\right].$ (38)
Here, the particle masses are given by $m_{b}$ (core), $m_{x}$ (valence), and $m_{A}$ (target) while $\hbar\vec{k}_{b}$ and $\hbar\vec{k}_{x}$ are the core and valence particle momenta in the final state. The total momentum of the system corresponds to $\hbar\vec{K}_{tot}$ and the connection with the momenta in Eq. (35) is made through:
$\vec{K}=\vec{k}_{b}+\vec{k}_{x}-\frac{m_{a}}{M_{tot}}\vec{K}_{tot};\quad\vec{k}=\frac{m_{b}}{m_{a}}\vec{k}_{x}-\frac{m_{x}}{m_{a}}\vec{k}_{b}$ (39)
with $m_{a}=m_{b}+m_{x}$ and $M_{tot}=m_{b}+m_{x}+m_{A}$ the total masses of the projectile and the three-body system, respectively.
An application of this formalism is presented in Fig. 19, corresponding to the breakup of 11Be on a proton target at $E_{p}=63.7$ MeV/u. The left panel corresponds to the differential cross section with respect to the final $n$+10Be relative energy and the right panel to the breakup with respect to the final 10Be energy (integrated in the 10Be and neutron angles). These observables were computed by transforming the XCDCC breakup transition amplitudes by means of Eqs. (36) and (37), and integrating over the unmeasured variables. The XCDCC calculations were performed with a 11Be model including the 10Be ground ($0^{+}$) and first excited ($2^{+}$) states. The formalism allows to separate and quantify the contribution of these two states of 10Be. Note that, due to energy conservation, in the relative energy distribution the contribution coming from the 10Be($2^{+}$) state contributes only above the excitation energy of this state ($E_{x}=3.367$ MeV).
Figure 19: Differential breakup cross sections, with respect to the $n$-10Be relative energy (a) and with respect to the final 10Be energy (b), for the breakup of 11Be on protons at 63.7 MeV/nucleon. The contributions of the 10Be ground ($0^{+}$) and first excited ($2^{+}$) states are shown in each panel. In the left panel, the result of the calculation neglecting the 10Be excitation mechanism (labelled “no DCE”) is also shown. The vertical arrow denotes the threshold for 10Be($2^{+}_{1}$)+n. Figure adapted from Ref. Die17 .
3.7.3 Pseudostates versus bins
As discussed in the preceding sections, discretization procedures employed in the CDCC method are either based on the pseudostate (PS) method or the binning method. Whereas reactions observables should be independent of the adopted discretization method (provided the calculations are fully converged), in practical applications one of the two methods might result more convenient than the other.
The PS discretization method becomes particularly suitable when dealing with narrow resonances. In the case of bins, one needs fine discretization in order to obtain a detailed description of the resonance region in the relative energy spectrum. By contrast, using PS’s one can obtain a detailed description of the resonance profile with a relatively small basis, with the aid of the convolution procedure discussed in Sec. 3.7.2. This is illustrated in Fig. 20, where we show the relative energy differential cross section corresponding to the 11Be+p $\to$ 10Be+n+p reaction at an incident energy of 63.7 MeV/nucleon. The 11Be bound and continuum states were generated with the structure model of Ref. Cap04 , which does not account for 10Be excitations. Continuum waves $s_{1/2}$, $p_{1/2}$, $p_{3/2}$, $d_{3/2}$, $d_{5/2}$ were included. For the CDCC-Bin calculation, the continuum was truncated at $E_{\mathrm{rel}}=12$  MeV and divided into 12 bins for each partial wave, evenly spaced in momentum. The relative-energy differential cross section was obtained by dividing the cross section of each bin by the width of that bin [see Eq. (31)]. The CDCC-PS calculations employ a THO basis with $N=30$ ($N=25$) oscillator functions for $\ell=1,2$ ($\ell=0$). After diagonalization of the 11Be Hamiltonian on this basis, only the eigenvalues below 12 MeV were retained. This results in 13 states for each partial wave (15 for $\ell=0$). The calculated breakup scattering amplitudes were then convoluted using Eq. (35), from which the differential cross sections, Eq. (36) was then evaluated. Finally, the latter was integrated over the solid angles $\Omega_{k}$ and $\Omega_{K}$.
Figure 20: Comparison of pseudostate (PS) and binning methods for the 11Be+p $\to$ 10Be+n+p reaction at 63.7 MeV/nucleon. Both methods use the same structure model Cap04 and include continuum states with configuration $s_{1/2}$, $p_{1/2}$, $p_{3/2}$, $d_{3/2}$, $d_{5/2}$. For the CDCC-PS calculation, the dominant individual contributions are shown separately.
As seen in Fig. 20, the binning and PS methods yield, as expected, almost identical results but the PS method allows for a finer description of the resonance region with a comparable, or even smaller, number of basis functions.
However, there are other situations where the binning procedure turns out to be more efficient than the PS method. One such situation is the case of reactions of highly polarizable weakly-bound nuclei (such as neutron-halo nuclei) on high-Z targets. These reactions are characterized by the presence of strong long-range Coulomb couplings, which tend to emphasize low-lying excitation states and probe large separations. In this case, the binning method has been shown to provide more stable results with respect to the basis size.
4 Transfer reactions with weakly bound nuclei
Transfer reactions are key spectroscopic tools for both stable and unstable nuclei. Angular distributions of outgoing fragments provide information on the angular momentum content of the populated states and the magnitude is closely related to the single-particle content of these states, quantified in terms of spectroscopic factors. The standard tool for analysing transfer reactions, the distorted-wave Born approximation (DWBA) method, assumes that the transfer occurs in a single-step. Possible excitations of the colliding or outgoing nuclei are ignored or, at most, taken into account effectively through the choice of the effective interactions. These interactions are typically chosen so as to reproduce the corresponding elastic scattering cross sections in the incident and exit channels.
Figure 21: Relevant coordinates for a transfer reaction of the form $A(d,p)B$.
To see how this method works in practice, let us consider as a particular case a process of the form $A(d,p)B$, schematically depicted in Fig. 21. Using the post-form representation, the transition matrix for this process can be written as
${\cal T}_{dp}=\langle\chi^{(-)}_{p}\Phi_{B}|V_{pn}+U_{pA}-U_{pB}|\Psi^{(+)}_{d}\rangle,$ (40)
where $V_{pn}$, $U_{pA}$ are the proton-neutron and proton-target interactions, $U_{pB}$ is an auxiliary (and, in principle, arbitrary) potential for the $p$-$B$ system, $\Phi_{B}$ is the internal wave function of the residual nucleus $B$ and $\Psi^{(+)}_{d}$ is the total wave function corresponding to an incident deuteron beam of kinetic energy $E_{d}$ and binding energy $\varepsilon_{d}$. If target excitation is not considered explicitly, this total wavefunction can be approximated as
$\Psi^{(+)}_{d}({\vec{r}},{\vec{R}},\xi_{A})\approx\Psi^{3b}({\vec{r}},{\vec{R}})\Phi^{(0)}_{A}(\xi_{A})$ (41)
where $\Phi^{(0)}_{A}(\xi_{A})$ denotes the target ground-state wavefunction.
Ignoring antisymmetrization for simplicity, the wave function $\Phi_{B}$ for a total angular momentum $J$ and projection $M$ can be expanded in $A$ states using the usual parentage decomposition
$\Phi^{JM}_{B}({\vec{r}}_{nA},\xi)=\sum_{I,\ell,j}[\Phi^{I}_{A}(\xi)\otimes\phi_{\ell j}({\vec{r}}_{nA})]_{JM},$ (42)
where $I$ is the spin of $A$, $l$ and $j$ the orbital and total ($\vec{j}=\vec{\ell}+s$) angular momentum of the valence particle and $\phi_{\ell j}({\vec{r}}_{nA})$ is a function describing the neutron-core relative motion. The normalization
[…]
can be regarded as a spectroscopic factor for the configuration $\{\ell,j\}$. [^2] Strictly, the functions $\phi_{lj}$ and the spectroscopic factors ${\cal S}_{\ell j}$ depend in general on the core state $I$ but, in the present case, this quantum number can be readily inferred from the $l,j$ values, so it is omitted for brevity.
If the interactions $U_{pA}$ and $U_{pB}$ are assumed to be independent of the target degrees of freedom ($\xi_{A}$), the integral in $\xi_{A}$ can be readily performed, transforming (40) into
${\cal T}^{3b}_{dp}=\sqrt{S_{\ell,j}}\langle\chi^{(-)}_{p}\phi_{\ell j}|V_{pn}+U_{pA}-U_{pB}|\Psi^{3b,(+)}_{d}\rangle,$ (44)
In DWBA, the three-body wavefunction $\Psi^{3b,(+)}_{d}$ is further approximated by $\Psi^{\mathrm{(+)}}\approx\chi^{(+)}_{d}({\vec{R}})\varphi_{d}({\vec{r}})$. When the transfer reaction involves weakly-bound nuclei, including the deuteron discussed here, this choice is not well justified. Due to the presence of the short-range $V_{pn}$ interaction, the evaluation of DWBA matrix element is mostly sensitive to small $p$-$n$ separations. These configurations do not only contain contributions coming from the deuteron ground state, but also from $p$-$n$ unbound states. By contrast, the deuteron optical potential describing deuteron-target elastic scattering is not restricted to small $p$-$n$ distances. Therefore, the use of the deuteron optical potential in the $(d,p)$ or $(p,d)$ transfer amplitude is likely to lack important deuteron breakup components. This problem was recognized long ago and several solutions have been proposed. One of the first and most popular ones is the adiabatic method of Johnson and Soper Ron70 .
The Johnson-Soper approximation
In the Johnson-Soper (JS) approximation, this effect is approximately taken into account by means of the choice
$\Psi_{d}^{(+)}(\vec{R},\vec{r})\simeq\chi_{JS}^{(+)}({\vec{R}})\phi_{d}(\vec{r}),$ (45)
where $\chi_{JS}^{(+)}(\vec{R})$ is the solution of a two-body scattering problem, on the coordinate $\vec{R}$, in which the interaction is given by a $p-n$ zero-range approximation:
$U_{JS}(R)=U_{pA}(R)+U_{nA}(R).$ (46)
We see that, in this limit, the adiabatic theory of the transfer amplitude adopts a form akin to that found in DWBA, but this analogy is only formal because the function $\chi_{JS}^{(+)}({\vec{R}})$ includes contributions from breakup and the potential $U_{JS}(R)$ may have little to do with the optical potential describing the deuteron elastic scattering. Due to the adiabatic approximation, the JS theory is not expected to be accurate at low incident energies.
The adiabatic approximation is equivalent to neglecting the excitation energy of the states of the projectile Ron70 . The adiabatic wave function takes into account the excitation to breakup channels, assuming that these states are degenerate in energy with the projectile ground state, as illustrated in Fig. 22(b). Therefore, the ADWA approach takes into account, approximately, the effect of deuteron break-up on the transfer cross section, within the adiabatic approximation. So, it should be well suited to describe deuteron scattering at high energies, around 100 MeV per nucleon. Systematic studies Har71 ; Sat71 ; Wal76 have shown that ADWA is superior to standard DWBA for $(d,p)$ scattering at these relatively high energies.
The Johnson-Tandy approximation
Although the zero-range adiabatic model of JS provides a systematic improvement over the conventional DWBA, there are situations in which the former fails to reproduce the experimental data Ste86 ; Ron89 . Models which go beyond the zero-range and adiabatic approximations are therefore needed. One of such models is the Weinberg expansion method of Johnson and Tandy (JT) JT74 . The idea is to expand $\Psi^{\mathrm{(+)}}$ in terms of a set of functions which are complete within the range of $V_{pn}$. A convenient choice is the set of Weinberg states (also called Sturmians), given by
$\Psi^{(+)}(\vec{R},\vec{r})=\sum_{i=0}^{N}\phi^{W}_{i}(\vec{r})\chi^{W}_{i}(\vec{R})\,,$ (47)
where $\phi^{W}_{i}(\xi)$ are the Weinberg states, which are solutions of the eigenvalue equation
$[T_{\vec{r}}+\alpha_{i}V_{pn}]\phi^{W}_{i}({\vec{r}})=-\varepsilon_{d}\phi^{W}_{i}({\vec{r}}),$ (48)
where $\varepsilon_{d}=2.225$ MeV is the deuteron binding energy and where $\alpha_{i}$ are the eigenvalues, to be determined along with the eigenfunctions. Beyond the range of the potential all the Weinberg states decay exponentially, like the deuteron ground-state wave function. For $i=0$, $\alpha_{0}=1$ and so $\phi^{W}_{0}({\vec{r}})$ is just proportional to the deuteron ground state. As $i$ and $\alpha_{i}$ increase, they oscillate more and more rapidly at short distances. The Weinberg states form a complete set of functions of ${\vec{r}}$ within the range of the potential $V_{pn}$. They are well suited to expand $\Psi^{(+)}$ in this region, as it is required by the amplitude (40). They do not satisfy the usual orthonormality relation but the less conventional one
$\langle\phi^{W}_{i}|V_{pn}|\phi^{W}_{j}\rangle=-\delta_{ij}.$ (49)
If we retain in (47) only the leading term, $\Psi^{(+)}(\vec{R},\vec{r})\approx\phi^{W}_{0}\chi^{W}_{0}(\vec{R})$, one finds JT74 that $\chi^{W}_{0}$ verifies the single-channel equation
$[T_{\vec{R}}+U_{JT}({\vec{R}})-E_{d}]\chi^{W}_{0}({\vec{R}})=0,$ (50)
with $E_{d}=E-\varepsilon_{d}$ and where the potential $U_{JT}$ is given by:
$U_{JT}(R)=\frac{\langle\phi_{pn}(\vec{r})|V_{pn}(U_{nA}+U_{pA})|\phi_{pn}(\vec{r})\rangle}{\langle\phi_{pn}(\vec{r})|V_{pn}|\phi_{pn}(\vec{r})\rangle}.$ (51)
The bra and ket in this equation mean integration over ${\vec{r}}$, with fixed ${\vec{R}}$. Interestingly, in the zero-range limit, $U^{JT}(R)$ reduces to the JS potential, eq. (46). Therefore, the zero-order result given by eq. (51) can be regarded as a finite-range version of the adiabatic (JS) potential. These two models are globally referred to as Adiabatic Distorted Wave Approximation (ADWA). However, it is worth noting that the full Weinberg expansion makes no reference to the incident energy and, as such, does not involve the adiabatic approximation. This suggests that a stripping theory based on this Weinberg expansion can be used at low energies, where the adiabatic condition is not well satisfied. The inclusion of higher-order terms ($i>0$) in the Weinberg expansion has been investigated in refs. Lai93 ; Pan13 .
Figure 22: Schematic representation of DWBA, ADWA and CDCC-BA approaches for a $(d,p)$ transfer reaction.
In Fig. 23 we present a comparison between the DWBA and ADWA methods for the 56Ni(p,d)57Ni (left) and 48Ca(d,p)49Ca reactions. For the ADWA calculations, the prescriptions of JS and JT are shown. It is clearly seen that the ADWA, in both its zero-range and finite-range forms, provides an improved description of the data as compared to the DWBA method.
Figure 23: Comparison of the ADWA and DWBA methods for the 56Ni($p$,$d$)57Ni (left) and 48Ca($d$,$p$)49Ca reactions at incident energies of 37 MeV and 10 MeV, respectively. The nucleon optical potentials are taken from the global CH89 systematics while Daenick global potential is used for the deuteron optical potential in the DWBA calculations. For the ADWA calculations, the prescriptions of Johnson-Soper (JS) and Johnson-Tandy (JT) are shown. Quoted from Ref. Tim20 .
An appealing feature of the ADWA method is that its ingredients are completely determined by experiments. These ingredients are the proton-target and neutron-target optical potentials, evaluated at half of the deuteron incident energy, as well as the well-known proton-neutron interaction. On the negative side, the ADWA approach does not consistently describe elastic scattering and nucleon transfer. Although, physically, elastic scattering, transfer and break-up should be closely related by flux conservation, this connection is not present in ADWA. Furthermore, the arguments leading to ADWA are strongly dependent on the assumption that the transfer is governed by a short-range operator. Thus, it is not obvious that the approximations remain valid for other weakly bound systems, like 11Be. Even in the case of $(d,p)$ scattering, the transfer matrix element is determined, in addition to the $n$-$p$ interaction, by the proton-target and proton-composite interactions which define the remnant term. The role of these terms, which would have contributions of three-body configurations in which proton and neutron are not so close together, is not clear a priori. The latter problem can be avoided by using an alternative expression for the scattering amplitude. Following Goldberg and Watson Gol64 , one can choose the auxiliary interaction $U_{pB}$ that appears in the remnant term of Eq. (44) to cancel exactly the core-target interaction, $U_{pA}$. This results in an alternative, but still exact, scattering amplitude (the SF is omitted for simplicity), namely,
${\cal T}^{\mathrm{3b_{GW}}}_{dp}=\langle\tilde{\Phi}^{(-)}_{pB}|V_{pn}|\Psi^{3b,(+)}_{d}\rangle,$ (52)
where $\tilde{\Phi}^{(-)}$ is a solution of the three-body equation:
$[T_{{\vec{r}}^{\prime}}+T_{{\vec{R}}^{\prime}}+U_{pB}+U-E]\tilde{\Phi}^{(-)}_{pB}=0.$ (53)
In Tim99 Timofeyuk and Johnson use an adiabatic approximation for $\tilde{\Phi}^{(-)}_{pB}$ to produce a tractable expression that still includes recoil excitation and breakup effects (REB), while keeping the matrix element constrained to the range of the $V_{pn}$ interaction:
$\tilde{\Phi}^{(-)}_{pB}\simeq\chi^{(-)}_{pA}(\boldsymbol{r}_{pA},\boldsymbol{k}_{\alpha})\phi_{nA}(\boldsymbol{r}_{nA})e^{-i\alpha\boldsymbol{k}_{\alpha}\boldsymbol{r}_{nA}}.$ (54)
In this expression, $\boldsymbol{k}_{\alpha}$ is the relative momentum between $B$ and $p$ and $\alpha=m_{n}/m_{B}$, and it was found to produce a good agreement between this calculation and experimental data for the 16O$(d,p)^{17}$ and 10Be$(d,p)^{11}$e reactions. Nonetheless, given that this expression is better suited for the transfer of weakly-bound and halo nuclei, the ADWA model has enjoyed more widespread use.
The CDCC-BA approximation
Another way of accounting for the breakup channels in transfer reactions is by insertion of the CDCC wavefunction, Eq. (8), in the transition amplitude (44). The resultant amplitude is formally analogous to that found in the CCBA method Sat83 , so we refer to it as CDCC-BA approximation. This is expected to be a good approximation since the CDCC expansion is accurate for small $p$-$n$ separations, for which the matrix transition amplitude presents the largest magnitude. The resultant amplitude, however, will be clearly much more computationally demanding than that obtained with the DWBA or ADWA methods.
In Fig. 24 we compare the performance of the DWBA, ADWA and CDCC-BA methods for the 58Ni($d$,$p$)59Ni reaction at $E_{d}=10$ MeV (upper panel) and $E_{d}=56$ MeV (lower panel), taken from Ref. Pan14 . It becomes clear that the ADWA and CDCC methods provide a more reliable description of the data compared to the traditional DWBA approach. In Cha17 , a systematic comparison between the CDCC-BA method and an adiabatic approximation (CDCC-AD) was performed for $(d,p)$ reactions on various targets, showing good agreement between both, except in some cases, such as when the deuteron beam energy is small and similar to the binding energy of the nucleon in the target or when the binding energy of the nucleon is very small ($S_{n}~{}\simeq 0.1$ MeV). It should be noted that in this work the adiabatic approximation is performed by setting the nominal energies of the continuum bins to that of the bound state, which is formally equivalent to the ADWA approach, although the calculation is rather different, so extensions of these results to standard ADWA calculations should be made with caution.
The CDCC-BA approximation is reasonable as long as the coupling to the rearrangement (transfer) channels is weak, typically, when the transfer cross section is much smaller than the reaction cross section. Otherwise, other methods that treat transfer on equal footing to breakup, such as the Coupled-Reaction-Channel method (CRC) Sat83 , are required.
Figure 24: Comparison of the DWBA, ADWA and CDCC-BA methods for 58Ni(d,p)59Ni reaction at $E_{d}=10$ MeV (upper panel) and $E_{d}=56$ MeV (lower panel). Adapted from Ref. Pan14 .
4.1 Transfer reactions populating unbound systems
So far, we have considered transfer reactions as a tool for investigating the bound states of a given nucleus. However, in a rearrangement process, the transferred particle can also populate unbound states of the final nucleus. This opens the possibility of studying and characterizing structures in the continuum, such as resonances or virtual states. In fact, since weakly-bound nuclei usually have almost no excited bound states and their binding energy is low, the matching conditions for transfer reactions Bri72 favours the population of highly excited states of the residual nucleus, including continuum states.
As in the case of transfer to bound states, the simplest formalism to analyze these processes is the DWBA method. In this case, the bound wavefunction $\phi_{\ell,j}({\vec{r}}^{\prime})$ appearing in the final state in Eq. (44) should be replaced by a positive-energy wavefunction describing the state of the transferred particle (neutron in this case) with respect to the core target. In principle, for this purpose, one could use the suitable scattering state of the $v$+$b$ system at the appropriate relative energy. However, this procedure tends to give numerical difficulties in evaluating the transfer amplitude due to the oscillatory behaviour of both the final distorted wave and the wavefunction $\phi_{\ell,j}(\vec{r}^{\prime})$. To avoid this problem, several alternative methods have been used. We enumerate here some of them:
  1. (i)
    The bound state approximation Cok74 . In the case of transfer to a resonant state, this method replaces the scattering state $\varphi_{\ell,j}({\vec{r}}^{\prime})$ by a weakly bound wavefunction with the same quantum numbers $\ell$ and $j$. In practice, this can be achieved by starting with the potential that generates a resonance at the desired energy and increase progressively the depth of the central potential until the state becomes bound.
  2. (ii)
    Huby and Mines Hub65 used a scattering state for $\phi_{\ell,j}({\vec{r}}^{\prime})$ modulated by a convergence factor $e^{-\alpha r^{\prime}}$ (with $\alpha$ a positive real number), which is intended to eliminate its contribution to the integral coming from large $r^{\prime}$ values, and then extrapolate numerically to the limit $\alpha\rightarrow 0$.
  3. (iii)
    Vincent and Fortune VF70 put into question the validity of the bound state approximation arguing that, in general, the bound state and resonant form factors can be very different and, even in those cases in which the fictitious form factor gives the correct shape, they can lead to very different absolute cross sections. They suggest using the actual scattering state, but choosing an integration contour along the complex plane in such a way that the oscillatory integrand is transformed into an exponential decay, thus palliating the slow convergence problem of the post-form transfer amplitude.
  4. (iv)
    In a real transfer experiment leading to positive-energy states, one does not have access to a definite final energy, but to a certain region of the continuum. That is to say, the extracted observables, such as energy differential cross sections, are integrated over some energy range which, at least, is of the order of the energy resolution of the experiment. This suggests a method of dealing with the unbound states consisting of discretizing the continuum states in energy bins, as in the CDCC approximation.
Figure 25: Radial part of the $d_{3/2}$ single-particle resonance wavefunction in 17O at $E_{r}=0.95$ MeV compared with a slightly bound wavefunction ($E=-0.1$ MeV) and a bin wavefunction, centered at the nominal energy of the resonance and with a width of 0.5 MeV.
In Fig. 25, we show as an example the radial part of a $3/2^{+}$ resonance in 17O, described in terms of a $d_{3/2}$ neutron coupled to a zero-spin 16O core. The solid line is a scattering wavefunction evaluated at the nominal resonance energy ($E_{rel}=0.95$ MeV). Note the oscillatory behaviour at large distances. The dotted line is a bin wavefunction, constructed by a superposition of scattering states, within the range of 0.5 MeV around the resonance energy. It is seen that, asymptotically, the oscillations are damped with respect to the original scattering states. Finally, the dot-dashed line is a bound state wavefunction, with a $1d_{3/2}$ single-particle configuration, and a separation energy of 0.1 MeV. This wavefunction is very similar to the scattering state at short distances but decays exponentially at large distances.
An advantage of the method (iv) is that it can be applied to situations in which one is interested in the description of the population of a range of continuum energies possibly covering both resonant and non-resonant states. For that, it is convenient to resort to the prior-form expression of the transition amplitude. Considering again the $A(d,p)B$ reaction for simplicity, this amplitude reads
$\mathcal{T}^{\mathrm{prior}}_{if}=\langle\Psi_{f}^{(-)}(\vec{R}\mkern 2.0mu\vphantom{R}^{\prime},\vec{r}\mkern 2.0mu\vphantom{r}^{\prime})|V_{nA}+U_{pA}-U_{dA}|\phi_{d}(\vec{r})\chi_{dA}^{(+)}(\vec{R})\rangle.$ (55)
The function $\chi^{(+)}_{dA}$ is the distorted wave generated by the optical potential $U_{dA}$ and $\Psi^{(-)}_{f}(\vec{R}\mkern 2.0mu\vphantom{R}^{\prime},\vec{r}\mkern 2.0mu\vphantom{r}^{\prime})$ is the exact three-body wave function for the final $p$+$n$+A state, with $\vec{r}\mkern 2.0mu\vphantom{r}^{\prime}$ and $\vec{R}\mkern 2.0mu\vphantom{R}^{\prime}$ denoting the $n$-$A$ and $p$-$B$ relative coordinates, respectively. The wave function $\Psi_{f}^{(-)}(\vec{R}\mkern 2.0mu\vphantom{R}^{\prime},\vec{r}\mkern 2.0mu\vphantom{r}^{\prime})$ is the time-reversed of $\Psi_{f}^{(+)}(\vec{R}\mkern 2.0mu\vphantom{R}^{\prime},\vec{r}\mkern 2.0mu\vphantom{r}^{\prime})$, which satisfies the three-body equation:
$[T_{\vec{R}\mkern 2.0mu\vphantom{R}^{\prime}}+T_{\vec{r}\mkern 2.0mu\vphantom{r}^{\prime}}+V_{pn}+U_{pA}+V_{nA}-E]\Psi_{f}^{(+)}(\vec{R}\mkern 2.0mu\vphantom{R}^{\prime},\vec{r}\mkern 2.0mu\vphantom{r}^{\prime})=0,$ (56)
with $E$ the total energy of the system. To solve this equation, the wavefunction $\Psi^{(-)}_{f}(\vec{R}\mkern 2.0mu\vphantom{R}^{\prime},\vec{r}\mkern 2.0mu\vphantom{r}^{\prime})$ can be expanded in $n+A$ states with well-defined energy and angular momentum, as in the continuum-discretized coupled-channels (CDCC) method.
An example is shown in Fig. 26, which corresponds to the differential cross section, as a function of the $n$-9Li relative energy, for the reaction 2H(9Li,$p$)10Li at 2.36 MeV/u (upper) and 11.1 MeV/u (lower panel), corresponding to the measurements of Refs. Jep06 and TRIUMF Cav17 , respectively. The lines are the results of transfer-to-the-continuum calculations populating 10Li continuum states using the same 10Li structure in both cases. The figure shows the separate contribution of the s-wave ($1^{-}$, $2^{-}$) and p-wave ($1^{+}$, $2^{+}$) continuum states. The strength of the measured cross section close to zero energy is due to the presence of a virtual state in the 10Li continuum, whereas the peak around 0.4 MeV is due to a $p_{1/2}$ resonance. This is an example of how the use of transfer reactions can provide information on the continuum structure of weakly-bound or even unbound systems. For more details on the calculations, see Ref. Mor19 .
Figure 26: Illustration of the transfer-to-the-continuum method, using a binning discretization, for the reaction 2H(9Li,p)10Li. Calculations are compared with the experimental data from Refs. Jep06 and Cav17 . Figure taken from Ref. Mor19 .
4.2 Transfer reactions involving three-body projectiles
The formalism presented in the previous sections can be extended and applied to more complex systems. We discuss here the interesting case of the stripping reaction induced by a three-body Borromean system on a proton target, in which one of the projectile fragments is transferred to the target, leaving a two-body residual unbound system. To be more specific, we consider the case of a two-neutron halo projectile, such as 11Li, 6He. In these cases, given the unstable nature of the projectile, the reaction experiment must be performed in inverse kinematics. This process can be schematically represented as (see Fig. 27).
$\underbrace{(C+N_{1}+N_{2})}_{A}+p\rightarrow\underbrace{(C+N_{2})}_{B}+d,$ (57)
The prior-form transition amplitude for this process can be formally reduced to an effective few-body problem, leading to
$\mathcal{T}_{if}=2~{}\langle\Psi_{f}^{(-)}(\vec{x},\vec{R}^{\prime},\vec{r}^{\prime})|V_{pN_{1}}+U_{pB}-U_{pA}|\Phi_{A}(\vec{x},\vec{y})\chi_{pA}^{(+)}(\vec{R})\rangle,$ (58)
where $\Phi_{A}$ represents the ground-state wave function of the initial three-body composite, $\chi^{(+)}_{pA}$ is the distorted wave generated by the auxiliary potential $U_{pA}$, and $\Psi^{(-)}_{f}$ is the exact four-body wave function for the outgoing $d$-$B$ system. Notice the explicit factor 2 coming from the two identical neutrons in the initial wave function.
Figure 27: Diagram for a $(p,d)$ or $(p,pN)$ reaction induced by a three-body projectile in inverse kinematics. Taken from Ref. Gom17c .
In writing the transition amplitude in the form (58), one implicitly uses a participant/spectator approximation, assuming that the reaction occurs due to the interaction of the incident proton with a single neutron $(N_{1})$ of $A$ (the participant), leaving the residual $B=N_{2}+C$ system unperturbed.
To reduce (58) to a tractable form, one can approximate the exact wave function $\Psi^{(-)}_{f}$ by the factorized form,
$\Psi_{f}^{(-)}(\vec{x},\vec{R}^{\prime},\vec{r}^{\prime})\approx\varphi^{(-)}_{\vec{q},\sigma_{2},\zeta}(\vec{x})\chi_{dB}^{(-)}(\vec{R}^{\prime})\phi_{d}(\vec{r}^{\prime}),$ (59)
where $\phi_{d}$ is the deuteron wave function, $\varphi^{(-)}_{\vec{q},\sigma_{2},\zeta}$ is a two-body continuum wave function with wave number $\vec{q}$ and spin projections ${\sigma_{2},\zeta}$ of the binary subsystem $B$, and $\chi_{dB}^{(-)}$ is a distorted wave describing the $d$-$B$ relative motion in the exit channel. The function $\varphi^{(-)}_{\vec{q},\sigma_{2},\zeta}$ is the time-reversed of $\varphi^{(+)}_{\vec{q},\sigma_{2},\zeta}$, and presents incoming scattering boundary conditions.
The ground state wavefunction of the initial three-body composite $A$ ($\Phi_{A}^{j\mu}(\vec{x},\vec{y})$) is conveniently described using a three-body model Des03 ; FaCE ; Rod05 , as introduced for the four-body CDCC calculations in Sec. 3.5. Typically, the three-body wave function is most naturally obtained in the so-called Jacobi-T set, but for the purpose of computing the required overlap functions, it is then transformed into the Jacobi-Y set, where the $x$ coordinate relates the core and one neutron (see Fig. 27).
Consistently with the spectator approximation for the composite $B$, it is assumed that the interaction does not change the state of this system, and the overlap between the two- and three-body wave functions ($\psi(\vec{q},\vec{y})$) contains all the relevant structure information:
[…]
from which the following transition amplitude can be defined:
$\mathcal{T}_{if}\equiv\langle\chi_{dB}^{(-)}\phi_{d}|V_{pN_{1}}+U_{pB}-U_{pA}|\psi~{}\chi_{pA}^{(+)}\rangle.$ (61)
From the transition amplitude the differential cross section of the $C$-$N_{2}$ relative momentum $\vec{q}$ and the scattering angle of $B$ with respect to the incident direction can be written as
$\frac{d\sigma^{2}}{d\Omega_{B}d\vec{q}}$ $=\frac{2}{(2j+1)}\frac{\mu_{i}\mu_{f}}{(2\pi\hbar^{2})^{2}}\frac{k_{f}}{k_{i}}\sum_{\nu}\left|\mathcal{T}_{if}\right|^{2},$ (62)
where $\mu_{i,f}$ the projectile-target reduced mass in the initial and final partitions, and $\nu\equiv\{M_{T}\sigma_{d}\}$ represents the spin projections of the final products.
After integration over the angle of $\vec{q}$ and transforming from momentum to energy, expression (62) provides double differential cross section with respect to the scattering angle and the excitation energy of the residual $N_{2}+C$ system. This observable can in principle be reconstructed experimentally from the measured energy and angle of the outgoing deuteron. To illustrate this formalism, we show in Fig. 28 results for the reaction 11Li(p,d)10Li Cas17 , recently measured at TRIUMF by Sanetullaev et al. San16 .
Figure 28: (a) Relative $n$-9Li energy spectrum, integrated in angle, for $s_{1/2}$ (green) and $p_{1/2}$ contributions (red) for the 11Li(p,d )10Li reaction at 5.7 MeV/u. evaluated with the prior-form DWBA calculations described in the text. The summed contributions, before (blue dashed) and after (blue solid) convoluting with the experimental resolution are also shown. The $1^{-}$ and $2^{-}$ contributions are shown separately as well as $1^{+}$ and $2^{+}$. (b) Energy-integrated angular differential cross section populating the $s_{1/2}$ and $p_{1/2}$ orbitals in 10Li. The data of San16 , corresponding to the resonant contribution ($p$-wave) is also shown. Figure adapted from Ref. Cas17 .
For the calculations presented, it was assumed that the two halo neutrons of 11Li have the same angular momentum $\vec{J}=\vec{L}+\vec{s}_{2}$. This condition is strictly fulfilled when the spin of the core is neglected and is also a good approximation when using $I=3/2^{-}$. Only the configurations $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ were considered, as other contributions, while present in the three-body calculations, give negligible components in the ground state of 11Li. The considered 9Li+n potential contains terms that couple the spin of the remaining neutron and 9Li in 10Li leading to splitting of the $p$-wave resonance into $1^{+}$ and $2^{+}$ states and of the virtual state in the $s$-wave in $1^{-}$ and $2^{-}$ components. These components are shown in the top panel of Fig. 28, which depicts the relative energy spectrum between $n$ and 9Li. In red, the $p-$wave components are presented, where the splitting of the $p$-wave resonance into $1^{+}$ and $2^{+}$ resonances at 0.37 and 0.61 MeV is evident, while in green the $s$-wave conributions are shown, presenting a virtual state (with scattering length $a=-37.9$ fm) only in the $2^{-}$ component. The total is shown in dashed blue, while its folding with the experimental resolution given in Ref. San16 is presented in the solid blue line. It is rather remarkable how the low-energy peaks are smoothed out by the experimental resolution leading to a relatively featureless distribution, which could explain why the fine details of 10Li are not observed in this experiment. In the bottom panel, the energy-integrated angular distribution of the outgoing deuteron is presented. It has been assumed that the subtraction of the non-resonant background performed in the analysis of the data completely excludes the $s$-wave component, so experimental data are compared only with the cross section corresponding to neutrons transferred from the $p_{1/2}$ orbitals, finding an excellent agreement both in shape and magnitude. Note that the experimental data have been measured at angles for which the $s_{1/2}$ contribution is minimal. This may explain why in our energy distribution the contribution from $s$ waves is quite important at low energies, while experimentally there is no direct sign of this effect.
Using the methods described in Sec. 4.1 to consider continuum states of the $p-N$ subsystem, a similar formalism can be used to study $(p,pN)$ reactions on three-body nuclei, simply considering the auxiliary amplitude (61) for continuum $p-N$ states, which are typically chosen to have a definite $p-N$ energy, angular momentum and parity:
$\mathcal{T}_{if}^{J^{\pi}_{pn}}(E_{pN})\equiv\langle\chi_{pNB}^{(-)}\phi^{J^{\pi}_{pn}}_{pN}|V_{pN_{1}}+U_{pB}-U_{pA}|\psi~{}\chi_{pA}^{(+)}\rangle.$ (63)
and where the outgoing $\chi_{pNB}^{(-)}$ functions can be computed using a CDCC solution for the final 3-body system Mor15 . Figures 29 shows results for the $n$-${}^{9}\mathrm{Li}$ relative energy distribution in the 11Li$(p,pn)^{10}$i reaction at 280 MeV/nucleon, using the same model for 11Li and 10Li as for the $(p,d)$ reaction in Fig. 28. The contributions of the different components are shown on the top panel, where again a splitting of the $p-$wave resonance is visible, as well as the virtual state only in the $2^{-}$ component, the main difference to the $(d,p)$ case being the increased importance of the $s$-wave components. In the bottom panel, the theoretical distribution is presented convoluted with the experimental resolution and compared with the experimental data, showing an excellent agreement with it, although the experimental resolution again significantly smooths the distribution.
It is a success of the model model that it is able to reproduce experimental data at $\sim$6 MeV per nucleon (the $(p,d)$ reaction) and at $\sim$280 MeV per nucleon (the $(p,pN)$ reaction), although perhaps not unexpected, given the similarities of both reactions when viewed as transfer to bound and continuum states of the $p-n$ system.
Figure 29: $n$-${}^{9}\mathrm{Li}$ relative energy distribution for 11Li$(p,pn)^{10}$i at 280 MeV/nucleon. In the top panel, the different contributions for the $s$ ($1^{-},2^{-}$) and $p$ ($1^{+},2^{+}$) waves within a three-body 11Li model including the 9Li spin Cas17 . The total cross section is given by the solid blue line. In the bottom panel, the results are compared, with and without convoluting with the experimental resolution, with the data from Ref. Aks08 . Figure adapted from Ref. Gom17c , with authorization from APS.
4.3 Simultaneous inclusion of projectile breakup and target excitation in transfer reactions
The models discusses in the previous subsections are based on a strict three- or four-body description of the transfer reaction in which possible excitations of the target ($A$) are only accounted for approximately by the choice of the effective interactions. From the analysis of reactions with ordinary nuclei, it is known that an explicit account of collective excitations of the colliding nuclei is required for an accurate description of the reaction. For $(d,p)$ reactions on deformed targets, one may anticipate that both projectile breakup and target excitation can play a role and may require their explicit inclusion. This is not possible in the standard formulations of the aforementioned CCBA and ADWA methods, which tend to emphasize one of the two mechanisms. As a way of including the two mechanisms simultaneously, one may resort to the extended CDCC wavefunction (24) which was devised to include in the model space the projectile breakup channel as well as some excited states of the target nucleus.
For the particular case of a $A(d,p)B$ reaction, and considering for simplicity only one excited state of the target, this extended CDCC wavefunction can be schematically expressed as
$\Psi^{(+)}_{d}({\vec{r}},{\vec{R}},\xi)$ $=\phi_{d}({\vec{r}})\Phi^{0}_{A}(\xi)\chi^{(+)}_{d,0}({\vec{R}})+\phi_{d}({\vec{r}})\Phi^{0}_{A}(\xi)\chi^{(+)}_{d,1}({\vec{R}})$
$+\sum_{n}\phi^{n}_{pn}({\vec{r}})\Phi^{\text{1}}_{A}(\xi)\chi^{(+)}_{n,0}({\vec{R}})+\sum_{n}\phi^{n}_{pn}({\vec{r}})\Phi^{\text{1}}_{A}(\xi)\chi^{(+)}_{n,1}({\vec{R}})$ (64)
where $\{\phi_{d},\phi^{n}_{pn}\}$ denote the deuteron ground state and (discretized) continuum states, and $\{\chi^{(+)}_{n,0}({\vec{R}}),\chi^{(+)}_{n,1}({\vec{R}})\}$ the functions describing the projectile-target relative motion with the target in either its ground state $\Phi^{0}_{A}(\xi)$ or in the excited state $\Phi^{\text{1}}_{A}(\xi)$, respectively. Therefore, the first two terms of (4.3) describe, respectively, elastic and inelastic scattering with the deuteron remaining in its ground state. The third and fourth terms describe deuteron breakup with respect to the target in its ground state or first excited state, respectively. When inserted into Eq. (55) this gives rise also to four terms,
${\cal T}_{dp}={\cal T}^{\mathrm{el}}_{dp}+{\cal T}^{\mathrm{inel}}_{dp}+{\cal T}^{\mathrm{elbu}}_{dp}+{\cal T}^{\mathrm{inbu}}_{dp},$ (65)
which may be interpreted as (I) elastic transfer, i.e., direct transfer from the deuteron ground state leaving the target in its ground state, (II) inelastic transfer, i.e., target excitation followed by one-neutron transfer, (III) elastic breakup transfer, i.e., deuteron breakup followed by transfer, leaving the target in the ground state and (IV) inelastic breakup transfer, i.e., deuteron breakup, accompanied by target excitation, and followed by neutron transfer.
Note that these four terms are to be added coherently, giving rise to interference effects, as discussed below. Note also that, in DWBA, only the first term (elastic transfer) is explicitly taken into account. One of our goals here is to study the contribution of the other terms.
As an example, we consider the reaction 10Be(d,p)11Be(g.s.), including both the 10Be(g.s.) and its first excited state ($2^{+}$; $E_{x}=3.367$ MeV). The different reaction paths described above are schematically shown in Fig. 30.
Figure 30: (Color online) Pictorial representation of the 10Be(d,p)11Be reaction, highlighting the different paths considered in this work. The initial $d+$10Be system includes breakup of the deuteron and excitation of 10Be coupled to all orders. The transfer channel is then coupled in first order to these channels, and four paths are distinguished: I (elastic transfer), II (inelastic transfer), III (breakup transfer) and IV (inelastic breakup transfer). Figure taken from Gom17b , with authorization from APS.
Calculations for this reaction have been reported in Ref. Gom17b . The structure of the 11Be nucleus is treated within the particle-rotor model of Ref. Nun96 , which assumes a deformation length of $\delta_{2}=1.664$ fm for 10Be and includes the ground state ($0^{+}$) and first excited state ($2^{+}$) of this nucleus.
In Fig. 31 we show the calculated angular distributions for the deuteron incident energies of $E_{d}=20$ MeV (top) and $E_{d}=60$ MeV (bottom). The different lines shown in each panel correspond to the paths of Fig. 30. In each panel, we show the full calculation (thick solid black line) along with the calculations keeping only one of the terms in Eq. (65). The red solid line corresponds to ${\cal T}^{el}_{dp}$ (path I in Fig. 31), the blue solid only to ${\cal T}^{inel}_{dp}$ (II), the red dashed only to ${\cal T}^{elbu}_{dp}$ (III) and the blue dashed only to ${\cal T}^{inbu}_{dp}$ (IV). As before, for the $d$+10Be partition all channels are coupled to all orders, while the transfer channel is coupled in first order. It can be seen that for both energies the main contributor is the elastic transfer ${\cal T}^{el}_{dp}$, while, at small angles, for $E_{d}=20$ MeV the second main contributor is the breakdown of the deuteron, ${\cal T}^{elbu}_{dp}$, whereas at 60 MeV it is the excitation of 10Be, ${\cal T}^{inel}_{dp}$
Figure 31: Differential transfer cross section for $E_{d}=$20 MeV (top) and 60 MeV (bottom). All results correspond to full CDCC calculations for the elastic channel, while the transfer channel has been calculated through Born approximation from all paths, and paths I, II, III and IV (see Fig. 30) for the black solid, red solid, blue solid, red dashed and blue dashed lines respectively. Figure taken from Gom17b , with authorization from APS.
In Fig. 32 we plot, as a function of incident energy, the ratio of the cross section at the peak for the calculation where one of the transfer paths was selected divided by the cross section at the peak for the full calculation, containing all transfer paths. As can be seen in the figure, path II, corresponding to ${\cal T}^{inel}_{dp}$ gains relevance as the energy increases,
Figure 32: Ratio between cross sections at the peak as a function of the deuteron incident energy. For each energy the ratio is between the calculation where one of the transfer paths is selected (as in Fig. 31) and the full calculation. As can be seen, path II, corresponding to the inelastic excitation of 10Be, gains importance with increasing energy so that at the higher energies, where the $R_{x}$ factors are smaller, it is of the same magnitude as path III, corresponding to deuteron breakup. Figure taken from Gom17b , with authorization from APS.
5 Inclusive breakup reactions
In previous sections, we have considered only exclusive breakup processes in which the three fragments ($b$, $x$ and $A$) survive after the collision and are observed in a definite internal state. In particular, when all fragments end up in their ground state, the process is called elastic breakup (EBU). In many experiments, however, one or more fragments are not detected in the final state. This is the case, for example, of reactions of the form $A(a,b)X$, in which only one of the projectile constituents ($b$ in this case) is observed. The angular and energy distributions of the $b$ fragments will contain contributions from all possible final states of the $x+A$ system. This includes the EBU channel, in which $x$ and $A$ remain in their ground state, but also $x$ transfer, breakup accompanied by excitations of $A$, and $x$+$A$ fusion [named incomplete fusion (ICF)]. These processes are schematically depicted in Fig. 33 for a deuteron-induced reaction. These non-elastic breakup components (NEB) must be added to the EBU component to give the total inclusive breakup. Whereas the EBU contribution can be accurately calculated within the CDCC method and other approaches, the evaluation of NEB is more involved due to the large number of accessible states.
Figure 33: Pictorial representation of elastic and non-elastic breakup channels for a $d+A$ reaction.
5.1 The Ichimura-Austern-Vincent (IAV) model
The problem of inclusive breakup was addressed in the 1980s by several groups, which developed formal techniques to reduce the sum over final states to a closed form HM85 ; Ich85 . These models are based on a spectator-participant picture. The $b$ fragment, which acts as spectator, is assumed to scatter elastically by the target nucleus and hence its relative motion with the target is represented by some optical potential $U_{bA}$. The fragment $x$, the participant, is allowed to interact in any possible way with the target. When $x$ scatters elastically by the target, we simply have the elastic breakup defined above. Conversely, NEB encompasses those processes in which $x$ scatters nonelastically by the target. Interestingly, all these early theories arrived at a similar expression for the NEB differential cross section with respect to the angle and energy of $b$ fragments, given by:
$\left.\frac{d^{2}\sigma}{dE_{b}d\Omega_{b}}\right|_{\mathrm{NEB}}=-\frac{2}{\hbar v_{a}}\rho_{b}(E_{b})\langle\varphi_{x}|W_{x}|\varphi_{x}\rangle,$ (66)
where $\rho_{b}(E_{b})=k_{b}\mu_{b}/[(2\pi)^{3}\hbar^{2}]$ is the density of states (with $\mu_{b}$ the reduced mass of $b+B$ and $k_{b}$ their relative wave number), $W_{x}$ is the imaginary part of the optical potential $U_{x}$, which describes $x+A$ elastic scattering. The function $\varphi_{x}$ ($x$-channel wavefunction) describes the final state of the $x-A$ system when the spectator $b$ is scattered with energy $E_{b}$ within the solid angle $\Omega_{b}$. A sketch with the relevant coordinates is shown in Fig. 34.
Figure 34: Relevant coordinates for a $A(a,b)X$ reaction.
The proposed models differ in the prescription for the $x$-channel wavefunction. Austern et al. Aus87 proposed a formula based on the CDCC wavefunction for the $d+A$ system,
$\varphi^{\mathrm{3b}}_{x}({\vec{r}}_{x})=\langle{\vec{r}}_{x}\chi_{b}^{(-)}|\Psi^{\mathrm{3b(+)}}\rangle.$ (67)
This expression has the appealing feature of connecting the NEB cross section with the CDCC three-body wavefunction. As noted in Sect. 3.7, the CDCC does not provide detailed NEB breakup cross sections by itself. However, with the aid of Eqs. (67) and (66) one can effectively extract the NEB content of the CDCC wavefunction.
Expression (67), albeit formally simple, possesses the difficulty of requiring an accurate representation of the three-body wavefunction $\Psi^{\mathrm{3b(+)}}$ in the full configuration space, since there is no natural cutoff in the integration variable ${\vec{r}}_{b}$. This difficulty can be sorted out recasting $\varphi^{\mathrm{3b}}_{x}({\vec{r}}_{x})$ as a solution of the inhomogeneous equation
$(E^{+}_{x}-K_{x}-U_{xA})\varphi^{\mathrm{3b}}_{x}=\langle{\vec{r}}_{x}\chi_{b}^{(-)}|V_{\mathrm{post}}|\Psi^{\mathrm{3b}^{(+)}}\rangle.$ (68)
Although this equation is a priori more difficult to implement than Eq. (67), it has the advantage that the $\Psi^{\mathrm{3b(+)}}$ function appears multiplied by $V_{\mathrm{post}}$, that will tend to emphasize small $b-x$ separations and hence one requires only an approximate three-body wavefunction accurate within that range. This can be achieved, for instance, expanding $\Psi^{\mathrm{3b(+)}}$ in terms of $b-x$ eigenstates, as done in the CDCC method, or in terms of Weinberg states Pan13 , already discussed in Sec. 4. The implementation of the method with CDCC wavefunctions is numerically challenging and the first calculation of this kind was reported only recently Jin19b .
Ichimura, Austern and Vincent Ich85 proposed a simpler DWBA version of the 3B formula above. In DWBA, the exact wavefunction $\Psi^{(+)}$ is approximated by the factorized form:
$|\Psi^{(+)}\rangle\approx|\chi_{a}^{(+)}\phi_{a}\phi^{0}_{A}\rangle.$ (69)
With this approximation, the NEB component of the $b$ singles cross section becomes
$\left.\frac{d^{2}\sigma}{dE_{b}d\Omega_{b}}\right|^{\mathrm{IAV,post}}_{\mathrm{NEB}}=-\frac{2}{\hbar v_{a}}\rho_{b}(E_{b})\langle\varphi^{\mathrm{post}}_{x}|W_{x}|\varphi^{\mathrm{post}}_{x}\rangle,$ (70)
with
$(E^{+}_{x}-K_{x}-{U}_{xA})\varphi^{\mathrm{post}}_{x}({\vec{r}}_{x})=\langle{\vec{r}}_{x}\,\chi_{b}^{(-)}|V_{\mathrm{post}}|\chi_{a}^{(+)}\phi_{a}\rangle\,.$ (71)
which can be also expressed in integral form as
$\varphi^{\mathrm{post}}_{x}({\vec{r}}_{x})=G^{(+)}_{x}\langle{\vec{r}}_{x}\,\chi_{b}^{(-)}|V_{\mathrm{post}}|\chi_{a}^{(+)}\phi_{a}\rangle.$ (72)
This latter form is more convenient for the application of Green’s function techniques Pot15 .
The IAV model has recently been revisited by several groups Car16 ; Jin15 ; Pot15 and its accuracy has been assessed against experimental data with considerable success Jin17 ; Pot17 .
As an example, in Fig. 35 we show the experimental and calculated energy distribution of $\alpha$ particles emitted in the reaction induced by 6Li on 118Sn at the incident energies indicated by the labels (adapted from Ref. Jin17 ). The EBU contribution (dashed line) was evaluated with the CDCC method, whereas the NEB part (dotted line) was obtained with the IAV method, in its DWBA form. Interestingly, one can see that the inclusive $\alpha$ yield is largely dominated by the NEB mechanism. The EBU is only important at small scattering angles (distant collisions, in a classical picture).
Figure 35: Energy distribution of $\alpha$ particles produced in the reaction 118Sn(6Li,$\alpha$)$X$ at the incident energies indicated by the labels. The dashed and dotted lines are the EBU and NEB contributions, and the solid line is their incoherent sum (TBU). Quoted from Jin17 , with authorization from APS.
Application to three-body projectiles
The model has also been applied to projectiles with a genuine three-body structure, such as Borromean nuclei, in which one of the fragments is removed from the projectile, leaving a residual two-body unbound system. An example is the reaction A(9Be,8Be)X which produces two alpha particles in the final state.
Using the simplifying assumption that the transition operator $V_{\mathrm{post}}$ does not alter the state of the unbound two-body core, one can use the IAV formalism with the new $x$-channel wavefunction [c.f. Eq. (72)]:
$\varphi_{x}(\vec{k}_{b},{\vec{r}}_{x})=G^{\mathrm{opt}(+)}_{xA}\langle\vec{r}_{x}\,\chi^{(-)}_{b}(\vec{k}_{b})|V_{\mathrm{post}}|\chi_{a}f_{ab}(\vec{r}_{bx})\rangle$ (73)
where we have introduced the overlap function between the three-body projectile and the two-body residual nucleus after the particle transfer, i.e.,:
[…]
The three-body wavefunction can be evaluated with the techniques discussed in previous sections, such as the hyperspherical expansion method employed in four-body CDCC calculations. An application is shown in Fig. 36, taken from Vil24 . The top panel (a) shows the norm of the $\langle{{}^{9}}\mathrm{Be}|^{8}\mathrm{Be}\rangle$ overlaps as a function of the $\alpha-\alpha$ relative energy. The middle panel is the cross section, as a function of $\alpha-\alpha$ relative energy, leading to bound states of the 198Au system for an incident energy of $E_{\mathrm{c.m.}}$. Finally, the bottom panel is the stripping cross section, integrated in the relative energy of the two $\alpha$’s, as a function of the incident energy. The separate contributions due to the 8Be $0^{+}$ and $2^{+}$ states are shown, together with their sum. The dashed line is the result of rescaling the IAV calculations by the spectroscopic factors obtained in the ab-initio Variational Monte Carlo (VMC). See Vil24 for further details.
Figure 36: IAV calculations for the reaction 197Au(9Be,8Be*)198Au populating bound states of the residual nucleus. (a) Norm of the overlaps between the 9Be g.s. and the 8Be continuum pseudostates for different relative angular momenta $L$, as a function of the $\alpha$-$\alpha$ relative energy up to 20 MeV. Dashed lines are included as a guide. (b) Differential cross section, at $E_{\mathrm{cm}}=36$ MeV, as a function of the $\alpha$-$\alpha$ relative energy in the final 8Be system, for 0+ states (solid line) and $2^{+}$ states (dashed line). The symbols indicate the nominal energies of the g.s. ($0^{+}$) and the first $2^{+}_{1}$ resonance of 8Be. (c) Angle-integrated total transfer cross section as a function of $E_{\mathrm{cm}}$. The contribution from the $0^{+}$ (dot-dashed line) and $2^{+}$ (dotted) states of 8Be, together with their sum (blue solid), are compared to the experimental data of Ref. gollan21 . The IAV result rescaled by VMC spectroscopic factors (dashed line) is also shown (see text). Adapted from Ref. Vil24 .
5.2 The Eikonal Hussein-McVoy formula (EHM)
Hussein and McVoy HM85 obtained an alternative formula for the inclusive breakup cross section. Although this was not the rationale followed by Hussein and McVoy, this formula can be readily derived making the same factorized approximation used by IAV, but using such approximation in Eq. (67):
$\varphi^{\mathrm{HM}}_{x}(\vec{r}_{x})=\langle{\vec{r}}_{x}|\varphi^{\mathrm{HM}}_{x}\rangle=\langle{\vec{r}}_{x}\chi^{(-)}_{b}|\chi_{a}^{(+)}\phi_{a}\rangle.$ (75)
This result is referred to as Hussein-McVoy (HM) formula for inclusive breakup (or, more precisely, for NEB). At sufficiently high energies, the HM formula can be simplified using the following three approximations:
  1. (i)
    the distorted waves $\chi^{(+)}_{a}(\vec{r}_{a})$ and $\chi^{(-)}_{b}(\vec{r}_{b})$ are treated in the Glauber (also known as eikonal) approximation. The Glauber approximation to an elastic-scattering distorted wave is:
    $\chi_{\vec{k}}^{(+)}({\vec{r}})=\mathrm{e}^{i\vec{k}\cdot\vec{r}}\exp\left[+i\int_{-\infty}^{z}\Delta k\left(z^{\prime},b\right)\mathrm{d}z^{\prime}\right]$ (76)
    where the incident momentum $\vec{k}$ points along the positive $z$-axis, and $\vec{b}$ is the component of $\vec{r}$ perpendicular to $z$. The exponent in the second factor is
    $\Delta k\left(z^{\prime},b\right)\equiv-\frac{k}{2E}U(z,b),$ (77)
    which, when integrated along the entire trajectory, gives the optical phase shift
    $2\delta(b)=\int_{-\infty}^{\infty}\Delta k\left(z^{\prime},b\right)\mathrm{d}z^{\prime}=2\int_{0}^{\infty}\Delta k\left(z^{\prime},b\right)\mathrm{d}z^{\prime},$ (78)
    that is related to the partial-wave optical S-matrix, i.e., $S(b)=\exp[{2i\delta(b)}]$.
  2. (ii)
    The $U_{a}$ potential, distorting the incident wave, is taken as the sum of the corresponding fragment-target potentials:
    $U_{a}=U_{bA}+U_{xA}$ (79)
    This particular choice has the virtue of taking into account breakup effects in the entrance channel. However, since each potential in (79) is evaluated in the corresponding fragment-target coordinate, the associated initial state wavefunction of the system would be a solution of a complicated three-body equation.
  3. (iii)
    The above complication vanishes thanks to the use of the eikonal approximation: the $x$ and $b$ fragments move with the same average velocity as the projectile and hence their momenta are given by
    $\vec{k}_{x}=\left(m_{x}/m_{\mathrm{a}}\right)\vec{k}_{\mathrm{a}},\quad\vec{k}_{b}=\left(m_{b}/m_{a}\right)\vec{k}_{a}.$ (80)
With the particular choice (79) and the assumption (80) one obtains the following result for the $x$-channel wavefunction:
$\varphi^{\mathrm{EHM}}_{x}({\vec{r}}_{x})=$ $\int\mathrm{d}^{3}{\vec{r}}_{b}\chi_{b}^{(-)*}({\vec{r}}_{b})\chi_{a}^{(+)}({\vec{r}}_{a})\phi_{a}({\vec{r}}_{bx})$ (81)
$=$ $\mathrm{e}^{i\vec{k}_{x}\cdot{\vec{r}}_{x}}\exp\left[i\int_{-\infty}^{z_{x}}\Delta k_{x}\left(z^{\prime},b_{x}\right)\mathrm{d}z^{\prime}\right]$
×$\times\int\mathrm{d}^{3}{\vec{r}}_{b}\mathrm{e}^{i\vec{q}\cdot{\vec{r}}_{b}}S_{bA}\left(b_{b}\right)\phi_{\mathrm{a}}\left({\vec{r}}_{bx}\right)$
with $\vec{q}=\vec{k}_{b}-\vec{k}^{\prime}_{b}$, the average momentum transferred in $b-A$ elastic scattering.
Inserting this expression into the general expression (70) (see details in HM85 ; Mar02 ) one obtains for the double differential cross section:
$\left.\frac{d^{2}\sigma}{dE_{b}d\Omega_{b}}\right|^{\mathrm{EHM}}_{\mathrm{NEB}}$ $=\frac{2}{\hbar v_{a}}\rho_{b}(E_{b})\frac{E_{x}}{k_{x}}$
×$\times\int\mathrm{d}^{2}\vec{b}_{x}\left|\tilde{\phi}_{a,b}(\vec{q},\vec{b}_{x})\right|^{2}\left[1-\left|S_{xA}(b_{x})\right|^{2}\right].$ (82)
It should be noticed that the NEB depends only on the asymptotic properties, this is, the $S$ matrices, of the interaction of $b$ and $x$ with the target. There is no sensitivity on the wavefunctions in the interaction region. This is a result of the eikonal approximation, plus the particular choice of the distorted interaction, which included the imaginary potential $W_{xA}$ which ultimately generates the NEB.
In many applications, one is interested in the total yield of fragment $b$, which is obtained upon integration of the previous formula over the angular and energy variables, resulting:
$\sigma_{\mathrm{NEB}}^{\mathrm{EHM}}=$ $\frac{2}{v_{\mathrm{a}}}(2\pi)^{3}\frac{E_{x}}{\hbar k_{x}}\int\mathrm{d}^{3}\vec{r}_{b}\mathrm{d}^{3}\vec{r}_{x}\left|\phi_{a}\left(\vec{r}_{bx}\right)\right|^{2}$
$\times$ $\left|S_{bA}(b_{b})\right|^{2}\left[1-\left|S_{xA}\left(b_{x}\right)\right|^{2}\right].$ (83)
This equation has an appealing and intuitive form: the integrand contains the product of the probabilities for the core being elastically scattered by the target, $|S_{bA}(b_{b})|^{2}$, times the probability of the valence particle being absorbed, $(1-|S_{xA}(b_{x})|)^{2}$. These probabilities are weighted by the projectile wave function squared and integrated over all possible impact parameters. Due to the Glauber approximation, Eq. (5.2) is expected to be accurate at high energies (above $\sim$100 MeV per nucleon). In fact, this formula has been extensively employed in the analysis of intermediate-energy knockout reactions (see e.g. Han03 ; Tos01 ; Tos14 and references therein) mostly aimed at obtaining spectroscopic information of nucleon hole states.
5.3 Interpretation of inclusive breakup data
In the previous sections, we have considered the elastic breakup and nonelastic breakup (with the latter possibly including transfer and incomplete fusion) taking place in a breakup reaction. In actual inclusive breakup experiments these contributions will appear entangled in the data, although they produce some distinctive features. For example, the energy distribution of the observed fragment at a given scattering angle will exhibit a characteristic shape, as shown in Fig. 37 for a hypothetical $A(d,pX)$ reaction. To understand this spectrum, it is important to recall that a given proton energy and angle will univocally determine the excitation energy of the residual $n+A$ system. According to the proton energy (or, equivalently, the residual system excitation energy), we may distinguish the following regions:
Figure 37: Proton energy spectrum from a $A(d,pX)$ inclusive breakup reaction.
  1. (i)
    The highest proton energies will be characterized by some narrow peaks corresponding to bound states of the $n+A$ system. Depending on the experimental resolution, these peaks will appear separated or, instead, will merge with neighbouring peaks. Theoretically, the cross section for these isolated bound states can be evaluated with the standard formalisms for transfer reactions, such as DWBA, CCBA, ADWA or CRC.
  2. (ii)
    As the excitation energy of the residual nucleus increases, so does the density of states and the low-lying bound states will merge to form a quasi-continuum. Here, the treatment with the aforementioned methods becomes more troublesome because of the impossibility of disentangling unambiguously the contribution of each state from the data. When one is interested in energy averaged cross sections, the combination of the IAV model of Sec 5.1 with the dispersive models provides an appealing alternative to the traditional methods.
  3. (iii)
    At a certain excitation energy, the residual system will reach the neutron separation threshold ($E_{x}=S_{n}$). Just above this excitation energy, the proton spectrum will exhibit narrow peaks, corresponding to low-lying resonances, superimposed to a nonresonant continuum background. It is important to realize that the properties of the system just above the threshold ($E_{x}>S_{n}$) and just below it ($E_{x}<S_{n}$) states are qualitatively similar. In particular, the discrete energy levels extend above $S_{n}$ even if those states are not strictly stationary. Therefore, the wavefunctions of the narrow resonances resemble very much those for the bound states below threshold and correspond to situations in which the system remains bound for a long time before decaying by barrier penetration. This is important from the reaction point of view because it indicates that the transfer cross section should evolve smoothly from negative to positive energies, as it actually happens in the distribution shown in Fig. 37
  4. (iv)
    As the excitation energy of the residual nucleus increases above the threshold, the narrow resonances disappear, the spectrum becomes continuous and structureless, giving rise to a bell-shaped bump. The most probable energy of the emerging proton can be obtained by assuming that the proton gets half of the kinetic of the deuteron at the point of breakup, plus the Coulomb energy of the deuteron at that place, and half the internal energy of the deuteron ($E_{B}=-2.22$ MeV) (see Ref. Bla91 , p.509), i.e.,
    $\langle E_{p}\rangle\simeq\frac{1}{2}\left(E_{d}-\frac{Ze^{2}}{R_{\text{bu}}}\right)+\frac{Ze^{2}}{R_{\text{bu}}}-\frac{1}{2}E_{B}$ (84)
    where $E_{d}$ is the kinetic energy of the relative motion of the deuteron and the target nucleus A in the c.m. frame and $R_{bu}$ the deuteron-target separation at the time of the deuteron breakup. The bump will contain contributions coming from both EBU and NEB components discussed in previous sections. The EBU can be conveniently evaluated with the DWBA or CDCC methods whereas, for the NEB part, the IAV model provides a very convenient framework.
  5. (v)
    In the former contributions, we have implicitly assumed that the observed particle (proton in this case) scatters elastically by the target nucleus. There will be situations in which this this not the case; for example, when the projectile fuses completely with the target nucleus forming a compound nucleus that will eventually thermalize by emitting particles and gamma rays. Among these particles, there will be protons that will add up incoherently to the proton spectrum. These protons are typically emitted with low energy in the c.m. frame and will therefore contribute to the low energy part of the proton spectrum. Some phenomena related to the fusion of weakly-bound nuclei will be discussed in the next section.
6 Fusion involving weakly bound nuclei
The previous sections have been devoted to the modeling of direct nuclear reactions. Fusion between weakly-bound nuclei display also distinctive features which require adequate reaction formalisms and considerations with respect to the case of well-bound nuclei. The topic of fusion with both stable and unstable nuclei has motivated many works and excellent review papers have been published in recent years, so we refer the reader to these works for a detailed account of the present status of the description of fusion (e.g. Can06 ; Can15 ; Can21 ). Here, we focus on two phenomena which are the focus of many theoretical and experimental efforts by several groups. One is the phenomenon of complete fusion suppression. Complete fusion is conveniently defined as the process in which the whole charge of the projectile and target nuclei merge, giving rise to an excited compound nucleus that will subsequently decay by particle and/or gamma emission. The other phenomenon is the large observed yields compatible with the partial fusion of the projectile (incomplete fusion, ICF). Experimental results and theoretical calculations indicate that these two phenomena are actually related since they appear simultaneously and so a plausible explanation of the CF suppression might be in fact the leak of flux going to the ICF channels.
A variety of models have been proposed to evaluate the CF cross section, from the simple single-barrier penetration model to more sophisticated coupled-channels methods, in which collective excitations of the projectile and/or target nucleus are taken into account explicitly. These models are very successful at predicting the cross section for well bound nuclei, but tend to overestimate them for weakly bound projectiles at energies above the barrier. For example, for the light weakly bound nuclei 6,7,8Li, 9Be the experimental CF cross sections are found to be suppressed by $\sim$20-30compared to the case of tightly bound nuclei Das99 ; Tri02 ; Das02 ; Das04 ; Muk06 ; Rat09 ; LFC15 .
Early analyses of these experiments tried to explain the phenomenon using coupled-channels calculations, including the coupling to low-lying excited states of the projectile and target Das99 ; Das02 ; Kum12 ; Zha14 ; Fan15 . Yet, these calculations systematically failed to reproduce such suppression. This failure has been attributed to the omission in these calculations of the breakup of the projectile; a scenario was suggested in which the weakly bound projectile breaks up prior to reaching the fusion barrier, with the subsequent reduction of probability of complete capture. This interpretation is supported by the presence of large $\alpha$ yields as well as target-like residues which are consistent with the capture of one of the fragment constituents of the projectile, that is, ICF. To account for these observations, some authors have proposed a two-step scenario Das02 ; Dia07 consisting on the elastic dissociation of the projectile followed by the capture of one of the fragments by the target. However, calculations based on a three-dimensional classical dynamical model Dia07 , which incorporates this two-step breakup-fusion mechanism, can only explain a small fraction of the observed CF suppression for 9Be Coo16 and 8Li Coo18 reactions. More encouraging results have been obtained with different methods based on the CDCC formalism, as described in the following subsections.
6.1 Computation of CF and ICF with CDCC
Although the CDCC method was originally envisaged as a practical tool to evaluate the elastic and breakup observables, some works have been made to use this method to obtain complete and incomplete fusion cross sections.
Some authors Hag00 ; Dia02 have proposed to identify the total fusion with the amount of flux that leaves the coupled channels set due to a short-range imaginary potential $iW_{F}(R)$, while the CF is identified with the absorption due to such potential, but restricted to bound states of the projectile only. Thus, the total fusion is computed as
$\sigma_{TF}=\frac{\pi}{2\mu E}\sum_{J_{T}}(2J_{T}+1)P_{J_{T}},$ (85)
where $E$ is the bombarding energy and $P_{J_{T}}$ is the complete fusion probability for total angular momentum $J_{T}$
[…]
where $\chi^{J_{T}}_{\beta,\beta_{i}}$ are the radial solutions obtained from the coupled equations (13). In the case of CF, the expression for the cross section is identical but the sum in $\beta$ is restricted to those channels associated to bound states.
The CF obtained in this way represents a lower limit of the physical CF cross section, since one has assumed no capture of all projectile fragments (10Be and the halo neutron) from breakup channels. In reality, these events should contribute to the CF, but cannot be distinguished in this model from the capture of only one projectile fragment. In this model, the incomplete fusion $\sigma_{ICF}$ is therefore defined as the absorption from breakup channels.
Figure 38: Calculated total fusion cross sections for 11Be + 208Pb (full stars) compared with the experimental data from Sig98 for 11Be + 209Bi (full squares) using the CDCC wavefunction. Adapted from Ref. Dia02 , with permission from APS.
An application of this method is shown in Fig. 38, where the calculations for the total fusion of 11Be + 208Pb (full stars) are compared with the total fusion data for the nearby reaction 11Be + 209Bi from Ref. Sig98 . The agreement is reasonable around the Coulomb barrier, but the calculation underestimates the data by $\sim$ 41% for energies well above the Coulomb barrier.
A limitation of this method is that it can only be applied to projectiles composed of a heavy charged fragment and a light uncharged one (such as 11 Be), since it relies on the assumption that the center of mass of the projectile is close to that of the heavy fragment and far from the light one. Thus, it cannot be used for projectiles like 6,7Li that break up into two fragments of comparable masses. To overcome this difficulty, Hashimoto et al. Hash09 proposed an alternative approach based also on the CDCC method and applied it to the case of deuteron scattering. Their idea is to transform the CDCC wavefunction from its natural coordinates ${\vec{r}}$, ${\vec{R}}$} to the coordinates ${\vec{r}}_{p}$, ${\vec{r}}_{n}$},
$|\Psi({\vec{r}},{\vec{R}})|^{2}\,d{\vec{r}}d{\vec{R}}=|\widetilde{\Psi}({\vec{r}}_{p},{\vec{r}}_{n})|^{2}d{\vec{r}}_{p}d{\vec{r}}_{n}\,,$ (87)
and then associate the CF and ICF cross sections with the absorption taking place in different regions of the $\{r_{p},r_{n}\}$ space, as illustrated in Fig. 39. The distances $r^{\mathrm{ab}}_{p}$ and $r^{\mathrm{ab}}_{n}$ denote the absorption radii for the proton and neutron, such that for $r_{p}>r^{\mathrm{ab}}_{p}$ ($r_{n}>r^{\mathrm{ab}}_{n}$) the proton (neutron) absorption becomes negligible. The CF is identified with the absorption taking place when both the proton and neutron are inside their respective absorption radii, i.e.,
[…]
×$\times\{W_{p}({\vec{r}}_{p})+W_{n}({\vec{r}}_{n})\},$ (88)
where $K$ is the incident wave number, $E$ the energy and $N$ is a normalization constant. Similarly, the ICF cross section is obtained from the absorption occurring in the region where one of the two fragments is inside its absorption radius while the other is outside it, i.e.,
[…]
Figure 39: Schematic illustration of the four integration regions employed in the method of Hashimoto et al. Hash09 for evaluating complete and incomplete fusion from CDCC wavefunction Quoted with authorization from Oxford University Press.
Figure 40: a) Comparison of $\sigma_{\rm ICF}^{(p)}$ (short-dashed line) and $\sigma_{\rm ICF}^{(n)}$ (dash-dotted line) obtained in Hash09 with $\sigma_{\rm STR}^{(p)}$ (squares) and $\sigma_{\rm STR}^{(n)}$ (triangles) given by the Glauber model Ye09  b) Complete fusion cross sections calculated with the Glauber model (dots) and with the method of Hash09 (solid line). Quoted with authorization from Oxford University Press.
The method was applied to the reaction $d$+7Li Hash09 . Although no comparison with experimental data was attempted, the authors compared their results with those computed with the Glauber calculations of Ref. Ye09 . This comparison is shown in Fig. 40 with the upper and lower panels corresponding respectively to the ICF and CF cross sections as a function of the deuteron incident energy. The agreement for the ICF is found to be very satisfactory for energies as low as $E_{d}$=10 MeV. Since the Glauber model is a high-energy approximation, the agreement at these relatively low energies is somewhat unexpected. By contrast, for the CF part (bottom panel), the Glauber method is found to give significantly smaller cross sections. A discussion of these results is provided in Ye09 .
Another procedure to extract the CF and ICF cross section from the CDCC method was proposed by Parkar and co-workers Par16 . The central idea of their method is to perform a series of CDCC calculations with different choices of the fragment-target potentials. In particular, to obtain the ICF cross section for the capture of a given fragment, they perform a CDCC calculation using a short-range imaginary potential for the interaction of that fragment and the target, and a real potential for the other fragment. Using this method, these authors have been able to obtain a reasonable account of the CF, ICF and TF cross sections of reactions induced by 6,7Li projectiles Par16 ; Par18 , although the short-range fusion potential needs to be adjusted for each reaction.
More recently, the authors of Refs. Ran20 ; Cor20 have proposed an alternative method which requires a single CDCC calculation to compute the TF, CF and ICF cross sections. In this CDCC calculation, the fragment-target interactions are also modeled with optical potentials with a short-ranged imaginary part. For a two-body projectile, the total fusion cross section is computed as:
$\sigma_{\mathrm{TF}}=\frac{1}{|N|^{2}}\ \frac{K}{E}\ \ \left\langle\ \Psi^{(+)}\,\left|\,W^{(1)}+W^{(2)}\,\right|\Psi^{(+)}\ \right\rangle\,,$ (91)
where $W^{(1,2)}$ represent the imaginary parts of the fragment-target interactions. This CDCC wavefunction can be split into bound ($\Psi^{B}$) and continuum ($\Psi^{C}$) components:
${\Psi}^{(+)}({{\vec{R}}},{\vec{r}})={\rm\Psi}^{B}({{\vec{R}}},{{\vec{r}}})+{\Psi}^{C}({{\vec{R}}},{{\vec{r}}}),$ (92)
where ${\Psi}^{B}$ and ${\Psi}^{C}$ are given by the channel expansions
${\Psi}^{B}({{\vec{R}}},{{\vec{r}}})$ $=$ $\sum_{\beta\,\in\,B}\ \chi_{\beta}({{\vec{R}}})\phi_{\beta}({{\vec{r}}})$ (93)
${\Psi}^{C}({{\vec{R}}},{\bf r})$ $=$ $\sum_{\gamma\,\in\,C}\ \chi_{\gamma}({{\vec{R}}})\phi_{\gamma}({{\vec{r}}}),$ (94)
where $\phi_{\beta}$ and $\phi_{\gamma}$ are respectively bound and unbound states of the projectile, and $\chi_{\beta}$ and $\chi_{\gamma}$ are the corresponding wave function describing the projectile-target relative motion.
Assuming that matrix-elements of the imaginary potentials connecting bound channels to bins are negligible, Eq. (91) can be put in the form
$\sigma_{\mathrm{TF}}=\sigma_{\mathrm{TF}}^{B}\ +\ \sigma_{\mathrm{TF}}^{C},$ (95)
with
$\sigma_{\mathrm{TF}}^{B}$ $=$ $\frac{1}{|N|^{2}}\frac{K}{E}\sum_{\beta,\beta^{\prime}\,\in\,{B}}\ \left\langle\chi_{\beta}\left|W_{\beta\beta^{\prime}}^{(1)}+W_{\beta\beta^{\prime}}^{(2)}\,\right|\chi_{\beta^{\prime}}\right\rangle$ (96)
$\sigma_{TF}^{C}$ $=$ $\frac{1}{|N|^{2}}\frac{K}{E}\sum_{\gamma,\gamma^{\prime}\in\,{C}}\ \left\langle\chi_{\gamma}\left|W_{\gamma\gamma^{\prime}}^{(1)}+W_{\gamma\gamma^{\prime}}^{(2)}\,\right|\chi_{\gamma^{\prime}}\right\rangle.$ (97)
where $W_{\alpha\alpha^{\prime}}^{(i)}=\left(\phi_{\alpha}\left|{W}^{(i)}\right|\phi_{\alpha^{\prime}}\right),$ with $\alpha,\alpha^{\prime}$ standing for either $\beta,\beta^{\prime}$ or $\gamma,\gamma^{\prime}$, are the matrix-elements of the imaginary potentials.
Then, by performing an angular momentum expansion of the wave functions and the imaginary potentials, Eqs. (96) and (97) become
$\sigma_{TF}^{B}$ $=$ $\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}_{B}^{TF}(J_{T})$ (98)
$\sigma_{TF}^{C}$ $=$ $\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}_{C}^{TF}(J_{T}),$ (99)
with
$\mathcal{P}_{B}^{TF}(J_{T})$ $=$ $\mathcal{P}_{B}^{(1)}(J_{T})+\mathcal{P}_{B}^{(2)}(J_{T})$ (100)
$\mathcal{P}_{C}^{TF}(J_{T})$ $=$ $\mathcal{P}_{C}^{(1)}(J_{T})+\mathcal{P}_{C}^{(2)}(J_{T}).$ (101)
where $\mathcal{P}_{B}^{(i)}(J_{T})$ and $\mathcal{P}_{C}^{(i)}(J_{T})$ are the probabilities of absorption of fragment $c_{i}$ in bound channels and in the continuum, respectively, resulting from the contributions of ${W}^{(i)}$ to the TF cross section.
In terms of these probabilities, the authors of Ref. Ran20 ; Cor20 introduce the ICF probabilities
[…]
and the sequential complete fusion probability
${\mathcal{P}}^{\mathrm{SCF}}(J_{T})={\mathcal{P}}^{(1)}_{C}(J_{T})\times{\mathcal{P}}^{(2)}_{C}(J_{T}).$ (104)
In terms of the introduced probabilities, the following fusion cross sections are defined:
  • Direct complete fusion (CF):
    $\sigma_{DCF}=\sigma_{TF}^{B},$ (105)
    which describes the simultaneous capture of the two fragments.
  • Sequential complete fusion (SCF):
    $\sigma_{SCF}=\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}^{SCF}(J_{T}).$ (106)
  • ICF of fragment $c_{i}$ (ICFi)
    $\sigma_{ICFi}=\frac{\pi}{K^{2}}\,\sum_{J_{T}}(2J_{T}+1)\ \mathcal{P}^{ICFi}(J_{T}).$ (107)
In this formalism, the CF, ICF and TF cross sections are then given by
$\sigma_{\mathrm{CF}}$ $=$ $\sigma_{DCF}\,+\,\sigma_{SCF},$ (108)
$\sigma_{\mathrm{ICF}}$ $=$ $\sigma_{ICF1}\,+\,\sigma_{ICF2},$ (109)
$\sigma_{TF}$ $=$ $\sigma_{CF}+\sigma_{ICF}.$ (110)
An application of this model is shown in Fig. 41, where CF (left panel) and ICF (right panel) data for the reaction 7Li+209Bi Das02 ; Das04 are compared with the predictions of the model. The upper and lower panels display the same results in logarithmic and linear scale, respectively. For the CF, the agreement with the data is very satisfactory, both above and below the barrier (indicated by the arrow). For the ICF, the separate contributions for triton capture and $\alpha$ capture are shown, with the former giving a much larger cross section. The calculations are found to reproduce well the data up to $E_{\mathrm{c.m.}}\approx 34$ MeV, but overestimate them at higher energies. A detailed discussion of these results can be found in Ref. Cor20 .
Figure 41: CF (left) and ICF (right) cross sections for 7Li+209Bi. Experimental data from Refs. Das02 ; Das04 are compared with the calculations of Ref. Cor20 . With permission of APS.
6.2 Evaluation of CF and ICF cross sections with the IAV model
As discussed in Sec. 5, the IAV model provides the total inclusive cross section corresponding to the detection of the $b$ fragment in reactions of the form $A(a,b)X$. This results from the fact that the imaginary part that appears in the expectation value of Eq. (70) accounts in principle for all processes in which the participant fragment $x$ interacts nonelastically with the target nucleus. This will include the ICF cross section, but also other NEB processes not associated with the formation of a compound nucleus of the $x+A$ system, such as target excitation. The isolation of the ICF cross section from the total NEB cross section is indeed not a non-trivial problem. An intuitive approach consists in identifying the ICF with the absorption due to a short-ranged imaginary potential. This strategy has been however barely explored so far within the IAV model.
A tentative application of this idea is shown in the bottom panel of Fig. 42, quoted from Mor22 , corresponding to the reaction 7Li+209Bi at energies below and above the barrier. The symbols correspond to the ICF data of Dasgupta et al. Das02 ; Das04 . For the calculations we show the individual contributions to the ICF cross section, namely, $\alpha$-ICF (i.e., $\alpha$ absorbed) and $t$-ICF ($t$-absorbed) as well as their sum. To compute the $\alpha$-ICF ($t$-ICF), the imaginary part of the $\alpha$+209Bi ($t$+209Bi) system was replaced by a short-range imaginary potential of Woods-Saxon form and parameters $W_{0}=-50$ MeV, $r_{i}=1.0$ fm, $a=0.2$ fm. The results are very similar to those reported in Ref. Cor20 and shown in Fig. 41, in which the authors made use of the absorption and survival probabilities extracted from the CDCC calculations. Further calculations are needed to elucidate the usefulness and applicability of the IAV model to evaluate ICF cross sections.
Figure 42: CF (top) and ICF (bottom) cross sections for 7Li+209Bi. Experimental data from Dasgupta et al. Das02 ; Das04 are compared with the calculations based on the IAV model. In the CF plot, we include also the reaction cross section (from a CDCC calculation) and the fusion computed with the barrier penetration model (BPM). Top panel adapted from Jin19b . Bottom panel quoted from Mor22 .
The IAV model has also been used to infer CF cross sections of weakly bound nuclei Jin19a ; Jin19b . The idea of this method is to decompose the reaction cross section as follows
$\sigma_{R}\approx\sigma_{\mathrm{CF}}+\sigma_{\mathrm{inel}}+\sigma_{\mathrm{EBU}}+\sigma^{(b)}_{\mathrm{NEB}}+\sigma^{(x)}_{\mathrm{NEB}}.$ (111)
In this expression, $\sigma_{\mathrm{inel}}$ corresponds to the excitation of the projectile and/or target without dissociation (i.e., inelastic scattering). The terms $\sigma_{\mathrm{EBU}}$ and $\sigma^{(b,x)}_{\mathrm{NEB}}$ correspond to the EBU and NEB contributions. In the latter, one distinguishes the cases in which either the fragment $x$ or $b$ interacts nonelastically with the target whereas the other scatters elastically. A successful determination of the CF cross section from the decomposition (111) requires that all other quantities involved in this formula can be evaluated accurately. The pure inelastic scattering cross sections ($\sigma_{\mathrm{inel}}$) are standardly computed by means of coupled-channels calculations including low-lying collective excitations of the projectile and target. The EBU part can be accurately calculated with the continuum-discretized coupled-channels (CDCC) method. Finally, the NEB contributions can be evaluated with the IAV model, at least for projectiles with a developed two-boy structure, such as 6,7Li.
An application of the method to the 7Li+209Bi reaction is shown in the top panel of Fig. 42, adapted from Ref. Jin19b . The circles are the CF data from Ref. Das04 , the solid green line is the reaction cross section obtained from a CDCC calculation and the solid red line the calculated CF cross section inferred from Eq. (111) assuming a two-body model ($\alpha+t$) for 7Li. For comparison, a single-channel barrier penetration model (BPM) calculation (dashed lined) is shown. As can be seen, the data are largely suppressed with respect to the BPM. In contrast, the CF extracted from Eq. (111) explains very well the data. As discussed in Jin19a , the reduction with respect to the BPM is found to be mainly due to the competition due to the 209Bi(7Li, $\alpha$)X channel, which includes, among others, the $t$-ICF channel (see the bottom panel).
6.3 Application to surrogate reactions
The gradual improvement in the models oriented to the computation of ICF cross sections has driven their applicability of the so-called surrogate method (SRM). This method provides an indirect way of evaluating compound-nucleus cross sections in reactions for which the direct measurement is difficult or even not possible. An example is the extraction of neutron-induced cross sections of the form $(n,\chi)$, where $\chi$ is a given decay channel product ($\gamma$, fission fragment, etc). Following the Bohr hypothesis, it is customarily assumed that in these reactions the formation and decay of a compound nucleus take place independently of each other. To obtain information on the decay of the compound nucleus ($B^{*}$) occurring in the reaction of interest ($a+A\to B^{*}\to c+C$), one uses the alternative (surrogate) reaction $a+A\to b+B^{*}$ that involves a projectile-target combination ($a+A$) that is experimentally more accessible.
We consider as an example the extraction of $(n,\chi)$ cross sections from a $(d,pX)$ surrogate reaction. Compound nuclear reactions are properly described in the Hauser-Feshbach formalism, which considers the conservation of angular momentum $J$ and parity $\pi$. The cross section for the “desired” reaction $A(n,\chi)$ is given by
$\sigma_{(n,\chi)}(E_{n})$ $=$ $\sum_{J_{T},\pi}\sigma^{CN}(E_{ex},J_{T},\pi)\;\;G_{\chi}^{CN}(E_{ex},J_{T},\pi)\;,$ (112)
where $\sigma_{J_{T},\pi}^{CN}(E_{ex},J_{T},\pi)$ is the cross section for the CN formation and $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ the branching ratio for the decay in the channel $\chi$. Note that the excitation energy $E_{ex}$ of the compound nucleus $B^{*}$ is related to the center-of-mass energy $E_{a}$ in the entrance channel through the energy needed to separate $a$ from $B$: $E_{a}=E-S_{a}(B)$. The objective of the surrogate method is to experimentally determine or constrain the decay probabilities $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$, which are often difficult to calculate accurately.
In the surrogate reaction, the same CN nucleus $B^{*}$ is formed and the decay product of interest ($\chi$) is measured in coincidence with the outgoing particle $b$. The probability for this process can be written as:
$P_{S,\chi}(E_{ex})=\sum_{J_{T},\pi}F_{S}^{CN}(E_{ex},J_{T},\pi)\;G_{\chi}^{CN}(E_{ex},J_{T},\pi)\;,$ (113)
where the subscript $S$ denotes the specific surrogate reaction (in this case, $A(d,p)B^{*}$), $F_{S}^{CN}(E_{ex},J_{T},\pi)$ is the probability of forming $B^{*}$ in this surrogate reaction (with specific values of $E_{ex}$, $J$ and $\pi$) and $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ is the same branching ratio appearing in the direct reaction (112). The probability $P_{S,\chi}(E_{ex})$ can be obtained experimentally as the ratio between the number of coincidences between the $b$ particle and the decay particle $\chi$, $N_{S,\chi}$, and the total number of surrogate events, $N_{S}$, i.e.:
$P^{exp}_{S,\chi}(E_{ex})=\frac{N_{S,\chi}}{N_{S}\,\epsilon_{\chi}}\;,$ (114)
where $\epsilon_{\chi}$ is the efficiency of detecting the exit-channel $\chi$ for the reactions in which $b$ is detected.
Ideally, if a reliable prediction of $F_{S}^{CN}(E_{ex},J_{T},\pi)$ is possible, with an accurate determination of $P^{exp}_{S,\chi}(E_{ex})$ for a range of energies and angles of $b$, it might be possible to extract the branching ratios $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ which can then be used to calculate the desired cross section using (112). In practice, this approach is not always feasible due to the lack of some of this required information and the approach has relied on additional approximations. In particular, most practical applications have made use of the so-called “Weisskopf-Ewing approximation” (WEA) , which assumes that the branching ratios $G_{\chi}^{CN}(E_{ex},J_{T},\pi)$ are independent of the angular momentum and spin, giving rise to the simplified cross section:
$\sigma_{(n,\chi)}(E_{n})$ $=$ $\sigma^{CN}(E_{ex})\;G_{\chi}^{CN}(E_{ex})\;,$ (115)
where $\sigma^{CN}(E_{ex})$ is the CN cross section summed over all possible $J_{T},\pi$ values. Applying the same approximation to the surrogate reaction, and using $\sum_{J_{T},\pi}F_{S}^{CN}(E_{ex},J_{T},\pi)=1$ we have
$P_{S,\chi}(E_{ex})=G_{\chi}^{CN}(E_{ex})\;,$ (116)
allowing the determination of the desired cross section as
$\sigma_{(n,\chi)}(E_{n})$ $=$ $\sigma^{CN}(E_{ex})\;P_{S,\chi}(E_{ex})\;,$ (117)
which avoids the need of the probabilities $F_{S}^{CN}(E_{ex},J_{T},\pi)$.
So, under the validity of the WEA, the neutron $(n,\chi)$ cross section can be readily inferred from the measured probabilities for the surrogate reaction. Note also that the CN cross section must be estimated in some way, using, for example, an optical model calculation.
In practice, the validity of the WEA is not guaranteed in most cases so one needs to resort to the more general expressions (112) and (113). The probabilities $F_{S}^{CN}(E_{ex},J_{T},\pi)$ that appear in the latter can be estimated with the IAV model discussed in Sec. 5.1. This idea has been successfully applied to the 95Mo$(n,\gamma)$ reaction Rat19 . The direct $(n,\gamma$) cross section for this reaction are compared in Fig. 43 with the cross section extracted from the surrogate reaction 95Mo(d,p) using the aforementioned formalism. As can be seen, they are in excellent agreement. As also shown in this figure, the result using the WEA departs significantly from the direct measurement.
Figure 43: Application of the surrogate method to the 95Mo$(n,\gamma)$ reaction. Top (panel (A)): Calculations of the cumulative probability of forming the CN through 95Mo$(d,p)$ ($\sum F^{CN}_{{}^{95}Mo(d,p)}$). The shaded region, which spans excitation energies in the range from $E_{x}=8.55$ MeV to $E_{x}=10.65$ MeV, indicates the excitation energies over which the surrogate data are fit. The vertical dotted line represents the neutron evaporation threshold. The middle (panel (B)) histogram of the total contribution to CN formation over the shaded range in (A) as a function of angular momentum, decomposed into positive and negative parities, and normalized to one over the integration region. Bottom: The solid blue region is the $(n,\gamma)$ cross section obtained from the SRM, the red circles and black squares are the direct measurements Abr08 ; Mus76 . The uncertainty resulting from experimental data and fitting error is indicated by the shaded band. The result obtained using the Weisskopf-Ewing approximation is also shown (gold diamonds). Quoted from Ref. Rat19 , with permission from APS.
7 Semiclassical description of breakup and transfer reactions
[…]
7.1 The semiclassical formalism of Alder and Winther
In its simplest form, the theory of Alder and Winther Ald75 assumes that the projectile moves along a classical trajectory, which is weakly affected by the internal excitations of the colliding nuclei. This means that:
$\frac{\Delta\ell}{\ell}\ll 1~{}~{}\textrm{and}~{}~{}\frac{\Delta\varepsilon_{n}}{E}\ll 1.$
The projectile–target interaction is split as: $V(\vec{R},\xi)=V_{0}(\vec{R})+V_{\mathrm{coup}}(\vec{R},\xi)$, where $V_{0}(\vec{R})$ is independent of the internal coordinates and determines the classical trajectory $\vec{R}(t)$. The time evolution of the total wave function of the system verifies the Schrödinger equation
$i\hbar{d\Psi(\xi,\theta,t)\over dt}=\left[V({\vec{R}}(\theta,t),\xi)+H_{p}(\xi)\right]\Psi(\xi,\theta,t)$ (118)
subject to the initial condition $|\Psi(-\infty)\rangle=|0\rangle$.
In the spirit of the coupled-channels method, the total wave function is expanded in a basis of internal states of the projectile Hamiltonian $[H_{p}(\xi)-\varepsilon_{n}]\phi_{n}(\xi)$=0 :
$\Psi(\xi,\theta,t)=\sum_{n=0}c_{n}(\theta,t)e^{-i\varepsilon_{n}t/\hbar}\phi_{n}(\xi)$ (119)
which, when inserted in (118), gives rise to a set of coupled equations for the expansion coefficients:
$i\hbar{dc_{n}(\theta,t)\over dt}=\sum_{m}e^{-i(\varepsilon_{m}-\varepsilon_{n})t/\hbar}V_{nm}(\theta,t)c_{m}(\theta,t)$ (120)
with the initial condition $c_{n}(\theta,-\infty)=\delta_{n0}$. The time-dependent coupling potentials $V_{nm}(\theta,t)$ are given by:
[…]
Once the coefficients are obtained, the excitation probability for a $0\rightarrow n$ transition is given by:
$P_{n}(\theta)=|c_{n}(\theta,\infty)|^{2},$
and the differential cross section
$\left({d\sigma\over d\Omega}\right)_{0\to n}=\left({d\sigma\over d\Omega}\right)_{\mathrm{clas}}P_{n}(\theta).$
where $(d\sigma/d\Omega)_{\mathrm{clas}}$ is the classical differential elastic cross section which, for a pure Coulomb case, coincides with the Rutherford cross section.
Due to the conservation of the total probability (flux), one has
$\sum_{n}P_{n}(t)=\sum_{n}|c_{n}(t)|^{2}=1.$
When the couplings are weak, one may solve (120) perturbatively, assuming that $c_{0}\approx 1$ and $c_{n}\ll 1$ for $n>0$. This gives the first-order solution
[…]
$\left({d\sigma\over d\Omega}\right)_{0\rightarrow n}=\left({Z_{t}e^{2}\over\hbar v}\right)^{2}{B(E\lambda,0\to n)\over e^{2}a_{0}^{2\lambda-2}}f_{\lambda}(\theta,\xi)$ (123)
which is valid only for angles smaller than the grazing one ($\theta<\theta_{\text{gr}}$) and where $a_{0}$ is half of the distance of closest approach in a classical head-on collision, $\xi_{0\rightarrow n}=\frac{(\varepsilon_{n}-\varepsilon_{0})}{\hbar}\frac{a_{0}}{v}$ is the adiabaticity parameter and $f_{\lambda}(\theta,\xi)$ is an analytic function, depending on the kinematical conditions, but independent of the structure of the projectile.
For weakly bound nuclei, the excitation will typically populate unbound (continuum) states. The previous formula can be generalized to:
$\frac{d\sigma(E\lambda)}{d\Omega d\varepsilon}=\left({Z_{t}e^{2}\over\hbar v}\right)^{2}{1\over e^{2}a_{0}^{2\lambda-2}}\frac{dB(E\lambda)}{d\varepsilon}{df_{\lambda}(\theta,\xi)\over d\Omega}\quad\quad\ (\theta<\theta_{\text{gr}})$ (124)
where $df_{\lambda}(\theta,\xi)/d\Omega$ is also a well-defined analytic function and $dB(E\lambda)/d\varepsilon$ is the electric reduced probability to the continuum states.
It is common to express (124) in terms of the so-called number of equivalent photons,
$\frac{d\sigma(E\lambda)}{d\Omega d\varepsilon}=\frac{16\pi^{3}}{9\hbar c}\frac{dN_{E\lambda}(E_{x}\theta)}{d\Omega}\frac{dB(E\lambda)}{d\varepsilon}$ (125)
where ${dN_{E\lambda}(E_{x}\theta)}/{d\Omega}$ is the number of equivalent photons per solid angle.
Figure 44, quoted from Ber09 , shows the number of equivalent virtual photons, integrated for impact parameters beyond $b=12.3$ fm, for the collision of a projectile with a $Z=82$ target, for three different incident energies. It can be seen that increasing the incident energy decreases the population of low-lying states and enhances the population of high-energy ones.
Figure 44: Total number of virtual photons for the E1 multipolarity, corresponding to the scattering of a projectile by a 208Pb target at different incident energies and integrated for impact parameters $b>12.3$ fm. Quoted from Ref. Ber09 .
7.2 Application to the extraction the electric transition probabilities from Coulomb dissociation data
Coulomb excitation has been used for many years as a powerful spectroscopic tool to extract the electromagnetic response of nuclei to a Coulomb field, provided by a heavy nucleus. For stable nuclei and low bombarding energies (tens of MeV per nucleon), Coulomb excitation is dominated by transitions to low-lying excited states connected by quadrupole and octupole multipolarities with the ground state, whereas at sufficiently high energies, Coulomb excitation is dominated by the so-called giant dipole resonance (GDR), characterized by collective motion of protons and neutrons moving out-of-phase motion. In the case of weakly bound nuclei and, more specifically, for neutron halo nuclei, the differential force of the Coulomb field acting on the valence neutron(s) with respect to the core produces a strong dipole force that may eventually break the system. This mode has been customarily termed as “soft dipole E1” excitation, as opposed to the giant dipole excitation, see Fig. 45.
Figure 45: B(E1) distribution for nucleus 11Li. At low energies, the peak associated to the weakly-bound neutrons forms the so-called “pygmy” or “soft dipole” resonance. Figure taken from Ref. Bro19 .
The concentration of dipole strength at low-excitation energies is a general property of weakly-bound nuclei, particularly those with a neutron-halo structure. This can be verified considering a simple situation of a two-body system described with a single-particle model in which the $B(E1)$ can be readily evaluated numerically. Considering such model, Fig. 46 displays the $B(E1)$ transition for a hypothetical (and unrealistic) system with mass $A=8$ and a nucleon binding energy $E_{b}$ = 0.7 MeV for different assumptions of the initial state orbital angular momentum. In the valence proton case, the strength distribution is shifted to higher energies, together with a damping of the absolute value. The concentration of strength becomes very large near the breakup threshold for an initial state with $\ell=0$.
In the limit of vanishing binding energy, one can derive analytical expressions for the $B(E1)$ distribution. The basic approximation is that most of the contribution of the $B(E1)$ comes from the external part of the initial and final wavefunctions so that they can be replaced by their asymptotic forms. For example, in the case of a $s$ to $p$ transition this leads to the simple expression
$\frac{dB(E1)}{d\varepsilon}(s\rightarrow p)=\frac{3\hbar^{2}}{\pi^{2}\mu}\left(Z_{\text{eff}}e\right)^{2}\frac{\sqrt{E_{b}}\varepsilon^{3/2}}{\left(\varepsilon+E_{b}\right)^{4}}$ (126)
where $Z_{\text{eff}}=-Z/A$ is the effective charge for the case of a neutron halo. The expression above predicts a maximum at $\varepsilon=3/5E_{b}$ and hence, the smaller the binding energy, the closer the position of the maximum to the breakup threshold.
Figure 46: B(E1) strength distribution for different situations, all associated with single-particle excitation from a bound state to the continuum. In all cases the initial particle is bound with the same energy ($E_{b}$ = 0.7 MeV). The labels “neutron” and “proton” denote, respectively, the neutron and proton excitations. For both cases, the plot shows the situations of initial $s$-state (with dipole transition to $p$-state in the continuum) and of initial $p$-state (with final $s$-state in the continuum). Bound and continuum single-particle states were generated from a Woods-Saxon potential, with parameters adjusted to reproduce the correct binding energy of the initial bound state. Taken from Nag05 .
The aim of a Coulomb dissociation experiment with drip-line nuclei is the extraction of the electric response to the continuum that, in the case of halo nuclei, is expected to be largely dominated by its $E1$ distribution.
An appealing feature of the semiclassical formulas outlined in the previous section is that they provide a neat separation between the structure of the projectile (through $B(E\lambda)$) and the reaction dynamics. This separation allows in principle to extract the ${dB(E\lambda)}/{d\varepsilon}$ distribution from the analysis of small-angle Coulomb dissociation data. For that, one can integrate the double differential cross section (124) up to a maximum angle ($\theta_{\mathrm{max}}$) such that for $\theta<\theta_{\mathrm{max}}$ the breakup can be assumed to be purely Coulomb and hence (124) is valid ($\theta_{\mathrm{max}}<\theta_{\text{gr}}$). This gives:
[…]
Figure 47: Top: angular distribution of the c.m. of the outgoing $n$+10Be system, following the breakup of 11Be on 208Pb at 69 MeV/u. Bottom: relative energy distribution between the $n$+10Be fragments, integrated up to two different maximum angles, as indicated by the labels. The curves are semiclassical calculations based on the theory of Alder and Winther with a model $B(E1)$ distribution obtained from a simple two-body potential model and smeared with the experimental resolution. Experimental data are from Fuk04 .
Figure 48: Electric transition probability for 11Be as a function of the excitation energy extracted from Coulomb dissociation experiments of 11Be on 208Pb at 69 MeV/u Fuk04 and 520 MeV/u Pal03 , measured at RIKEN and GSI, respectively.
An application of this method to the 11Be+208Pb reaction discussed above is shown in Fig. 47. The data correspond to the measurements of Fukuda et al. Fuk04 performed at RIKEN at 69 MeV/u. The curves are the result of semiclassical calculations using the EPM with a two-body structure model for 11Be and convoluted with the experimental energy and angular resolutions. In the top panel, it is apparent that the semiclassical calculation reproduces well the data up to $\theta_{\text{c.m.}}\sim 3^{\circ}$. For larger angles, the assumption of pure Coulomb scattering is no longer valid (the grazing angle is estimated to be $\theta_{\text{gr}}=3.8^{\circ}$). In the bottom panel, the corresponding relative energy distributions are compared with the data, for two angular cuts. It is seen that, for the wider angular range ($\theta_{\mathrm{c.m.}}<6^{\circ}$), there is some underestimation of the data, which may be attributed to the omission of nuclear effects and higher Coulomb multipoles.
Restricting the analysis to the lower cut ($\theta_{\mathrm{c.m.}}<1.3^{\circ}$), and assuming that for these angles the dissociation probability is dominated by the Coulomb dipole couplings, one can use the approximation (127) and extract the $B(E1$) distribution. The result of such an analysis, performed in Fuk04 , is shown in Fig. 48, which shows the $B(E1)$ extracted in Fuk04 (black squares).
In the particular case of 11Be the ground state is known to contain admixtures of $s$ and $d$-waves for the neutron-core relative motion (see Sec. 3.3). As noted above, the $B(E1)$ is largely dominated by the asymptotic part of the ground-state wavefunction and hence most of the contribution is due to the $s$-wave component. Consequently, the magnitude of the $B(E1)$ is roughly proportional to the weight of this component or, more correctly, by the spectroscopic factor associated with the overlap $\langle^{11}\mathrm{Be}(g.s.)|^{10}\mathrm{Be}(0^{+})\otimes 2s_{1/2}\rangle$. This way, Coulomb dissociation experiments constitute complementary spectroscopic tools to transfer or knockout reactions. As in those cases, a possible caveat of this methodology is the dependence of the spectroscopic factor on the assumed single-particle model. A detailed discussion of this technique can be found in recent topical reviews Aum13 ; Nak17 .
If the approximations implied by the EPM are well fulfilled, the extracted $B(E\lambda$) distributions should be independent of the incident energy and the target. This is not always the case, as can be seen by comparing the $B(E1)$ discussed above with that obtained from the analysis of another 11Be+208Pb experiment performed at GSI by Palit et al. Pal03 at the much higher energy of 520 MeV/u, shown in Fig. 48 by the yellow circles. The origin of this discrepancy is further discussed in Sec. 7.6).
[…]
The assumption of pure Coulomb trajectories can be relaxed, at the expense of losing some of the simplicity of the method. A compelling application is shown in Fig. 49 (taken from Ref. Mar02 ) where semiclassical coupled-channels calculations, using trajectories modified by the nuclear interaction, are compared with CDCC calculations for the breakup angular distribution of the 8B+58Ni reaction at the near-barrier energy of 26 MeV, finding a nice agreement between both.
7.3 Dynamic Coulomb polarization potential from the AW theory
Feshbach theory of nuclear reactions provides formal expressions for the contribution on the optical potential of nonelastic channels. These contributions are such that, when added to the bare interaction between the colliding nuclei (that is, the static part of the interaction due to the ground-state densities for the projectile and target) yield the “correct” elastic scattering cross section. In general, this dynamical contribution of the optical potential, termed polarization potential, is nonlocal, and energy and angular momentum dependent. Under suitable approximations, it is possible to derive simple, analytical expressions for specific parts of the polarization potential.
Figure 50: Second-order process describing the excitation and de-excitation of a nucleus moving along a classical trajectory determined by the potential $V_{0}(t)$, due to the action of the potential $V_{1}(t)$.
In the case of Coulomb excitation to either bound or unbound states (breakup) the semiclassical theory of Alder and Winther outlined in the previous section provides a very convenient framework for obtaining these simple forms of the Coulomb polarization potential. One such example is provided by the works of Andrés, Gómez and Nagarajan And94 ; And95 .
These works consider the second-order process in which the projectile is excited to a continuum state due to the dipole Coulomb interaction ($V_{1}(t)$) and then de-excites, returning to the ground state. The situation is schematically depicted in Fig. 50. The amplitude probability for this process is given by:
[…]
×$\times\int_{-\infty}^{t}dt^{\prime}\langle z|V_{1}(t^{\prime})|0\rangle\exp\left\{{i\over\hbar}(\varepsilon_{z}-\varepsilon_{0})t^{\prime}\right\},$ (128)
where $|z\rangle$ denotes the intermediate states populated during the reaction. Then, one requires that this second-order amplitude coincides with the first-order amplitude associated with the polarization potential for all classical trajectories corresponding to a given scattering energy. This gives rise to the following expresssion:
$U_{\mathrm{pol}}(R)$ $=-\frac{4\pi}{9}\frac{Z_{t}^{2}e^{2}}{\hbar v}\frac{1}{(R-a_{o})^{2}R}$
×$\times\int^{\infty}_{0}d\varepsilon\frac{dB(E1,\varepsilon)}{d\varepsilon}\left[g\left(\frac{R}{a_{o}}-1,\xi\right)+if\left(\frac{R}{a_{o}}-1,\xi\right)\right],$ (129)
where g and f are analytic functions defined as
$f(z,\xi)$ $=$ $4\xi^{2}z^{2}\exp{(-\pi\xi)}K_{2i\xi}^{\prime\prime}\left(2\xi z\right),$ (130)
[…]
and $\xi$ is the Coulomb adiabaticity parameter corresponding to the excitation to the continuum energy $\varepsilon$ of the nucleus. An important feature of this potential is that when the breakup energy $\varepsilon_{b}$ is large enough, the purely real adiabatic dipole potential given by Eq. (2). In the opposite limit, for small breakup energies, $f\left(\frac{R}{a_{o}}-1,\xi\right)\rightarrow 1$ and $g\left(\frac{R}{a_{o}}-1,\xi\right)\rightarrow 0$, and the polarization potential becomes purely imaginary, depending on $R$ as ${1}/[(R-a_{o})^{2}R].$
Figure 51: Elastic scattering differential cross section for 6He+208Pb at $E_{\mathrm{lab}}=22$ MeV. The dashed line is a single-channel calculation performed with a bare potential, given by a 6Li+208Pb optical potential. The solid line is the optical model calculation obtained with the bare plus dynamic Coulomb polarization potential of Eq. (7.3).
Figure 52: Real and imaginary parts of the bare and CDP potentials for 6He+208Pb at $E_{\mathrm{lab}}=22$ MeV computed with dynamic Coulomb polarization potential of Eq. (7.3).
As an example, we show in Fig. 51 the elastic scattering data for 6He+208Pb at 22 MeV compared with the optical model calculations obtained with a bare potential alone (dashed) and with the bare plus the Coulomb dipole polarization (CDP) potential (solid line). For the bare interaction, we have used an optical potential for the nearby projectile 6Li, which has a structure similar to that of 6He but which does not exhibit the strong dipole excitation mechanism. The elastic cross section predicted by the bare potential presents a Fresnel-like diffraction pattern, not observed in the data. The inclusion of the CDP significantly reduces the elastic cross section and suppresses the Fresnel peak. In Fig. 52, the real and imaginary parts of the bare and CDP potentials are shown. Their most remarkable feature is their long range, which was anticipated from the optical model analysis of Sec. 2, and is associated with the long range of the dipole Coulomb couplings, as we have already discussed in the context of the CDCC method.
7.4 Semiclassical transfer-to-the-continuum model
The semiclassical STC model Bon88 developed by Bonaccorso and Brink is a generalization of the model of Brink and collaborators Has78 ; Has79 ; Mon85 ; Bon87 to the case of final unbound states. These methods provide a more transparent and often analytical interpretation of the quantum-mechanical counterparts. The method is valid for peripheral reactions based on the core spectator model as in the IAV model discussed in previous sections. To obtain an analytic expression for the transfer amplitude the STC replaces the internal state wavefunctions by their asymptotic forms, given analytically in terms of Hankel functions.
Considering the case of a neutron transfer, the starting point of the STC is the semiclassical transfer to the continuum scattering amplitude, given by Bon88
[…]
with $\bar{\omega}=(\omega t-mvz/\hbar)$. The time dependent nucleon initial and final wave functions in their respective reference frames are
$\psi_{i,f}(\boldsymbol{r}_{n},t)=\phi_{i,f}(\boldsymbol{r}_{n})e^{-i\varepsilon_{i,f}t/\hbar}$ (133)
The initial and final radial wave functions are approximated by Hankel functions according to Bon88 . The initial state is therefore approximated as:
$\phi_{l_{i}}(\boldsymbol{r}_{Cn})\simeq-C_{i}i^{l}\gamma h^{(1)}_{l_{i}}(i\gamma_{i}r_{Cn})Y_{l_{i},m_{i}}(\Omega_{i}).$ (134)
with $C_{i}$ the asymptotic normalization coefficient of the initial bound state wavefunction. The the final (continuum) state is approximated by:
$\phi_{l_{f}}(\boldsymbol{r}_{n})\simeq C_{f}k_{f}\frac{i}{2}(h^{(+)}_{l_{f}}(k_{f}r_{n})-S_{l_{f}}h^{(-)}_{l_{f}}(k_{f}r_{n}))Y_{l_{f},m_{f}}(\Omega_{f}),$ (135)
where $\ S_{l_{f}}(\varepsilon_{f})$ is the neutron-target elastic scattering S-matrix at energy $\varepsilon_{f}$ and $C_{f}$ a normalization constant.
With these approximate initial and final wavefunctions, the one-neutron removal probability results:
$\frac{dP_{-n}}{d\varepsilon_{f}}$ $\approx\frac{1}{2}\sum_{j_{f}}(2j_{f}+1)|1-S_{j_{f}}|^{2}(1+R_{if})\left[\frac{\hbar}{mv}\right]^{2}$
×$\times\frac{m}{\hbar^{2}k_{f}}|C_{i}|^{2}\frac{e^{-2\eta b_{c}}}{2\eta b_{c}}M_{l_{f}l_{i}},$ (136)
The factor $M_{l_{f}l_{i}}$ is due to the overlap of the angular parts, and $R_{if}$ are spin-coupling coefficients. Further definitions and discussion can be found in Refs. Bon88 .
To calculate the elastic $S$-matrix a choice must be made for the neutron-target potential. In general, such potential will be complex and energy dependent. At low neutron-target relative energies, narrow resonances are expected that will progressively become broad and overlapping at higher energies. In the spirit of the optical model, an energy averaging procedure is assumed, resulting:
$\frac{dP_{-n}}{d\varepsilon_{f}}(j_{f},j_{i})=\left(|1-\bar{S}_{j_{f}}|^{2}+(1-|\bar{S}_{j_{f}}|)^{2}\right)B_{if}$ (137)
The first and second terms correspond to the elastic and nonelastic breakup contributions discussed in the context of the IAV model. The $\bar{S}_{j_{f}}$ are nucleon-target S-matrices calculated for each nucleon final energy according to the optical model with an energy-dependent optical potential. The $B_{if}$ are kinematical factors depending on the energies ($\varepsilon_{i}$ and $\varepsilon_{f}$), momenta ($\gamma_{i}$ and $k_{f}$), and orbital angular momenta ($l_{i}$ and $l_{f}$) of the initial and final states of a single neutron particle, respectively, and on the incident energy per particle, $mv^{2}/2$ at the distance of closest approach $d$ , which coincides with the projectile impact parameter in a straight line trajectory. Their explicit form can be found in the original works Bon88 .
To obtain the relative energy distribution of the final states, one integrates the probabilities over the core impact parameters
[…]
where $S_{c}(b_{c})$ is the core-target S-matrix and $|S_{c}(b_{c})|^{2}$ is therefore a core-survival probability.
For sufficiently high incident energies, the $S$-matrix can be evaluated in the eikonal approximation and one can also use the connection between the angular momentum and impact factor to transform the sum in Eq. (136) by an integral over the impact parameter, i.e, $\Sigma_{j_{f}}\to\int d^{2}\boldsymbol{b}_{n}$.
Figure 53: n-10Be relative energy spectrum, including Coulomb and nuclear breakup for the reaction 11Be+12C $\to$ n+10Be+X at 69 A.MeV. Only the contributions from an s initial state with spectroscopic factor C2S= 0.84 are calculated. The triangles are the total calculated result after convolution with the experimental resolution function. The dots are the experimental points from Fuk04 . Figure taken from Ref. Blan07 , with authorization from Elsevier.
An example of the application of the method is given in Fig. 53. It corresponds to the reaction 11Be+12C $\to$ $n$+10Be+12C. The STC calculations for different angular momenta of the final state are compared with the data from Fukuda et alFuk04 .
7.5 Dynamical Eikonal approximation
The dynamical eikonal approximation (DEA) developed by the group at the ULB Bay05 ; Gol06 unifies the semiclassical time-dependent and eikonal methods. It allows for the calculation of differential cross sections for elastic scattering and breakup in a quantal way by taking into account interference effects.
As in the CDCC method, the idea of the method is to provide an approximate solution of the three-body Schrodinger equation with the effective Hamiltonian (5). The three-body scattering solution of this equation is expressed in the factorized form:
$\Psi^{\rm DEA}(\vec{R},\vec{r})$ $=e^{iK_{0}Z}\widehat{\Psi}(\vec{R},\vec{r}).$ (139)
where $K_{0}$ stands for the incident momentum and $Z$ is the third component of the $\vec{R}$ vector. At sufficiently high energies, the deviation from the initial plane wave $e^{iK_{0}Z}$ of the projetile-target relative motion is expected to be small and hence the dependence on $\vec{R}$ of $\widehat{\Psi}$ is expected to be smooth. This enables to neglect its second-order derivative in $\vec{R}$ with respect to its first-order derivative
$\nabla^{2}_{\vec{R}}\widehat{\Psi}(\vec{R},\vec{r})\ll K_{0}\frac{\partial\widehat{\Psi}(\vec{R},\vec{r})}{\partial Z}.$ (140)
Therefore, introducing the factorization (139) into the three-body Schrodinger equation leads to the so-called DEA equation
$i\frac{\hbar^{2}K_{0}}{\mu_{aA}}\frac{\partial}{\partial Z}\widehat{\Psi}(Z,\vec{b},\vec{r})=$
$\left[(H_{0}-\varepsilon_{0})+U_{bA}(\vec{r}_{bA})+U_{xA}(\vec{r}_{xA})\right]\widehat{\Psi}(Z,\vec{b},\vec{r}),$ (141)
where the three-body wavefunction has been conveniently expressed in terms of the $Z$ and transverse $\vec{b}$ parts of the projectile-target coordinate $\vec{R}$.
The DEA equation (140) is mathematically equivalent to a time-dependent Schrodinger equation for a straight-line trajectory. It can therefore be solved using similar numerical techniques as in the time-dependent model (see e.g. Kid94 ; Cap03 ) The DEA equation must be solved for each transverse component $\vec{b}$ of the $a$-$A$ coordinate with the initial boundary condition
$\widehat{\Psi}(Z\rightarrow-\infty,\vec{b},\vec{r})=\phi_{0}(\vec{r}).$ (142)
Breakup amplitudes can then be extracted from the wave function as Bay05 :
[…]
where $\vec{q}=\vec{K}-K\hat{Z}$ is the transferred momentum and
$S(\vec{k},\vec{b})=\langle\phi_{\vec{k}}^{(-)}(\vec{r})|\hat{\Psi}(\vec{R},\vec{r})\rangle_{Z=+\infty}$ (144)
where $\phi_{\vec{k}}^{(-)}(\vec{r})$ is a $b+x$ scattering state with incoming boundary conditions evaluated at the desired final energy.
Despite its formal analogy, it is important to note that there is no semiclassical approximation involved in the derivation of the method and so the coordinates $Z$ and $\vec{b}$ are quantal variables. This enables us to take into account interferences between trajectories.
The DEA method has been tested against the CDCC and time-dependent semiclassical methods for the breakup of halo nuclei. These studies indicate that the DEA predictions agree very well with those of the CDCC method at energies of around 100 MeV/u and even below. Moreover, it is able to reproduce the angular pattern of breakup, that is not given by the semiclassical approximations. An example of this benchmark comparison is given in Fig. 54.
Figure 54: Top: Differential breakup cross sections, with respect to the $n$-14C relative energy, for the breakup of 15C on Pb at 68 MeV/nucleon. The CDCC, DEA and a semiclassical time-dependent method (TD) are compared. Bottom: Breakup angular distributions from the same calculations. Quoted from Cap13 .
7.6 Coulomb dissociation revisited: quantum-mechanical and nuclear effects
In Sec. 7 we presented the standard procedure of extracting electric transition probabilities from Coulomb dissociation data, making use of the first-order semiclassical theory of Alder and Winther (i.e., the equivalent photon method). This procedure faces several difficulties, namely
  1. (i)
    in addition to Coulomb couplings, nuclear couplings are always present to a greater or lesser extent;
  2. (ii)
    even in situations in which the nuclear contribution is small, other multipoles, other than the dipole one, may be present;
To overcome (i), two alternative strategies have been commonly adopted. The first consists in measuring the breakup data at sufficiently small c.m. scattering angles, under the assumption that, for angles well below the grazing one, nuclear effects are negligible (the so-called “safe-Coulomb” region). This method has been adopted, for example, in exclusive breakup experiments performed at RIKEN, at typical energies around 100 MeV per nucleon (e.g. Nak99 ; Nak06 ; Fuk04 ). A second, alternative, strategy is to collect the breakup data in a wider angular range. This has the advantage of increasing the statistics, but the disadvantage of including nuclear contributions. The latter must be estimated and subtracted from the breakup data for a meaningful extraction of the electric contribution.
Each of these strategies has pros and cons. The “safe-Coulomb method” has the disadvantage of considering only a fraction of the breakup data. Moreover, nuclear breakup can extend down to very small angles. For example, in the analysis of the 11Li+208Pb Coulomb dissociation experiment performed at RIKEN at an incident energy of 70 MeV per nucleon Nak06 , the authors considered data below $1.46^{\circ}$.
The “safe-Coulomb” assumption is tested in Fig. 55 for the 11Be +208Pb reaction, which shows the n-10Be relative energy distribution for two upper cuts of the scattering angle, $\theta_{\mathrm{max}}=6^{\circ}$ (top) and $1.3^{\circ}$ (bottom). The lines are XCDCC calculations based on a particle-plus-rotor model of 11Be, including both Coulomb and nuclear couplings or only nuclear couplings. It is evident from the figure that nuclear effects are still very relevant for $\theta_{\mathrm{max}}=6^{\circ}$ but are almost suppressed (but not negligible) for the lower cut. This gives an idea of the accuracy and validity of the EPM analysis performed in Fuk04 to extract the $B(E1)$ distribution for 11Be.
Figure 55: Relative energy distribution of the fragments following the breakup of 11Be on a lead target at 69 MeV/u, integrated in the scattering angle up to a maximum value of $\theta_{\mathrm{max}}=6^{\circ}$ (top) and $1.3^{\circ}$ (bottom). Full CDCC calculations, including both Coulomb and nuclear couplings are compared with CDCC calculations including only nuclear couplings.
The alternative strategy, in which angle-integrated cross sections are considered for the analysis, has the advantage of profiting from the entire breakup data, but deals with the uncertainties associated with the estimation of the nuclear contribution. For that, measurements with a high-$Z$ target nucleus are combined with those for a light nucleus, like 12C, for which the nuclear contribution is expected to be dominant. Thus, the nuclear contribution in the reaction with the heavy target is accounted for by subtracting a properly scaled cross section measured with the carbon target. The scaling factor can be estimated from the respective radii of the heavy and light targets. Alternatively, one may resort to eikonal formulas for stripping and diffraction to estimate the cross section ratio between heavy and light targets, as done in Pal03 .
Besides the additional source of uncertainty added by the need of estimating the nuclear contribution, another caveat of this method is the neglect of possible interference effects between the nuclear and Coulomb contributions. The importance of these effects can be assessed by means of quantum-mechanical calculations, such as those given by the CDCC method.
As an example, in Fig. 56 the effect of Coulomb/nuclear interference is illustrated for the 11Be+208Pb reaction at 520 MeV/u. The plot shows EPM and XCDCC calculation. In the latter, different choices for the nuclear and Coulomb couplings are shown. It is seen that (i) the nuclear breakup is rather significant (consistently with the estimate of Pal03 ) and, more importantly, that (ii) the incoherent sum of the nuclear and Coulomb contributions overestimates the full XCDCC calculations in which Coulomb and nuclear couplings are included simultaneously and hence interference effects appear. From these calculations, it becomes apparent that the procedure of summing incoherently the nuclear and Coulomb contributions introduces some error in the extraction of the electric transition probabilities from Coulomb dissociation experiments.
Figure 56: Relative energy distribution of the fragments following the breakup of 11Be on a lead target at 520 MeV/u. EPM and XCDCC calculations, based on the same structure model, are compared. The effect of the nuclear/Coulomb interference is illustrated.
In addition of nuclear contribution and its interference with the Coulomb contribution, higher-order couplings effects may also affect the extraction of electric transition probabilities from cross section data. In principle, one might rely on more sophisticated calculations, such as those provided by the CDCC and XCDCC methods, in which nuclear and Coulomb couplings are included to all orders and their interference is properly accounted for. The price to pay is that these calculations are more computationally demanding, and more importantly, they do not provide a simple connection between the reaction observables (cross sections) and the underlying $B(E\lambda)$ distributions. A recent work Mor20 proposed a method to extract $B(E1)$ distribution by comparing exclusive breakup data with CDCC calculations based on some reasonable trial structure model of the weakly-bound nucleus, which is subsequently corrected by comparing the CDCC and experimental cross sections.
An application of this method is shown in Fig. 57, adapted from Mor20 . The solid black line is the $B(E1)$ distribution provided by the starting $B(E1)$ model used in the XCDCC calculation. This $B(E1)$ is corrected for each excitation energy by comparing the calculated and measured cross sections, as explained in Mor20 . The procedure is applied separately to the data of Ref. Fuk04 , for the two angular cuts discussed above, as well as to the data of Pal03 . The corrected $B(E1)$ distributions are displayed in this figure with the symbols. It is seen that the three of them agree reasonably well within the estimated errors. This is in contrast to the $B(E1)$ extracted in the original works and shown in Fig. 48. It is important to note that the latter are affected by the experimental resolution, while those shown in Fig. 57 do not include the effect of the experimental energy resolution. Differences between these experimental resolutions in the two experiments are partially responsible for the differences observed in the $B(E1)$ distributions reported in the original works.
Figure 57: Unfolded $B(E1)$ distributions extracted from the experimental breakup cross sections measured in Pal03 (squares) and Fuk04 (circles) using the procedure based on the XCDCC method proposed in Mor20 . For comparison, the starting $B(E1)$ distribution is shown by the black line and NCSMC ab-initio calculation of Ref. Cal16 (red solid line) is also shown. Figure adapted from Mor20 .
8 Study of nucleon-nucleon correlations from nucleon removal reactions
One particular point of interest for nucleon-Borromean systems of the form N+N+C is the study of the correlation between their two valence nucleons, since the nucleon-nucleon interaction is essential to hold the Borromean system bound. As many Borromean systems also present nucleon halos, the study of the correlation between their valence nucleons permits the exploration of the nucleon-nucleon interaction in low-density, isospin-asymmetric environments. A quantity related to the nucleon-nucleon correlation is the average opening angle between nucleons with respect to the core $\left\langle\theta_{NN}\right\rangle$. If both nucleons follow independent uncorrelated orbits, this angle will on average be $90^{\circ}$ Esb92 ; Ber07 . A smaller angle corresponds to a small average distance between nucleons compared to the distance of their center of mass to the core, in which is known as a “dineutron” (or diproton) configuration mig73 , which is associated with an attractive correlation between nucleons. Meanwhile, an angle larger than 90 indicates a larger internucleon distance compared to their distance to the core, which corresponds to both nucleons being on opposite sides of the core, in the so-called “cigar” configuration, which corresponds to a repulsive nucleon-nucleon correlation.
An experimental probe for $\left\langle\theta_{NN}\right\rangle$ in the case of neutron-Borromean systems is the dipole electric strength $B(E1)$: assuming point-like particles, its expression reduces to:
$\begin{split}B(E1)&=\dfrac{3}{\pi}\left(\frac{Z_{c}}{A}\right)^{2}e^{2}\left\langle r_{n}^{2}+r_{n^{\prime}}^{2}+2r_{n}r_{n^{\prime}}\cos\theta_{NN}\right\rangle\\ &\simeq\dfrac{6}{\pi}\left(\frac{Z_{c}}{A}\right)^{2}e^{2}\left\langle r_{n}^{2}\right\rangle(1+\cos\left\langle\theta_{NN}\right\rangle),\end{split}$ (145)
where $\left\langle r_{n}^{2}\right\rangle$ is the average square distance between neutron and core, $Z_{c}$ is the charge of the core and $A$ the mass of the nucleus Bert91 , so an enhanced $B(E1)$ indicates a smaller angle and dineutron correlations while a reduced $B(E1)$ indicates the existence of a cigar-like configuration. $B(E1)$ strengths have been measured for various Borromean nuclei, from which values for $\left\langle\theta_{NN}\right\rangle$ can be extracted. The values extracted in Ber07 are presented in Table 3, where the angles can be seen to suffer from significant uncertainties, due both to experimental uncertainties and the model dependence due to $\left\langle r_{n}^{2}\right\rangle$.
6He 11Li 14Be 17Ne
B(E1) ($e^{2}fm^{2}$) 1.20(20) 1.42(18) 1.69 1.56
Aum99 Nak06 Zhu95 Gri06
[…]
Table 3: B(E1) strength function and deduced average opening angle $\left\langle\theta_{NN}\right\rangle$ for various Borromean nuclei. Table adapted from Ber07 .
Another method of extracting the opening angle is the use of nucleon knockout reactions, where one of the valence nucleons is removed via a sudden reaction, such as nucleon-knockout with 9Be or 12C targets or a $(p,pN)$ reaction with proton targets, assuming the validity of the spectator approximation for the remaining $(C+N)$ system. Within this approximation, in the projectile rest frame the Borromean nucleus start at rest, so that, when the nucleon is removed, the final momentum of the $(C+N)$ system will be equal and opposite to that of the removed nucleon within the Borromean system ($\boldsymbol{k}_{y}$). Since the $(C+N)$ system is unbound, the relative momentum between $C$ and the remaining nucleon can also be measured $(\boldsymbol{k}_{x})$, and the angle between them $\theta_{k}$ determined. A scheme of the quantities of interest is shown in Fig. 58.
Figure 58: Opening angle and momenta measured in nucleon-removal reactions on Borromean systems. Figure adapted from Cor23 .
It should be noted that, as $\theta_{k}$ is an angle between the momenta associated to the nucleons instead of their positions, it can be viewed as a Fourier transform of the opening angle $\theta_{NN}$. As such, a value of $\theta_{k}$ larger than $90^{\circ}$ is associated to a value of $\theta_{NN}$ lower than $90^{\circ}$, and thus to a dineutron configuration, while the cigar configuration corresponds to $\theta_{k}<90^{\circ}$ and $\theta_{NN}>90^{\circ}$. Various measurements of $\theta_{k}$ have been performed for Borromean nuclei Chu97 ; Sim99 ; Mar01 ; Sim07 , finding for its average value $\left\langle\theta_{k}\right\rangle$ the following results (in degrees): $90$ (6He), $90$ (8He), $103.6^{+0.7}_{-0.9}$ (11Li), $96.2^{+1.8}_{-1.6}$ (14Be) Sim07 . It can be seen that, indeed, values for $\left\langle\theta_{k}\right\rangle>90^{\circ}$ correspond to $\left\langle\theta_{NN}\right\rangle<90^{\circ}$, as shown in Table 3. It should also be noted that an asymmetric distribution in either $\theta_{k}$ or $\theta_{NN}$ (resulting in $\left\langle\theta\right\rangle\neq 90^{\circ}$) requires the interference of waves with opposite parities Cat84 , which neatly explains the symmetry for 6,8He, since the valence neutrons find themselves in the negative-parity $p_{3/2}$ and $p_{1/2}$ waves, so their angular distribution must be symmetric. This also sets the average opening angle as a measure of the admixture of opposite parity components Kub20 .
As seen above, for neutron-Borromean systems, the dineutron correlation seems to dominate over the cigar distribution. Recent experiments have tried to explore the spatial location of this dineutron correlation, as some nuclear matter calculations point to the low-density surface of Borromean systems as a region where these correlations should be favoured Mats06 . Recent $(p,pn)$ experiments on Borromean nuclei Kub20 ; Cor23 have used the modulus of the momentum of the removed neutron ($k_{y}$) as a proxy for the nuclear location (smaller momenta correspond to the nuclear halo and surface, while larger momenta correspond to the nuclear interior) and the average opening angle as a proxy for dineutron correlations. Through the use of an eikonal zero-range approximation for the $(p,pn)$ nucleon-nucleon collision process, the cross section can be expressed as Kik16 ; Cas21 :
$\sigma\propto\left|\left\langle\phi_{Cn}(\boldsymbol{k}_{x},\boldsymbol{x})\otimes e^{i\boldsymbol{k}_{y}\cdot\boldsymbol{y}}|S(y)\phi_{gs}(\boldsymbol{x},\boldsymbol{y})\right\rangle\right|^{2},$ (146)
where $\boldsymbol{x}$ and $\boldsymbol{y}$ are defined as in Sec. 4.2, $\phi_{gs}$ describes the bound state of the Borromean nucleus, $\phi_{Cn}$ the final-state wave-function between the core and the remaining neutron, and $S(y)$ is a $S-$matrix that includes the absorptive potential between target proton and core $C$. As such, the cross section can be interpreted as the square of the Fourier transform of the bound state, distorted by the absorption of the target proton and the final-state interaction between neutron and core. In Kub20 , the maximum $\left\langle\theta_{k}\right\rangle$ for 11Li was found for small values of $k_{y}\sim 0.3$ fm-1, which was interpreted as the dineutron being located at the nuclear surface where the nuclear density is low, as predicted by previous models Mats06 ; Hag07 . As the low nuclear density is a feature of the surface of all neutron-Borromean nuclei, it is expected that the dineutron also appears in the surface of other such systems. Indeed, in Cor23 , the maximum of $\left\langle\theta_{k}\right\rangle$ was found for roughly the same value of $k_{y}$ also for 14Be and 17B, as shown in Fig. 59, making a tantalizing case for the universality of dineutron correlations in the surface of neutron-Borromean nuclei. as predicted in Mats06 , although experiments on more nuclei, such as 19B or 22C, are required in order to confirm this feature. In the figure the green bands correspond to systematic errors in experimental data, while statistical errors are indicated by the errorbars. The theoretical calculations are able to match experimental results for 11Li and 17B but overestimate the maximum average angle for 14Be. This has been associated to the contribution of excited states of the 12Be beyond the $2^{+}$ state included in the theoretical model Cor23 .
Figure 59: Average opening angle in momentum space $\left\langle\theta_{k}\right\rangle$ as a function of $k_{y}$. Green bands correspond to the systematic errors in the experimental measurement. In all nuclei the maximum average momentum is found at similar values of $k_{y}$ for all considered nuclei. Figure adapted from Ref. Cor23 .
9 Uncertainty evaluation
Given the amount of ingredients required for the computation of reaction observables (bound-state wavefuncions, optical potentials, spectroscopic factors,…) there has been a push for the quantification of the uncertainties in the calculated observables due to the ambiguities of these ingredients (the Igo ambiguity Igo58 of the optical potentials is a classical example of such ambiguities). While the classical response to this demand was the comparison of calculation using different ingredients (for example, optical potentials from different global parametrizations), this evaluation of the uncertainties is at most qualitative and highly dependent on the choice of inputs. Recently, some reaction theory groups, particularly the one at Michigan State University, have started to employ Bayesian methods to quantify the uncertainties that optical potentials suffer when they are obtained from the fitting of elastic scattering data (as is usual for optical potentials) and the propagation of these uncertainties to observables such as $(d,p)$ transfer angular or charge exchange $(p,n)$ differential cross sections Lov17b ; Lov18 ; Kin18 ; Kin19 ; Cat19 ; Cat20 ; Phi21 ; Whi22 or nucleon-knockout momentum distributions Heb23 . The determination of Bayesian posterior distributions and confidence intervals requires the evaluation of the reaction observables for a range of values for every considered parameter, which rapidly results in a humongous number of calculations, which are feasible in facilities with high computing power for light-computation models such as DWBA and ADWA, but become prohibitive for models which require heavier computations, such as CDCC or CRC. As such, these groups are exploring methods to accelerate these calculations using emulators Fur20 ; Dri21 to predict and interpolate the result of these heavier calculations for the required parameter values. Through these methods, uncertainty analyses have been performed using CDCC Sur22 , although admittedly only the uncertainties due to the valence-core interaction were studied, further proof of the computational magnitude of the task. In this regard, new global potential parametrizations with quantified uncertainties have recently been developed, with uncertainties originating either from the chiral interactions used to generate the potentials Whi21 or from fitting to extensive elastic scattering data Pru23 . The latter potential has been used to study the uncertainty in the reduction factors for transfer and knockout reactions Heb23b , which is still an open problem in Nuclear Physics Aum21 .
10 Inclusion of non-local potentials
It has long been known that nucleon-nucleus and nucleus-nucleus optical potentials are non-local Fes58a , both due to antisymmetrization and the effect of non-elastic channels (encoded in the second term of the Feshbach potential (3)). The use of non-local potentials transforms the Schrödinger equation into an integro-differential equation, which is far more cumbersome to solve, so local potentials have traditionally been favoured in reaction calculations. In fact, for a long time only a few non-local parametrizations for optical potentials have been published Per62 ; Gia76 ; Gia80 ; Tia15 and non-locality is not explicitly considered in the majority of potential parametrizations CH89 ; KD03 . In recent years there has been a development of dispersive optical potentials based on the nuclear self-energy, which is inherently non-local Cha07 ; Dic10 ; Mah16 ; Mor24 , thus reigniting the interest in fully non-local potentials.
For the study of elastic scattering, the neglect of non-locality is not problematic, as one can always find a local equivalent potential which reproduces the phase-shifts and observables of the non-local one Fie66 ; Fie67 . Analogously, reactions that are only sensitive to the asymptotics of the wave functions are usually well described by local potentials. However, non-locality reduces wavefunctions in the nuclear interior, in what is called the Perey effect Aus65 ; Fie66 , so non-locality should be taken into consideration in reactions that are sensitive to the nuclear interior. This has usually been approximated through the multiplication of wavefunctions obtained from a local potential by the so-called Perey factor Fie66 , which reduces the wavefunctions in the interior without altering their asymptotics. Although the Perey factor can be obtained for any non-locality shape, the prescriptions typically used assume a Gaussian dependence on non-locality Fra57 (usually referred to as Perey-Buck geometry). These approximations (local equivalent potential and Perey factor) have been compared to the computation of the wavefunction using the full non-local potential in Tit14 for $(d,p)$ transfer reactions, finding a difference in spectroscopic factors of $\sim 10\%$, although these results did not consider non-locality in the outgoing deuteron wavefunction.
[…]
11 Conclusions and perspectives
The present review aimed to provide a glimpse into the unique physics behind reactions involving weakly-bound nuclei and the formalisms and methods commonly used in the modeling and analysis of these reactions. We have shown the importance of breakup channels for weakly-bound systems and the methods developed to properly consider them, such as the CDCC method (Sec. 3), several semiclassical descriptions (Sec. 7), or the ADWA and CDCC-BA formalisms for transfer reactions (Sec. 4), as well as their extension to consider collective excitations and/or 3-body projectiles, where the particular characteristics of Borromean systems have been exploited to explore nucleon-nucleon correlations (Sec. 8). As well, inclusive breakup (Sec. 5) and incomplete and complete fusion (Sec. 6) reactions with weakly-bound systems have been explored due to the unique phenomenon of complete fusion suppression found for weakly-bound species. Thanks to the increase in computational power, demanding methods such as CDCC and IAV become more accessible and useful in the analysis of experimental data. In addition, questions which require larger numerical efforts, such as the inclusion of non-locality in these sophisticated methods (Sec. 10) and the study of theoretical uncertainties (Sec. 9), are starting to be explored.
Admittedly, the selection of reactions and methods presented in this work has been heavily curated by the expertise and preferences of the authors and many other methods for the study of reactions with weakly-bound nuclei exist and/or are currently being developed:
Some of the most promising advances are those related to few-body and ab-initio methods. The application of Faddeev method to nuclear reactions, while potentially providing an exact, rigorous solution of a given few-body Hamiltonian, has only been possible in recent years thanks to the increasing computational capabilities, the inclusion of complex optical potentials and the development of techniques to deal with the long-range Coulomb couplings. The method has been very useful for benchmarking some of the methods discussed in this review, such as the CDCC and ADWA methods Del07 ; Upa12 . It has also proven very useful for analyzing nucleon knockout reactions Cre11 ; Cre14 ; Cre16 .
There has also been remarkable progress in the No-Core-Shell-Model (NCSM) and its extension, the NCSM with Continuum (NCSMC) Bar13 . While originally developed to study nuclear structure, the method has been applied with great success to the calculation of scattering observables Vor19 ; Baroni13 ; Cal16 .
Ab-initio methods provide not only alternative ways for evaluating scattering observables by themselves, but they also furnish very useful structure inputs for traditional reaction frameworks. For example, the Variational Monte Carlo (VMC) Car15 and its variants, such as the Green’s function Monte Carlo and the Cluster Variational Monte Carlo Pie92 , provide microscopic overlap functions which can be directly plugged into reaction formalisms for transfer Fla13 and knockout Gri11 reactions.
The expansion of the experimental studies, reaching heavier exotic nuclei, has also shown the need to extend and upgrade the existing reaction models to incorporate the new features found in the structure and reactions of the newly discovered nuclei. For example, it has become clear that a proper description of newly discovered halo nuclei in the islands of inversion requires the simultaneous inclusion of deformation, pairing and Pauli blocking. Reaction formalisms, such as DWBA or CDCC, must therefore be extended to accommodate such effects. Admittedly, the XCDCC method discussed in this review has used so far relatively simple particle-plus-rotor models, and hence it is mandatory to resort to more sophisticated models.
The challenge is even larger for the case of three-body systems which, in addition to the properties cited above, require specific formalisms to account for their three-body nature. For example, knockout experiments with 14Be have clear evidenced the need of including the excitations of the 12Be “core” Cor19 ; Cor23 . Even beyond, models including four- and five-body systems may be necessary, as the recent discovery of a tetraneutron correlated system in 8He attests Due22 .
All of these avenues for advancement show that, despite its numerous and significant successes, the study of nuclear reactions with weakly-bound nuclei remains a developing and active field, pushed by theoretical and technical advances and by new experimental measurements in ever more exotic nuclei. It is only fitting that an overview on such a field will be incomplete and become quickly outdated, but we hope we have managed to transmit to the reader at least a glimpse of the craft behind the study of such a vast and fascinating topic.
Acknowledgements
The authors acknowledge financial support by MCIN/AEI/10.13039/501100011033 under I+D+i project No. PID2020-114687GB-I00. M.G.-R. acknowledges financial support by MCIN/AEI/10.13039/501100011033 under grant IJC2020-043878-I (also funded by “European Union NextGenerationEU/PRTR”).